### v1

# Random pairing MLE for estimation of item parameters in Rasch model

Abstract
The Rasch model, a classical model in the item response theory, is widely used in psychometrics to model the relationship between individuals’ latent traits and their binary responses on assessments or questionnaires. In this paper, we introduce a new likelihood-based estimator—random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$) and its bootstrapped variant multiple random pairing MLE ($\mathsf{MRP\text{-}MLE}$) that faithfully estimate the item parameters in the Rasch model. The new estimators have several appealing features compared to existing ones. First, both work for sparse observations, an increasingly important scenario in the big data era. Second, both estimators are provably minimax optimal in terms of finite sample $\ell_{\infty}$ estimation error. Lastly, $\mathsf{RP\text{-}MLE}$ admits precise distributional characterization that allows uncertainty quantification on the item parameters, e.g., construction of confidence intervals of the item parameters. The main idea underlying $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ is to randomly pair user-item responses to form item-item comparisons. This is carefully designed to reduce the problem size while retaining statistical independence. We also provide empirical evidence of the efficacy of the two new estimators using both simulated and real data.
1 Introduction
The item response theory (IRT) [ER13] is a framework widely used in psychometrics to model the relationship between individuals’ latent traits (such as ability or personality) and their responses on assessments or questionnaires. It is particularly useful in the development, analysis, and scoring of tests and assessments; see a recent survey [CLLY21] for a statistical account of IRT.
Among statistical models in IRT, the Rasch model [Ras60] is a simple but fundamental one for modeling binary responses. Specifically, for a user $t$ (e.g., test-taker) and an item $i$ (e.g., test problem), Rasch model assumes that the response user $t$ has to item $i$ is binary and obeys
$\mathbb{P}\left[\text{user $t$ ``beats'' item $i$}\right]=\frac{e^{\zeta_{t}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}},$
where $\zeta_{t}^{\star},\theta_{i}^{\star}\in\mathbb{R}$ are latent traits of user $t$ and item $i$, respectively. The term “beats” here refers to positive responses, such as answering an exam question correctly, writing a favorable review of a product, etc. We also call this response a comparison between user $t$ and item $i$.
In this paper, we focus on estimating the item parameters $\bm{\theta}^{\star}$, which is one of the four main statistical tasks surrounding the Rasch model (or IRT more generally) listed in [CLLY21]. Estimating the item parameters is also quite useful in practice. For instance, in education testing, $\bm{\theta}^{\star}$ could reveal the difficulty of the exam questions, while in product reviews, $\bm{\theta}^{\star}$ could reveal the popularity of the products. Various methods have been proposed for estimating the item parameters $\bm{\theta}^{\star}$, including the joint maximum likelihood estimator (JMLE), the marginal maximum likelihood estimator, the conditional maximum likelihood estimator (CMLE), as well as the spectral estimator recently proposed in [NZ22, NZ23]. We refer readers to a recent article [Rob21] for comparisons between different item parameter estimation methods. However, three main gaps remain in tackling item estimation in the Rasch model:
  • Non-asymptotic guarantee. Apart from the recently proposed spectral estimator [NZ22, NZ23], most theoretical guarantees for the likelihood-based estimators are asymptotic. Since all the estimation procedures are necessarily applied with finite samples, the asymptotic guarantee alone fails to inform practitioners about the performance of different estimators when working with a limited number of samples.
  • Sparse observations. It is not uncommon to encounter situations where each user only responds to a handful of questions (or items). This brings the challenge of incomplete and sparse observations. Among the likelihood based estimators, CMLE is known to have numerical issues with incomplete observations [Mol95]. While the spectral estimator [NZ22, NZ23] is capable of handling incomplete observations, it still requires the observations to be relatively dense. We will elaborate on this point later.
  • Uncertainty quantification. Beyond estimation, uncertainty quantification on the item parameters is central to realizing the full potential of the Rasch model. However, existing results do not address this problem under sparse observations. An exception is the recent work [CLOX23], which is based on joint estimation and inference on the item parameters $\bm{\theta}^{\star}$ and the user parameters $\bm{\zeta}^{\star}$. Their sampling scheme is restrictive and requires a relatively dense sampling rate.
In light of the above gaps, we raise the following natural question:
Can we develop an estimator for the item parameters $\bm{\theta}^{\star}$ that (1) enjoys optimal estimation guarantee in finite sample, and (2) is amenable to tight uncertainty quantification, when the observations are sparse?
1.1 Main contributions
The main contribution of our work is the proposal of a novel estimator named random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$ in short) that achieves the two desiderata listed above.
In essence, $\mathsf{RP\text{-}MLE}$ compiles user-item comparisons to item-item comparisons by randomly pairing responses of the same user to different items. This pairing procedure is carefully designed to extract information of the item parameters while retaining statistical independence. After this compilation step, item parameters $\bm{\theta}^{\star}$ are estimated by the MLE $\widehat{\bm{\theta}}$ given the item-item comparisons.
Even when the observations are extremely sparse, $\mathsf{RP\text{-}MLE}$ achieves the following:
  • Regarding estimation, we show that both $\mathsf{RP\text{-}MLE}$ and its bootstrapped version enjoy optimal finite sample $\ell_{\infty}$ error guarantee. Compared to the conventional $\ell_{2}$ error guarantee, the $\ell_{\infty}$ guarantee, as an entrywise guarantee is more fine-grained. Consequently, we also show that $\mathsf{RP\text{-}MLE}$ can recover the top-$K$ items with minimal sample complexity.
  • While the optimal $\ell_{\infty}$ error guarantee directly yields optimal $\ell_{2}$ guarantee, such guarantee is only correct in an order-wise sense. We provide a refined finite-sample $\ell_{2}$ error guarantee of $\mathsf{RP\text{-}MLE}$ that is precise even in the leading constant.
  • Supplementing the estimation guarantee, we also build an inferential framework based on $\mathsf{RP\text{-}MLE}$ $\widehat{\bm{\theta}}$. More specifically, we precisely characterize the asymptotic distribution of $\widehat{\bm{\theta}}$. This result facilitates several inferential tasks such as hypothesis testing and construction of confidence regions of $\bm{\theta}^{\star}$.
We test our methods on both synthetic and real data, which clearly show competitive empirical estimation performance. The inferential result on synthetic data also closely matches our theoretical predictions.
1.2 Prior art
Item response theory.
The item response theory is a popular statistical framework for modeling response data. It often involves a probabilistic model that links categorical responses to latent traits of both users and items. Early endeavors include [Ras60] that introduces the Rasch model studied herein and [LNB68] that describes a more general framework using parametric models. Popular IRT models include the Rasch model, the two-parameter model (2PL), and the three-parameter logistic model (3PL). As response data widely appears in real life, IRT finds application in numerous fields including educational assessment [DC10], psychometrics [LNB68], political science [VHSA20], and medical assessment [FBC05]. See also [CLLY21] for an overview of IRT from statisticians’ perspective.
Latent score estimation for Rasch model.
An important statistical question in the Rasch model is to estimate the latent parameters of the items. As the Rasch model is an explicit probabilistic model, many methods are based on the principle of maximum likelihood estimation. For instance, marginal MLE assumes a prior on the user parameters that is either given or optimized within a parametric distribution family. The item parameter is then estimated by maximizing the marginal likelihood. A drawback is that MMLE relies on a good prior. On the other hand, joint MLE (JMLE) makes no distributional assumption and maximizes the joint likelihood w.r.t. both the item and user parameters. However, it is not consistent for estimating the item parameters when the number $m$ of items is fixed [Gho95]. Interested readers may also consult [Lin99] for an overview of other classical estimators.
Several methods are more relevant to our proposed estimator $\mathsf{RP\text{-}MLE}$ as they follow a similar philosophy to form item-item pairs from user-item responses. Pseudo MLE (PMLE) [Zwi95] maximizes the sum of the log-likelihood of all pairs of responses from the same users to different items. However, due to the dependency issue, no satisfying finite sample performance guarantee has been established. Another related approach is the spectral method, in which a Markov chain on the space of items is formed and the item parameters are estimated via the stationary distribution of the Markov chain. The most recent works in this category are [NZ22, NZ23], which essentially use the same idea as pseudo MLE in forming item-item comparisons.
The Bradley-Terry-Luce model with sparse comparisons.
An informed reader may realize that the Rasch model resembles the Bradley-Terry-Luce (BTL) model [Luc59, BT52] in the ranking literature. Indeed, one can view the Rasch model as a special case of the BTL model that distinguishes the two groups of users and items, and only allows inter-group comparisons. There has been a recent surge in interest in studying top-$K$ ranking in the BTL model [CS15, JKSO16, CFMW19, CGZ22, GSZ23] and its extensions [FHY22, FLWY22, FLWY23], especially under sparse observations of the pairwise responses. Most notably, under a uniform sampling scheme, [CFMW19] shows that (regularized) MLE and spectral methods are both optimal in top-$K$ ranking and [GSZ23] provides inference results for both methods. Going beyond uniform sampling, [Che23, LSR22] investigate the performance of MLE in the BTL model with a general comparison graph and later [YCOM24] improves the analysis to show the optimality of MLE for the BTL model in both uniform and semi-random sampling.
Notation.
For a positive integer $n$, we denote $[n]=\{1,2,\ldots,n\}$. For any $a,b\in\mathbb{R}$, $a\wedge b$ means the minimum of $a,b$ and $a\vee b$ means the maximum of $a,b$. We use $\overset{\mathrm{d}}{\rightarrow}$ to denote convergence in distribution. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$, we use $\lambda_{1}(\bm{A})\geq\lambda_{2}(\bm{A})\geq\cdots\geq\lambda_{n}(\bm{A})$ to denote its eigenvalues and $\bm{A}^{\dagger}$ to denote its Moore-Penrose pseudo-inverse. For symmetric matrices $\bm{A},\bm{B}\in\mathbb{R}^{n\times n}$, $\bm{A}\preceq\bm{B}$ means $\bm{B}-\bm{A}$ is positive semidefinite, i.e., $\bm{v}^{\top}(\bm{B}-\bm{A})\bm{v}\geq 0$ for any $\bm{v}\in\mathbb{R}^{n}$. We use $\bm{e}_{i}$ to denote the standard unit vector with $1$ at $i$-th coordinate and 0 elsewhere. Unless specified otherwise, $\log(\cdot)$ denotes the natural log.
2 Problem setup and new estimator $\mathsf{RP\text{-}MLE}$
In this section, we first introduce the formal setup of the item parameter estimation problem in the Rasch model. Then we present the newly proposed estimator $\mathsf{RP\text{-}MLE}$ along with the rationale behind its development.
2.1 Problem setup
The Rasch model considers pairwise comparisons between elements of two groups: users and items. Let $n$ (resp. $m$) be the number of users (resp. items). Rasch assumes a user parameter $\bm{\zeta}^{\star}\in\mathbb{R}^{n}$ and an item parameter $\bm{\theta}^{\star}\in\mathbb{R}^{m}$ that measures the latent traits (e.g., difficulty of a problem) of users and items, respectively. For a subset of possible user-item pairs $\mathcal{E}_{X}\subset[n]\times[m]$, we observe binary responses $\{X_{ti}\}_{(t,i)\in\mathcal{E}_{X}}$ obeying
$\mathbb{P}[X_{ti}=1]=\frac{e^{\theta_{i}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}}.$ (1)
Here $X_{ti}=1$ means user $t$ has negative response against item $i$ (e.g., unable to solve a problem). The goal is to estimate $\bm{\theta}^{\star}$, the item parameters.
To model sparse observations/comparisons, we assume that $\mathbb{P}[(t,i)\text{ is compared}]=p$ independently for every $(t,i)\in[n]\times[m]$. To put it in the language of graph theory, we denote the associated bipartite comparison graph to be $\mathcal{G}_{X}=(\mathcal{V}_{X},\mathcal{E}_{X})$, where $\mathcal{V}_{X}$ consists of $n$ users and $m$ items. Then essentially, we are assuming that the bipartite graph follows an Erdős-Rényi random model. [^1] Alternatively, we can assume each user responds to $mp$ items uniformly at random. Our estimator and performance guarantee continue to work in this sampling scheme.
Before moving on, we introduce the condition numbers that characterize the range of the latent traits. Let $\kappa_{1}$, $\kappa_{2}$, and $\kappa$ be defined by $\log\;(\kappa_{1})=\max_{ij}\{|\theta_{i}^{\star}-\theta_{j}^{\star}|\}$, $\log(\kappa_{2})=\max_{ti}\{|\zeta_{t}^{\star}-\theta_{i}^{\star}|\}$, and $\kappa=\max\{\kappa_{1},\kappa_{2}\}$, respectively.
2.2 Random pairing maximum likelihood estimator
Algorithm 1 Random Pairing Maximum Likelihood Estimator ($\mathsf{RP\text{-}MLE}$)
  1. 1.
    For each tester $t$,
    1. (a)
      Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems.
    2. (b)
      For each $(i,j)\in[m]\times[m]$, do the following:
      1. i.
        If $(i,j)$ is selected as a pair in Step 1(a), $R_{ij}^{t}=1$. Furthermore, if $X_{ti}\neq X_{tj}$, let $Y_{ij}^{t}=\mathds{1}\{X_{ti}<X_{tj}\}$ and $L_{ij}^{t}=1$; if $X_{ti}=X_{tj}$, let $L_{ij}^{t}=0$.
      2. ii.
        If $(i,j)$ is not selected as a pair in Step 1(a), let $L_{ij}^{t}=0$ and $R_{ij}^{t}=0$.
  2. 2.
    Let $\mathcal{E}_{Y}$ be a set of edges defined by $\mathcal{E}_{Y}\coloneqq\{(i,j):\sum_{t=1}L_{ij}^{t}\geq 1\}$ and let $\mathcal{G}_{Y}=([m],\mathcal{E}_{Y})$. For each $(i,j)\in\mathcal{E}_{Y}$, let $L_{ij}\coloneqq\sum_{t=1}^{n}L_{ij}^{t}$ and $Y_{ij}\coloneqq(1/L_{ij})\sum_{\{t:L_{ij}^{t}=1\}}Y_{ij}^{t}$.
  3. 3.
    Compute MLE on $Y_{ij}$, i.e.,
    $\widehat{\bm{\theta}}\coloneqq\arg\min_{\bm{\theta}:\bm{1}_{m}^{\top}\bm{\theta}=0}\mathcal{L}(\bm{\theta})$
    where
    $\mathcal{L}(\bm{\theta})=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\left(-Y_{ji}(\theta_{i}-\theta_{j})+\log(1+e^{\theta_{i}-\theta_{j}})\right).$ (2)
  4. 4.
    Return the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}$.
In this section, we present our main method $\mathsf{RP\text{-}MLE}$; see Algorithm 1. The algorithm can be divided into two parts. The first part—Steps 1 and 2—uses random pairing to compile the observed user-item responses $\bm{X}\in\mathbb{R}^{n\times m}$ to item-item comparisons $\bm{Y}\in\mathbb{R}^{m\times m}$. The second part—Steps 3 and 4—computes a standard MLE on the item-item comparisons. Some intuitions regarding the development of $\mathsf{RP\text{-}MLE}$ are in order.
Random pairing to construct item-item comparisons.
The idea of pairing is that by matching the responses $X_{ti}$ with $X_{tj}$, we form a comparison between items $i$ and $j$ to directly extract information of item parameters $\theta_{i}^{\star}$ and $\theta_{j}^{\star}$. More specifically, the item-item comparisons $\bm{Y}$ follow the Bradley-Terry-Luce model [BT52, Luc59],i.e., $\mathbb{P}[Y_{ij}^{t}]=e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})$; see Section 4.1 for a formal argument.
By compiling user-item responses to item-item comparisons, we reduce the size of the data matrix from $n\times m$ to $m\times m$, and also the number of intrinsic parameters from $n+m$ to $m$, since the likelihood function (2) of $Y_{ij}^{t}$ is completely independent with the user parameter $\zeta_{t}^{\star}$.
More importantly, the pairing is performed in a disjoint fashion. This ensures that all constructed item-item comparisons $Y_{ij}^{t}$ are independent with each other; see Section 4.1 for a formal statement. This is the key ingredient that enables us to improve over previous implementation of item-item comparisons, such as pseudo-likelihood [Cho82, Zwi95] and spectral methods [NZ22, NZ23].
A variant via bootstrapping.
A drawback of this random pairing is that it potentially induces a loss of information since not every possible pairing is considered. Once $X_{ti}$ is paired with $X_{tj}$, we cannot pair $X_{ti}$ with another response $X_{tl}$. That being said, we will later show that the $\ell_{\infty}$ error of $\mathsf{RP\text{-}MLE}$ is still rate-optimal up to logarithmic factors. Hence the loss of information can at most incur a small constant factor in terms of estimation error. Nevertheless, we provide a remedy to this phenomenon in $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by running (in other words, bootstrapping) the $\mathsf{RP\text{-}MLE}$ multiple times with different random data splitting and averaging the resulting estimates. $\mathsf{MRP\text{-}MLE}$ trivially enjoys the same theoretical guarantee (cf. Theorem 1) while improving the estimation error in practice over $\mathsf{RP\text{-}MLE}$. See Figure 4 in Section 5.1 for the empirical evidence.
Algorithm 2 Multiple Random Pairing Maximum Likelihood Estimator ($\mathsf{MRP\text{-}MLE}$)
  1. 1.
    Let $n_{\mathrm{split}}$ be the number of runs. For $i=1,\ldots,n_{\mathrm{split}}$, run $\mathsf{RP\text{-}MLE}$ (Algorithm 1), each time with an independent random splitting in Step 1. Let the estimated latent scores be $\widehat{\bm{\theta}}^{(i)}$.
  2. 2.
    Estimate the latent score with
    $\widehat{\bm{\theta}}=\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}^{(i)}.$
  3. 3.
    Estimate the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}$.
3 Main results
In this section, we collect the main theoretical guarantees for $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$. Section 3.1 focuses on the finite sample $\ell_{\infty}$ error bound. While one can translate $\ell_{\infty}$ error into an $\ell_{2}$ bound, in Section 3.2, we present a much sharper characterization of the $\ell_{2}$ error of $\mathsf{RP\text{-}MLE}$. Lastly in Section 3.3, we provide a distributional characterization of $\mathsf{RP\text{-}MLE}$.
3.1 $\ell_{\infty}$ estimation error and top-$K$ recovery
Without loss of generality, we assume that the scores of the items are ordered, i.e., $\theta_{1}^{\star}\geq\theta_{2}^{\star}\geq\cdots\geq\theta_{m}^{\star}$, and denote $\Delta_{K}\coloneqq\theta_{K}^{\star}-\theta_{K+1}^{\star}$. In words, $\Delta_{K}$ measures the difference between the difficulty levels of items $K$ and $K+1$. The following theorem provides $\ell_{\infty}$ error bounds and top-$K$ recovery guarantee for both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$. We defer its proof to Section 4.2.
Theorem 1.
Suppose that $mp\geq 2$ and $np\geq C_{1}\kappa_{1}^{4}\kappa_{2}^{5}\log^{3}(n)$ for some sufficiently large constant $C_{1}>0$. Suppose that there exists some constant $\alpha>0$ such that $m\leq n^{\alpha}$. Let $\widehat{\bm{\theta}}$ be the $\mathsf{RP\text{-}MLE}$ estimator. With probability at least $1-O(n^{-10})$, $\widehat{\bm{\theta}}$ satisfies
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}\leq C_{2}\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{\log(n)}{np}}.$
Consequently, the estimator is able to exactly recover the top-$K$ items as soon as
$np\geq\frac{C_{3}\kappa_{1}^{2}\kappa_{2}^{1}\log(n)}{\Delta_{K}^{2}}.$
Here $C_{2},C_{3}>0$ are some universal constants. All the claims continue to hold for $\mathsf{MRP\text{-}MLE}$ as long as there exists some constant $\beta>0$ such that $n_{\mathrm{split}}\leq n^{\beta}$.
Some remarks are in order.
Finite sample minimax optimality.
Intuitively, as we have $\Theta(mnp)$ sampled comparisons and $m$ parameters to estimate, the optimal $\ell_{\infty}$ error should be $O(\sqrt{m/(mnp)})=O(1/\sqrt{np})$. This is formalized in the following result from [NZ23].
Proposition 1 (Minimax lower bound, Theorems 3.3 and 3.4 in [NZ23]).
Assume that $np\geq C_{1}$ for some suffiently large constant $C_{1}>0$. For any $n$ and $m$, there exists a class of user and item parameters $\Theta$ such that
$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{2}^{2}\geq\frac{C_{2}m}{np},\qquad\text{and}\qquad\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{\infty}^{2}\geq\frac{C_{2}}{np},$
where $C_{2}>0$ is some constant. Moreover if $np\leq C_{K}\log(m)/\Delta_{K}^{2}$ for some constant $C_{K}>0$, we have
$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{P}\left[\widehat{\bm{\theta}}\text{ fails to identify all top-}K\text{ items}\right]\geq\frac{1}{2}.$
Comparing our upper bounds with the lower bound in the proposition, we can see that both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ are rate-optimal in $\ell_{\infty}$ estimation error and top-$K$ recovery sample complexity, up to logarithmic and $\kappa$ factors.
Sample size requirement.
While the rates are optimal, it is worth noting that in Theorem 1 we have made several sample size requirements. We now elaborate on them.
First, the assumption $mp\geq 2$ is a mild requirement on the expected number of items compared by each user. This is required as we need user $t$ to compare at least two items to form a comparison between items. In fact, if a user only responds to one item, it is clear that this data point is not useful at all for item parameter estimation.
Second, it is a standard and necessary requirement to have $np\gtrsim\log(n)$ to make sure that each item is compared to at least one user with high probability. In Theorem 1 we require an extra $\log^{2}(n)$ factor to suppress a quadratic error term that comes up in the analysis. This cubic log factor can possibly be loose, but it is a minor issue and we leave it as a problem for future research.
Lastly, $m\leq n^{\alpha}$ and $n_{\mathrm{split}}\leq n^{\beta}$ are both minor as we only need these to allow union bounds over $m$ and $n_{\mathrm{split}}$.
Comparison with [NZ23].
The closest result to our paper in terms of $\ell_{\infty}$ guarantee for the Rasch model appears in the recent work [NZ23]. Their spectral method uses a similar construction of the item-item comparisons but without disjoint pairing. To provide detailed comparisons, we restate their results below.
Proposition 2 (Informal, Theorem 3.1 in [NZ23]).
Assume that $p\gtrsim\log(m)/\sqrt{n}$ and $mp\gtrsim\log(n)$, with probability at least $1-O(m^{-10}+n^{-10})$, spectral estimator $\widehat{\bm{\theta}}_{\mathrm{spectral}}$ satisfies
$\|\widehat{\bm{\theta}}_{\mathrm{spectral}}-\widehat{\bm{\theta}}\|_{\infty}\lesssim\kappa^{9}\sqrt{\frac{\log(m)}{np}}.$
This error rate is similar to ours. However, the required sample size is much larger as they require $p\gtrsim\log(m)/\sqrt{n}$. Our result makes a significant improvement by allowing a much smaller sampling rate $p$, cf. $mp\geq 2$ and $np\gtrsim\log^{3}(n)$. In fact, as we have argued earlier, it is nearly the sparsest possible. In addition, our methods enjoy a significantly better error rate dependency on $\kappa$. In Section 5.1, we provide empirical evidence for this improvement: when $\kappa$ is large, $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ outperform the spectral methods in [NZ23].
3.2 Refined $\ell_{2}$ error characterization
The $\ell_{\infty}$ error guarantee in Theorem 1 immediately implies an $\ell_{2}$ error bound
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|\leq C\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{m\log(n)}{np}},$ (3)
which is rate-optimal compared to the minimax lower bound in Proposition 1. However, this guarantee is only correct in an order-wise sense. In this section, we present a refined characterization of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ that is precise in the leading constant.
We start with necessary notation. Let $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$ and $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$. Define a matrix
$\bm{L}_{L\tilde{z}}=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (4)
for $\tilde{z}\in\{z,\widehat{z}\}$ and let $\bm{L}_{L\tilde{z}}^{\dagger}$ be its pseudo-inverse. One can view $\bm{L}_{L\tilde{z}}$ as a weighted graph Laplacian. In the following theorem, we show that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrates tightly around $[\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})]^{1/2}$. We defer the analysis to Section 4.3 and the complete proof to Section D.
Theorem 2.
Instate the assumptions of Theorem 1. Then for some constants $C_{1},C_{2}>0$, with probability at least $1-O(n^{-10})$, we have
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np},$ (5)
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}(\kappa_{1}^{6}+\kappa_{1}^{7/2}\kappa_{2}^{2})\sqrt{m}\log^{2}(n)}{np}.$ (6)
Theorem 2 is more refined compared to (3). First, it provides both upper and lower bounds for $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$. Second, there is no hidden constant in front of the leading term $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$. In addition, inspecting the proof of Theorem 2, we see that
$\sqrt{\frac{m-1}{np}}\leq\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\leq 4\kappa_{1}^{1/2}\kappa_{2}^{1/2}\sqrt{\frac{m}{np}},$
and the same holds for $[\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})]^{1/2}$. Consequently, the right hand sides of both (5) and (6) are lower order terms compared to $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$ when $n,m\rightarrow\infty$. Indeed this recovers the naive $\ell_{2}$ bound (3) under appropriate sample size assumptions.
3.3 Uncertainty quantification
Now we move on to uncertainty quantification about the $\mathsf{RP\text{-}MLE}$ estimator $\widehat{\bm{\theta}}$ in order to perform, say, hypothesis testing on the item parameters $\bm{\theta}^{\star}$. Our starting point is the following characterization of the limiting distribution of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$. The analysis is deferred to Section 4.3 and the complete proof is deferred to Section E.
Theorem 3.
Instate the assumptions of Theorem 1 and fix $m$. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
$\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
Though simple, we record the following immediate consequence on the asymptotic normality of $\widehat{\theta}_{i}-\theta_{i}^{\star}$ for any single coordinate $i$.
Corollary 1.
Instate the assumptions of Theorem 1. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$ (7)
We compare our result with that in [CLOX23]. Theorem 8 therein studies the Rasch model and provides an inference result for a joint estimator of $\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}$. However, it requires a dense sampling scheme with $p\gtrsim 1/\sqrt{\log(n)/n}$ and $p\gtrsim n\log^{2}(n)/m^{2}$. It also requires that both $m$ and $n$ tend to infinity. These two assumptions are more restrictive than ours.
Theorem 3 is quite general as it allows various applications to inference. As an illustrative example, we demonstrate how to construct confidence intervals for a particular item parameter $\theta_{i}^{\star}$.
Fix any $\alpha>0$, the level of significance. Define
$\mathcal{C}_{i}^{+}(\alpha)$ $\coloneqq\widehat{\theta}_{i}+\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
$\mathcal{C}_{i}^{-}(\alpha)$ $\coloneqq\widehat{\theta}_{i}-\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
where $\Phi$ is the cdf for the standard normal distribution. Corollary 1 implies the following proposition.
Proposition 3.
Suppose the statement in Theorem 3 holds. Then for any $(i,j)$, we have
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\geq\mathcal{C}_{i}^{+}(\alpha)\right]$ $=\alpha,$ (8a)
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\leq\mathcal{C}_{i}^{-}(\alpha)\right]$ $=\alpha,$ (8b)
and
$\lim_{np\rightarrow\infty}\mathbb{P}\left[\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\right]=\alpha.$ (9)
This proposition can be extended to confidence intervals for any linear combination of $\bm{\theta}^{\star}$, such as $\theta_{i}^{\star}-\theta_{j}^{\star}$. These confidence intervals naturally enable hypothesis testings such as
$\mathcal{H}_{0}:\theta_{j}^{\star}\geq\theta_{i}^{\star},\qquad\mathcal{H}_{1}:\theta_{j}^{\star}<\theta_{i}^{\star}$
and
$\mathcal{H}_{0}:\text{item $i$ is a top-$K$ item},\qquad\mathcal{H}_{1}:\text{item $i$ is not a top-$K$ item}.$
Various other inference tasks have been considered in the ranking literature [LFL23][FLWY22, FLWY23]. We expect our general distributional characterization in Theorem 3 to be readily applied for these tasks.
4 Analysis
In this section, we present the main steps in the analysis to obtain theoretical results in the previous section. Section 4.1 gives a complete argument on the reduction to the BTL model we alluded in Section 2.2, Section 4.2 provides the analysis of the $\ell_{\infty}$ error, and Section 4.3 provides the analysis of both the $\ell_{2}$ error and the asymptotic distribution as they share similar components in the analysis.
4.1 Reduction to Bradley-Terry-Luce model
A key idea in $\mathsf{RP\text{-}MLE}$ is the randomized pairing in Steps 1 and 2 of Algorithm 1. It compiles the user-item responses $\bm{X}$ to item-item comparisons $\bm{Y}$. In this section, we make a detailed argument that $\bm{Y}$ follows the Bradley-Terry-Luce model with a non-uniform sampling scheme.
Recall that $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}$ and $Y_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}<X_{tj}\}$. The following fact provides the distribution of $Y_{ij}^{t}$ conditional on $L_{ij}^{t}=1$. We defer its proof to Section A.
Fact 1.
Let $i,j$ be two items and $t$ be a user. Suppose that user $t$ has responded to both items $i$ and $j$. Let $X_{ti}$ and $X_{tj}$ be the responses sampled from the probability model (1). Then we have
$\mathbb{P}[X_{ti}<X_{tj}\mid L_{ij}^{t}=1]=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}},$
and
$\mathbb{P}[L_{ij}^{t}=1]\geq\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}.$ (10)
Fact 1 shows that conditional on $L_{ij}^{t}=1$, $Y_{ij}^{t}$ follows the BTL model with parameters $\bm{\theta}^{\star}$. More importantly, as we deploy random pairing (cf. Step 1a), each response $X_{ti}$ is used at most once. As a result, conditional on $\{L_{ij}^{t}\}_{ijt}$, $Y_{ij}^{t}$’s are jointly independent across users and items. In light of these, we can equivalently describe the data generating process of $\bm{Y}$ as follows:
  1. 1.
    For each user-item pair $(t,i)$, there is a comparison between them with probability $p$ independently.
  2. 2.
    Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems. (Step 1(a) of Algorithm 1)
  3. 3.
    For all $(t,i,j)$, items $i$ and $j$ are compared by user $t$ if $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}=1$.
  4. 4.
    Conditioned on $L_{ij}^{t}=1$, one observes the outcome $Y_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}<X_{tj}\}$.
Steps 1–3 generates a non-uniform comparison graph $\mathcal{E}_{Y}$ between items. Step 4 reveals the independent outcomes of these comparisons following the BTL model, conditional on the graph $\mathcal{E}_{Y}$. This justifies that (2) is truly the likelihood function of the BTL model conditional on the comparison graph $\mathcal{E}_{Y}$.
4.2 Analysis of $\ell_{\infty}$ error
We have seen that analyzing $\mathsf{RP\text{-}MLE}$ under the Rasch model can be reduced to analyzing the MLE under the BTL model. This reduction allows us to invoke the general theory of MLE in the BTL model established in the recent work [YCOM24].
To facilitate the presentation, we introduce the necessary notation. For any $i\in[m]$, let $d_{i}\coloneqq\sum_{j:j\neq i}L_{ij}$ be the weighted degree of item $i$ in $\mathcal{G}_{Y}$ and $d_{\max}=\max_{i\in[m]}d_{i}$. Let the weighted graph Laplacian $\bm{L}_{L}$ be
$\bm{L}_{L}\coloneqq\sum_{i,j:i>j}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
The following lemma adapts Theorem 3 of the recent work [YCOM24] to our setting.
Lemma 1 (Theorem 3 in [YCOM24]).
Assume that $\mathcal{G}_{Y}$ is connected, and that
$[\lambda_{m-1}(\bm{L}_{L})]^{5}\geq C_{1}\kappa_{1}^{4}(d_{\max})^{4}\log^{2}(n)$ (11)
for some large enough constant $C_{1}>0$. Then with probability at least $1-n^{-10}$, we have
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$
for some constant $C_{2}>0$.
To leverage this general result, we need to characterize the spectral and degree properties of the comparison graph $\mathcal{G}_{Y}$, which is achieved in the following two lemmas. The proofs are deferred to Section B.
Lemma 2 (Degree bound in $\mathcal{G}_{Y}$).
Suppose that $np\geq C\kappa_{2}^{2}\log(n)$ for some large enough constant $C>0$ and $m\leq n^{\alpha}$ for some sufficiently large constant $\alpha>0$. With probability at least $1-2n^{-10}$, for all $i\in[m]$,
$\frac{1}{24\kappa_{2}}np\leq d_{i}\leq\frac{3}{2}np.$ (12)
Lemma 3.
Suppose $mp\geq 2$, $np\geq C\kappa_{2}^{2}\log(n)$ for some large enough constant $C$, and $m\leq n^{\alpha}$ for some constant $\alpha>0$. With probability at least $1-10n^{-10}$, we have
$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L})\leq\lambda_{1}(\bm{L}_{L})\leq 3np$ (13)
and
$\frac{np}{16\kappa_{1}\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{Lz})\leq\lambda_{1}(\bm{L}_{Lz})\leq np.$ (14)
4.2.1 Proof of Theorem 1
Now we are ready to prove Theorem 1. We focus on analyzing $\mathsf{RP\text{-}MLE}$, as the analysis of $\mathsf{MRP\text{-}MLE}$ follows immediately from the union bound of the different data splitting and the triangular inequality:
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}\leq\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\|\widehat{\bm{\theta}}^{(i)}-\bm{\theta}^{\star}\|_{\infty}.$
By assumption we have $mp\geq 2$ and $np\geq C_{1}\kappa_{1}^{4}\kappa_{2}^{5}\log^{3}(n)$ for some constant $C_{1}>0$. Then we can apply Lemmas 3 and 2 to see that
$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L}),\qquad\text{and}\qquad d_{\max}\leq\frac{3}{2}np.$
We observe that (11) is satisfied as long as $np\geq C_{1}\kappa_{1}^{4}\kappa_{2}^{5}\log^{3}(n)$ for some constant $C_{1}$ that is large enough. Invoking Lemma 1, we conclude that
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}\leq 2C_{2}\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{\log(n)}{np}}.$
It remains to show the top-$K$ recovery sample complexity. As $\theta_{1}^{\star}\geq\cdots\geq\theta_{K}^{\star}>\theta_{K+1}^{\star}\geq\cdots\geq\theta_{m}^{\star}$ by assumption, it suffices to show $\widehat{\theta}_{i}-\widehat{\theta}_{j}>0$ for any $i\leq K$ and $j>K$. Using the $\ell_{\infty}$ error bound, we have that
$\widehat{\theta}_{i}-\widehat{\theta}_{j}$ $\geq\left(\theta_{i}^{\star}-\theta_{j}^{\star}\right)-\left|\widehat{\theta}_{i}-\theta_{i}^{\star}\right|-\left|\widehat{\theta}_{j}-\theta_{j}^{\star}\right|\geq\Delta_{K}-4C_{2}\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{\log(n)}{np}}.$
Then $\widehat{\theta}_{i}-\widehat{\theta}_{j}>0$ as long as
$np\geq\frac{16C_{2}^{2}\kappa_{1}^{2}\kappa_{2}\log(n)}{\Delta_{K}^{2}}.$
4.3 Analysis for refined $\ell_{2}$ error and uncertainty quantification
To make the main text concise, we sketch the key ideas that lead to both Theorems 2 and 3.
We start with some necessary notation. For convenience, for any $(i,j)$, renumber the set of comparisons $\{Y_{ji}^{t}\}_{t:L_{ij}^{t}=1}$ as $\{Y_{ji}^{(l)}\}_{l=1,\ldots,L_{ij}^{t}}$ in an arbitrary order. Let $\epsilon_{ij}^{(l)}\coloneqq Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$ where $\sigma(x)=e^{x}/(1+e^{x})$ is the sigmoid function. Let $\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ (repeat $L_{ij}$ times for edge $(i,j)$), where $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$ is the total number of observed effective comparisons in $\mathcal{G}_{Y}$. Define
$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{(l)}/\sqrt{z_{ij}},\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y},l=1,\ldots,L_{ij}^{t}}\in\mathbb{R}^{L_{\mathrm{total}}}.$
The key step in establishing Theorems 2 and 3 is the following refined characterization of the estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$.
Lemma 4.
Instate the assumptions of Theorem 1. With probability at least $1-O(n^{-10})$, the estimator $\widehat{\bm{\theta}}$ given by the Algorithm (1) can be written as
$\widehat{\bm{\theta}}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r},$ (15)
where $\bm{r}\in\mathbb{R}^{m}$ is a random vector obeying $\|\bm{r}\|_{\infty}\leq C\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C>0$.
As the residual term $\bm{r}$ has small magnitude, we may analyze the properties of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ by focusing on the leading term $-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$.
Proof outline of Theorem 2.
Using the approximation $\widehat{\bm{\theta}}-\bm{\theta}^{\star}\approx-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$, we can turn our attention to the concentration of
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|^{2}\approx\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}=\widehat{\bm{\epsilon}}^{\top}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}.$
This is a quadratic form of sub-Gaussian random variables. The Hanson-Wright inequality tells us that $\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}$ concentrates around its expectation—$\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})$. The complete proof is deferred to Section D for Theorem 2.
Proof outline of Theorem 3.
With Lemma 4 in hand, we can expand $\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ to reach
$\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ $\approx-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$
$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
The main component of the last expression is a sum of bounded random variables $\bm{u}_{ij}^{(l)}$ that are independent conditional on $\mathcal{G}_{Y}$. This sum has zero mean and variance
$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
Therefore applying multivariate CLT, we obtain
$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
See the complete proof in Section E.
4.4 Proof outline of Lemma 4
Due to the importance and usefulness of Lemma 4, we provide a sketch of its proof and leave the full one to Section C.
The proof is inspired by the proof of Theorem 1 in [Che23], which analyzes MLE via the trajectory of the preconditioned gradient descent (PGD) dynamic starting from ground truth. More precisely, letting $\bm{\theta}^{0}=\bm{\theta}^{\star}$, we consider the PGD iterates defined by
$\bm{\theta}^{t+1}=\bm{\theta}^{t}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{t}),$
where $\eta>0$ is the step size of PGD. [Che23] shows that this dynamic converges to $\widehat{\bm{\theta}}$. We proceed one step further by establishing precise distributional characterization of $\hat{\bm{\theta}}$ via analyzing PGD. With Taylor expansion, the gradient can be decomposed into
$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
where $\bm{r}^{t}$ is a residual vector with small magnitude. Then the PGD update becomes
$\bm{\theta}^{t+1}-\bm{\theta}^{\star}=\left(1-\eta\right)(\bm{\theta}^{t}-\bm{\theta}^{\star})-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right).$
We establish Lemma 4 by solving this recursive relation. More specifically, as $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ does not depend on $t$ and $\|\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\|_{\infty}$ can be controlled for each step $t$, taking $t\rightarrow\infty$, we see that
$\widehat{\bm{\theta}}-\bm{\theta}^{\star}=\lim_{t\rightarrow\infty}\bm{\theta}^{t}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
for some residual term $\bm{r}$ that is well controlled in $\ell_{\infty}$ norm.
5 Experiment
In this section, we demonstrate the empirical performance of $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$ using both simulated and real data.
5.1 Simulations
We begin by using simulated data to validate our theoretical results and compare our estimators with existing ones for the Rasch model.
The data generating process follows the model specified in Section 2.1. Unless specified otherwise, in each trial, the user and item parameters are randomly drawn from
$\widetilde{\bm{\zeta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{n}),\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{m}).$
Afterwards, $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ is computed by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean.
5.1.1 $\ell_{\infty}$ estimation error
Here, we investigate the $\ell_{\infty}$ estimation error. First, we validate the theoretical result in $\ell_{\infty}$ estimation error of $\mathsf{RP\text{-}MLE}$. Second, we show how much advantage the $\mathsf{MRP\text{-}MLE}$ brings through multiple runs of data splitting. Lastly, we compare our methods with existing comparison-based algorithms, including the case where $\kappa_{1},\kappa_{2}$ are large.
Validating the theoretical result.
Theorem 1 tells us that the $\ell_{\infty}$ error scales as $1/\sqrt{np}$. As shown in Figure 1, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ is very close to being linear w.r.t. both $1/\sqrt{n}$ and $1/\sqrt{p}$. This is consistent with our theoretical predictions.
(a) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ with v.s. $1/\sqrt{n}$. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $10000$ to $40000$. Each point represents the average of 1000 trials.
(b) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{p}$. The parameter is chosen to be $m=50,n=10000$ and $p$ varies from $1/9$ to $1$. Each point represents the average of 1000 trials.
Figure 1: Estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ of $\mathsf{RP\text{-}MLE}$ with varying $n$ and $p$.
Multiple runs in $\mathsf{MRP\text{-}MLE}$.
As we have discussed after Theorem 1, the random data splitting could incur a small loss of information. We have provided a remedy $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by averaging over multiple runs with independent data splitting. In Figure 2, we show the empirical evidence of its effectiveness. For each trial, we run 50 independent data splitting and obtain the 50 estimates $\widehat{\bm{\theta}}^{(i)}$. We then report the $\ell_{\infty}$ error of the running average, i.e., $\|\frac{1}{k}\sum_{i=1}^{k}\widehat{\bm{\theta}}^{(i)}-\bm{\theta}^{\star}\|$ for $k=1,\ldots,50$. As a baseline for comparison, we also run pseudo MLE on the same datasets, which uses all possible pairs of comparison. We can see that by averaging over more runs of data splittings, $\mathsf{MRP\text{-}MLE}$ gives a better estimation. In particular, one can see that with sufficient number of independent data splittings, the average estimation error matches and even beats pseudo MLE.
Figure 2: The $\ell_{\infty}$ estimation error of $\mathsf{MRP\text{-}MLE}$ with different numbers of data splitting. The dashed line is the performance of Pseudo MLE. The parameters are chosen to be $m=50,p=0.1,n=10000$. Each point is averaged over 1000 trials.
Comparison with existing estimator.
We then compare our algorithms with two other comparison-based algorithms: pseudo MLE (PMLE) and the spectral method in [NZ23]. We plot the $\ell_{\infty}$ error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ against $n$ in Figure 4. We can see that all four algorithms follow a similar trend. The performance of $\mathsf{MRP\text{-}MLE}$ is on par with PMLE and better than the spectral method. Our proposed algorithm not only enjoys an improved theoretical guarantee but is also competitive in practice.
Performance with large $\kappa_{1},\kappa_{2}$.
While we assume $\kappa=\max\{\kappa_{1},\kappa_{2}\}=O(1)$ is most of this article, the scenario with large $\kappa$ can be practically relevant. Here we compare the $\ell_{\infty}$ error of different methods when the condition number is large. Fixing $\kappa$, we draw the user and item parameters with
$\widetilde{\bm{\zeta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))$
and compute $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean. Figure 4 shows the performance of different estimators as $\kappa$ varies. The MLE-based approaches including $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ achieve better $\ell_{\infty}$ error than the spectral method when $\kappa$ is large.
Figure 3: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $n$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $5000$ to $20000$. The result is averaged over 1000 trials.
Figure 4: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $\log(\kappa)$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1,n=20000$ and $\kappa$ varies from $1$ to $e^{10}$. The result is averaged over 1000 trials.
5.1.2 Top-$K$ recovery
In this part, we investigate the performance of different algorithms in top-$K$ recovery. We set $\theta_{i}^{\star}=(1-K/m)\Delta_{K}$ for $i\leq K$ and $\theta_{i}^{\star}=(-K/m)\Delta_{K}$ otherwise. For any estimator $\widehat{\bm{\theta}}$, we define top-$K$ recovery rate to be
$\frac{1}{K}\left|\left\{i\leq K:i\in A_{K}\right\}\right|,$
where $A_{K}$ is an arbitrary $K$-element set such that $\widehat{\theta}_{i}\geq\widehat{\theta}_{j}$ for any $i\in A_{K},j\notin A_{K}$. We compare the top-$K$ recovery rate of PMLE and the spectral method in [NZ23] with $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ in Figure 6. The recovery rate of PMLE, spectral method and $\mathsf{MRP\text{-}MLE}$ is similar, indicating again that our algorithm performs well in practice.
Figure 5: Top-$K$ recovery rate using spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=10000,m=50,p=0.1,K=5$ and $\Delta_{K}$ varies from 0.1 to 0.7. The result is averaged over 1000 trials.
Figure 6: The relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ v.s. $n$ for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$. The parameter is chosen to be $p=0.1$, $n$ varies from $10000$ to $50000$ and $m=n/500$. The result is averaged over 1000 trials.
5.1.3 Refined $\ell_{2}$ estimation error
In this part of the section, we show the empirical evidence of Theorem 2. In Theorem 2 we have shown that the $\ell_{2}$ error concentrate around $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ for $\widetilde{z}\in\left\{\widehat{z},z\right\}$. In each trial we compute the following quantity
$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}$
for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$. This measures the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$. In Figure 6 we consider the regime where $p$ and $n/m$ is fixed. In this case, Theorem 2 implies that
$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}\lesssim\frac{\sqrt{\frac{1}{np}}+\frac{\sqrt{m}}{np}}{\sqrt{\frac{m}{np}}}\lesssim\frac{1}{\sqrt{n}}.$
In other words, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrate tightly around $\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})$. We can see that the deviation is very small between $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$. In both cases, the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ decreases as $n$ and $m$ increase as expected.
5.1.4 Inference
Lastly, we validate our inferential results in Section 3.3.
Asymptotic normality.
We first check the asymptotic normality presented in Theorem 3 and Corollary 1. For simplicity, we focus on the first coordinate. Theorem 3 and Corollary 1 have claimed that
$\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1-\frac{1}{m}\right)$
and
$\frac{\widehat{\theta}_{1}-\theta_{1}^{\star}}{\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$
For simplicity we set both $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ to be all zero vectors. In Figure 7, we present the QQ-plot comparing the generated instances and the theoretical distributions. In both cases, the spread of the term of interest matches the normality claimed in Theorem 3 and Corollary 1.
Figure 7: Quantile-quantile plots of $\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}$ and $(\widehat{\theta}_{1}-\theta_{1}^{\star})/\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}$ v.s. zero-mean normal distributions. The estimate $\widehat{\bm{\theta}}$ is computed with $\mathsf{RP\text{-}MLE}$. The parameters are set to be $n=10000,m=10,p=0.5$. Data is collected over 10000 trials.
Coverage of the confidence intervals.
We compare the empirical coverage rate of the two-sided confidence intervals in Proposition 3 with the theoretical ones. Without loss of generality we consider the confidence interval for the first coordinate. In each trial, we compute the empirical coverage rate
$1-\widehat{\alpha}=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\{\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\}$
and report the average $1-\widehat{\alpha}$ over the trials. We do this for multiple $\alpha$’s and compare $1-\alpha$ with $1-\widehat{\alpha}$ in Table 1. The empirical coverage rate $1-\widehat{\alpha}$ is very close to the theoretical coverage rate $1-\alpha$.
$1-\alpha$ 0.8 0.9 0.95 0.98 0.99 0.995 0.999
$1-\widehat{\alpha}$ 0.801145 0.900635 0.95108 0.980395 0.99025 0.99523 0.99904
Table 1: Theoretical coverage rate $1-\alpha$ and empirical coverage rate $1-\widehat{\alpha}$ of the two-sided confidence intervals $[\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$. The parameters are set to be $n=10000,m=20,p=0.1$. The result is averaged over $10000$ trials.
5.2 LSAT real data
We study a real dataset (LSAT) on the Law School Admissions Test from [DBL70]. LSAT has full observation of 1000 people answering 5 problems. Foro each person-item pair it records whether the answer is correct or not. The second row in Table 2 lists how many people answer each problem correctly. From the first look, Problem 3 seems to be the hardest question. We proceed to quantify the hardness of these problems under the Rasch model and infer how confident we are in claiming it is the hardest. We use $\mathsf{RP\text{-}MLE}$ to get an estimate of the latent scores and calculate two-sided confidence intervals at level $\alpha=0.01$ for each coordinate with the framework in Section 3.3. See Table 2 for the result. Higher latent score here means higher difficulty. The estimated parameters agree with the total number of correct answers in reverse order. We can also see that the confidence interval lower bound of $\theta_{3}^{\star}$ is larger than the largest confidence interval upper bound of $\theta_{1}^{\star},\theta_{2}^{\star},\theta_{4}^{\star},\theta_{5}^{\star}$. With Bonferroni correction, we have a 0.95 confidence level to say Problem 3 is the most difficult problem.
Problem 1 2 3 4 5
Total correct 924 709 553 763 870
$\theta$ estimate -1.2139 0.4093 1.1123 0.2977 -0.6053
CI lower bound -1.5957 0.1449 0.8321 0.0234 -0.9050
CI upper bound -0.8322 0.6736 1.3926 0.5720 -0.3057
Table 2: Latent score estimate calculated using $\mathsf{RP\text{-}MLE}$ and confidence interval calculated with the framework in Section 3.3. The significance level is chosen to be $\alpha=0.01$ for each coordinate.
Now we assume Problem 3 is the top-$1$ item and investigate the top-$1$ recovery rate of different algorithms on LSAT under incomplete observation. In each trial we randomly select $\widetilde{n}$ people, and for each of them we randomly select their outcome on $\widetilde{m}$ problems. We then estimate $\bm{\theta}^{\star}$ using the subsampled data with different methods and compare their top-1 recovery rate, i.e., the proportion of trials where the top-$1$ item is correctly identified. In Figure 8 we see results similar to the simulation. Our algorithm $\mathsf{MRP\text{-}MLE}$ has a similar recovery rate compared to PMLE and spectral method.
Figure 8: Top-1 recovery rate using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameters are chosen to be $\widetilde{m}=4$ and $\widetilde{n}$ varies from 100 to 200. The result is averaged over 10000 trials.
6 Discussion
In this paper, we propose two new likelihood-based estimators $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ for item parameter estimation in the Rasch model. Both enjoy optimal finite sample estimation guarantee and $\mathsf{RP\text{-}MLE}$ is further amenable to tight uncertainty quantification. All this is achieved even when the user-item response data are extremely sparse (cf. [NZ23]). Below, we identify several questions that are interesting for further investigation:
  • Does PMLE achieve optimal theoretical guarantee? In our experiments, pseudo MLE has shown a similar performance to $\mathsf{MRP\text{-}MLE}$. This naturally leads to the question of whether PMLE can enjoy the same theoretical guarantee. This is relevant to our work because our methods can be viewed as a modification of pseudo MLE by incorporating random disjoint pairing to decouple statistical dependency among paired $Y_{ij}$’s. It remains unclear whether such dependency is a fundamental bottleneck.
  • Inference for multiple data splitting. In $\mathsf{MRP\text{-}MLE}$ (Algorithm 2), we average the estimates over multiple runs of data splitting to improve the empirical result over $\mathsf{RP\text{-}MLE}$. While this is effective for estimation, we do not have an immediate analog for inference as the estimations $\{\widehat{\bm{\theta}}^{(i)}\}_{i\in[n_{\mathrm{split}}]}$ are dependent on each other. It is an interesting question whether we can characterize the distribution of $\widehat{\bm{\theta}}$ in $\mathsf{MRP\text{-}MLE}$ and whether it leads to better inference, e.g., tighter confidence intervals.
  • Extending random pairing to other models in IRT. Some IRT models parameterize the latent score of users and items differently from the Rasch model. For instance, the two-parameter logistic model (2PL) assumes that $X_{ti}$, the response of user $t$ to item $i$, follows the law
    $\mathbb{P}[X_{ti}=1]=\frac{\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})}{1+\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})},$
    where $\bm{\theta}^{\star}$ is the latent score of the items while $\bm{a}^{\star}$ and $\bm{b}^{\star}$ are two parameter vectors for the users. Unlike the Rasch model, in the 2PL model,
    $\mathbb{P}[X_{ti}>X_{tj}\mid X_{ti}\neq X_{tj}]=\frac{\exp(b_{t}^{\star}\theta_{i}^{\star})}{\exp(b_{t}^{\star}\theta_{i}^{\star})+\exp(b_{t}^{\star}\theta_{j}^{\star})}$
    is not independent of the user parameter $\bm{b}^{\star}$. Therefore reduction to the BTL model is no longer true in this case. It is interesting and non-trivial to extend the idea of random pairing to the 2PL model.
Acknowledgement
YY and CM were partially supported by the National Science Foundation via grant DMS-2311127.
Appendix A Proof of Fact 1
Expanding the probability of the random events, we have that
$\mathbb{P}[X_{ti}<X_{tj}\mid X_{ti}\neq X_{tj}]$ $=\frac{\mathbb{P}[X_{ti}=0,X_{tj}=1]}{\mathbb{P}[X_{ti}=0,X_{tj}=1\text{\text{ or }}X_{ti}=1,X_{tj}=0]}$
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\Bigg{/}\left(\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\right)$
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{e^{\zeta_{t}^{\star}}(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}$
$=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}.$
Now consider $\mathbb{P}[X_{ti}\neq X_{tj}]$, we have
$\mathbb{P}[X_{ti}\neq X_{tj}]$ $=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}$
$=\frac{e^{\theta_{j}^{\star}-\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}}}{(1+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}})(1+e^{\theta_{j}^{\star}-\zeta_{t}^{\star}})}.$ (16)
Let $f:[1/\kappa_{2},\kappa_{2}]^{2}\rightarrow\mathbb{R}$ defined by
$f(a,b)\coloneqq\frac{a+b}{(1+a)(1+b)}.$
Its partial derivatives are
$\frac{\partial}{\partial a}f(a,b)=\frac{b^{2}-1}{(1+a)^{2}(1+b)^{2}}\quad\text{and}\quad\frac{\partial}{\partial b}f(a,b)=\frac{a^{2}-1}{(1+a)^{2}(1+b)^{2}}.$
It is now easy to see that the minimum or maximum of $f$ can only happen if $(a,b)=(1,1)$ or $(a,b)\in\{1/\kappa_{2},\kappa_{2}\}^{2}$. After comparing the value of $f$ at these points, we conclude that $f$ achieves minimum at
$f(1/\kappa_{2},1/\kappa_{2})=f(\kappa_{2},\kappa_{2})=\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}.$
By the definition of $\kappa_{2}$, $|\theta_{l}^{\star}-\zeta_{t}^{\star}|\leq\log(\kappa_{2})$ for any $l\in[m]$. Then (16) fits the definition of $f$ and the proof is completed.
Appendix B Degree and spectral properties of the comparison graphs
In this section, we present the analysis for lemmas that characterize the degree and spectral properties of the comparison graphs. We start with a lemma that controls the degrees in $\mathcal{G}_{X}$, and then prove Lemmas 2 and 3.
B.1 Degree range of $\mathcal{G}_{X}$
Recall that $n_{t}$ is the number of neighbors of user $t$ in $\mathcal{G}_{X}$. Furthermore, we denote $m_{i}$ as the number of users that is compared with problem $i$ and at least another item, i.e.,
$m_{i}\coloneqq\left|\{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\}\right|.$ (17)
The following lemma controls the size of $n_{t}$ and $m_{i}$.
Lemma 5 (Degree bounds in $\mathcal{G}_{X}$).
Suppose that $np\geq C\log(n)$ for some large enough constant $C>0$ and that $m\leq n^{\alpha}$ for some constant $\alpha>0$. Then with probability at least $1-2n^{-10}$, for all $i\in[m]$, we have
$\frac{1}{4}np\leq m_{i}\leq\frac{3}{2}np.$ (18)
Moreover, with probability at least $1-n^{-10}$, for all $t\in[n]$, we have
$n_{t}\leq\left(\frac{3}{2}mp\right)\vee 165\log(n).$ (19)
Proof.
We prove the two claims in the lemma sequentially.
Fix any $t,i$. One has
$\mathbb{P}\left[t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\right]$ $=\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]-\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\text{ and }n_{t}=1\right]$
$=p-p(1-p)^{m-1}$
$\geq p(1-e^{-(m-1)p})$
$\geq\frac{1}{2}p,$ (20)
as long as $mp\geq 2$. Let $\mu_{i}\coloneqq\mathbb{E}[m_{i}]$. By the linearity of expectation, we have
$np/2\leq\mu_{i}\leq\sum_{t}\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]=np.$ (21)
Fix $i\in[m]$. Since the sampling is independent with different $t$, by the Chernoff bound,
$\mathbb{P}[|m_{i}-\mu_{i}|\leq(1/2)\mu_{i}]$ $\leq 2e^{-\frac{1}{12}\mu_{i}}\leq 2e^{-\frac{1}{24}np}\leq m^{-1}n^{-10}$
as long as $np\geq C\log(n)$ for large enough constant $C$. Applying (21) and union bound on $i\in[m]$ yields (18).
Moving on to (19), we first consider the case where $mp\geq 110\log(n)$. By Chernoff bound,
$\mathbb{P}[n_{t}\geq(3/2)mp]\leq e^{-\frac{1}{10}mp}\leq n^{-11}.$
In the case of $mp<110\log(n)$, the quantity $\mathbb{P}\left[n_{t}\geq 165\log(n)\right]$ clear decreases as $p$ decreases. So we may use the case $mp=110\log(n)$ to bound this quantity and conclude that
$\mathbb{P}\left[n_{t}\geq 165\log(n)\right]\leq n^{-11}.$
Finally we apply union bound on $t\in[n]$ to reach (19).
B.2 Proof of Lemma 2
The assumption of this lemma allows us to invoke Lemma 5. For the upper bound of $d_{i}$, since (18) is true,
$d_{i}$ $=\sum_{j:j\neq i}L_{ij}$
$=\sum_{t:(t,i)\in\mathcal{E}_{X}}\sum_{j:j\neq i}L_{ij}^{t}$
=$\overset{(\text{i})}{=}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}L_{ij}^{t}$
$\overset{(\text{ii})}{\leq}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}R_{ij}^{t}\leq m_{i}\leq\frac{3}{2}np.$
Here (i) holds since $L_{ij}^{t}$ can only be 0 when $n_{t}\leq 1$, and (ii) holds since $L_{ij}^{t}\leq R_{ij}^{t}$ by definition. For the lower bound of $d_{i}$, notice that for any $(t,i)$, $\sum_{j}L_{ij}^{t}$ is either 00 or 1. Fix $\mathcal{E}_{X}$ and only consider randomness on $L_{ij}^{t}$. By Hoeffding’s inequality,
$\mathbb{P}\left\{d_{i}-\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j:j\neq i}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\leq-\frac{1}{12\kappa_{2}}np\right\}$ $\leq\exp\left(-\frac{(1/72)\kappa_{2}^{-2}n^{2}p^{2}}{m_{i}}\right)$
$\leq\exp\left(-\frac{\kappa_{2}^{-2}np}{108}\right)$
$\leq m^{-1}n^{-10}$ (22)
as long as $np\geq 1200\kappa_{2}^{2}\log(n)$. The second to last inequality uses (18). For each $(t,i)\in\mathcal{E}_{X}$,
$\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]$ $=\sum_{j}\mathbb{P}\left[R_{ij}^{t}=1\mid(t,i)\in\mathcal{E}_{X}\right]\mathbb{P}\left[\sum_{j}L_{ij}^{t}=1\mid R_{ij}^{t}=1\right]$
$\geq\sum_{j}\mathbb{P}\left[R_{ij}^{t}=1\mid(t,i)\in\mathcal{E}_{X}\right]\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}$
$\geq\frac{2\lfloor n_{t}/2\rfloor}{n_{t}}\cdot\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq\frac{1}{3\kappa_{2}}$
as long as $n_{t}\geq 2$. The first inequality here uses Fact 1. Then by definition of $m_{i}$ in (17),
$\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\geq\frac{1}{3\kappa_{2}}m_{i}.$ (23)
Combining (22), (23) and (18),
$d_{i}\geq\frac{1}{3\kappa_{2}}m_{i}-\frac{1}{12\kappa_{2}}np\geq\frac{1}{24\kappa_{2}}np.$
Applying union bound over $i\in[m]$ yields the desired result.
B.3 Proof of Lemma 3
We first consider the spectrum of $\bm{L}_{L}$. Recall that
$\bm{L}_{L}$ $=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
$=\sum_{t=1}^{n}\underbrace{\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}^{t}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}}_{\bm{L}_{L}^{t}}.$
For the upper bound, it is clear from Lemmas 10 and 2 that $\lambda_{1}(\bm{L}_{L})\leq 2\max_{i}d_{i}\leq 3np$. For the lower bound, we use the matrix Chernoff inequality (see Section 5 of [Tro15]). Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then $\lambda_{m-1}(\bm{L}_{L})=\lambda_{m-1}(\bm{R}\bm{L}_{L}\bm{R}^{\top})$. For any $t\in[n]$, by (19),
$0\leq\lambda_{m-1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})\leq\lambda_{1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})=\lambda_{1}(\bm{L}_{L}^{t})\leq 2$
The last inequality follows from Lemma 10 since $\sum_{j}\bm{L}_{ij}^{t}\leq 1$ for any $t\in[n]$ and $i\in[m]$. By Fact 1,
$\mathbb{P}[L_{ij}^{t}=1\mid R_{ij}^{t}=1]\geq\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq 1/(2\kappa_{2}).$
Then
$\lambda_{m-1}(\mathbb{E}\bm{R}\bm{L}_{L}\bm{R}^{\top})$ $=\lambda_{m-1}\left(\bm{R}\sum_{t=1}^{n}\sum_{i>j}\mathbb{E}L_{ij}^{t}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{R}^{\top}\right)$
$\geq\frac{1}{2\kappa_{2}}\lambda_{m-1}\left(\bm{R}\sum_{t=1}^{n}\sum_{i>j}\mathbb{E}R_{ij}^{t}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{R}^{\top}\right)$
Moreover $\sum_{i>j}\mathbb{E}R_{ij}^{t}\geq\frac{1}{2}(\mathbb{E}[n_{t}]-1)=(mp-1)/2$, where the $-1$ accounts for possible unpaired $X_{ti}$. By symmetry $\mathbb{E}R_{ij}^{t}$ is the same for any $(i,j)$. Then for any $(i,j)$,
$\mathbb{E}R_{ij}^{t}\geq\frac{mp-1}{2}/\binom{m}{2}\geq\frac{p}{2m}$
as long as $mp\geq 2$. Thus
$\lambda_{m-1}(\mathbb{E}\bm{R}\bm{L}_{L}\bm{R}^{\top})$ $\geq\frac{1}{2\kappa_{2}}\lambda_{m-1}\left(\bm{R}\sum_{t=1}^{n}\sum_{i>j}\frac{p}{2m}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{R}^{\top}\right)$
$=\frac{1}{2\kappa_{2}}\cdot n\cdot\frac{p}{2m}\cdot m$
$=\frac{np}{4\kappa_{2}}.$
Now invoke the matrix Chernoff inequality, we have
$\mathbb{P}\left\{[\lambda_{m-1}(\bm{R}\bm{L}_{L}\bm{R}^{\top})]\leq\frac{np}{8\kappa_{2}}\right\}$ $\leq m\cdot\left[\frac{e^{-1/2}}{(1/2)^{1/2}}\right]^{\frac{np}{4\kappa_{2}}/2}\leq n^{-10}$
as long as $np\geq C\kappa_{2}\log(n)$ for some large enough constant $C$.
The spectrum of $\bm{L}_{Lz}$ comes directly from the spectrum of $\bm{L}_{L}$. Recall
$\bm{L}_{Lz}=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
By Lemma 9,
$\lambda_{1}(\bm{L}_{Lz})$ $=\max_{\bm{v}:\|\bm{v}\|=1}\bm{v}^{\top}\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$\leq\frac{1}{4}\max_{\bm{v}:\|\bm{v}\|=1}\bm{v}^{\top}\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$=\frac{1}{4}\lambda_{1}(\bm{L}_{L})\leq np,$
and
$\lambda_{m-1}(\bm{L}_{Lz})$ $=\min_{\bm{v}:\|\bm{v}\|=1,\bm{v}^{\top}\bm{1}_{m}=0}\bm{v}^{\top}\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$\geq\frac{1}{4\kappa_{1}}\min_{\bm{v}:\|\bm{v}\|=1,\bm{v}^{\top}\bm{1}_{m}=0}\bm{v}^{\top}\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$=\frac{1}{4\kappa_{1}}\lambda_{m-1}(\bm{L}_{L})\geq\frac{np}{16\kappa_{1}\kappa_{2}}.$
Appendix C Proof of Lemma 4
We study the MLE $\widehat{\bm{\theta}}$ by analyzing the iterates of preconditioned gradient descent starting from the ground truth. Let $\bm{\theta}_{0}=\bm{\theta}^{\star}$ be the starting point and $\eta>0$ be the stepsize that is small enough. The preconditioned gradient descent iterates are given by
$\bm{\theta}^{t+1}=\bm{\theta}^{t}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{t}).$ (24)
Consider the Taylor expansion of $\nabla\mathcal{L}(\bm{\theta}^{t})$, we have
$\nabla\mathcal{L}(\bm{\theta}^{t})$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-Y_{ji}^{(l)}+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}+\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})(\delta_{i}^{t}-\delta_{j}^{t})+\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}\right)(\bm{e}_{i}-\bm{e}_{j})\right).$
Here $\bm{\delta}^{t}\coloneqq\bm{\theta}^{t}-\bm{\theta}^{\star}$ and for all $(i,j)$, $\xi_{ij}^{t}\in\mathbb{R}$ is some number that lies between $\theta_{i}^{\star}-\theta_{j}^{\star}$ and $\theta_{i}^{t}-\theta_{j}^{t}$. As $\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=z_{ij}$ and $\delta_{i}^{t}-\delta_{j}^{t}=(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{\delta}^{t}$, we have
$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}\bm{\delta}^{t}-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
where $\bm{r}^{t}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}[\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}(\bm{e}_{i}-\bm{e}_{j})]$. Feeding this into (24), we have
$\bm{\delta}^{t+1}=\left(1-\eta\right)\bm{\delta}^{t}-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right)$ (25)
By definition $\bm{\delta}^{0}=\bm{0}$. Applying this recursive relation $t-1$ times we obtain
$\bm{\delta}^{t}$ $=-\eta\sum_{i=0}^{t-1}(1-\eta)^{i}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}$
$=-\left[1-(1-\eta)^{t}\right]\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}.$
At this point, we invoke an existing result on these terms that have been studied in [YCOM24] for a more general setting. The proof of Lemma 2 combined with Lemmas 3 and 4 in [YCOM24] reveal the following properties of (24).
Lemma 6.
Instate the assumptions of Theorem 1. Suppose that
$\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{Lz}))^{3}}\leq C_{1}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{Lz})}},$ (26)
for some constant $C_{1}>0$. Then with probability at least $1-n^{-10}$, the precondition gradient descent dynamic satisfies the following properties:
  1. 1.
    There exists a unique minimizer $\widehat{\bm{\theta}}$ of (2).
  2. 2.
    There exist some $\alpha_{1},\alpha_{2}$ obeying $0<\alpha_{1}\leq\alpha_{2}$ such that any $t\in\mathbb{N}$,
    $\|\bm{\theta}^{t}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}}\leq(1-\eta\alpha_{1})^{t}\|\bm{\theta}^{0}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}},$
    provided that $0<\eta\leq 1/\alpha_{2}$.
  3. 3.
    For any $k,l$ and iteration $t\geq 0$,
    $\left|\left(\theta_{k}^{t}-\theta_{l}^{t}\right)-\left(\theta_{k}^{\star}-\theta_{l}^{\star}\right)\right|\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$ (27)
    for some constant $C_{2}>0$.
  4. 4.
    For any $k,l$ and iteration $t\geq 0$,
    $\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{3}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{L}))^{3}}$
    for some constant $C_{3}>0$.
Lemmas 3 and 2 imply that
$\frac{np}{4\kappa_{1}\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{Lz})\qquad\text{and}\qquad d_{\max}\leq\frac{3}{2}np.$
Then the condition (26) holds as long as $np\geq C_{4}\kappa_{1}^{2}\kappa_{2}\log^{3}(n)$ for some large enough constant $C_{4}$. Invoke Lemma 6 to see that for any $(k,l,t)$,
$\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np},$ (28)
where $C_{5}>0$ is some constant. Furthermore, the convergence given by Lemma 6 implies that
$\widehat{\bm{\delta}}$ $=\lim_{t\rightarrow\infty}\bm{\delta}^{t}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\underbrace{\eta\lim_{t\rightarrow\infty}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}}_{\eqqcolon\bm{r}}.$ (29)
[…]
$|r_{k}-r_{j}|\leq\eta\lim_{t\rightarrow 0}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}=C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
As $\bm{1}^{\top}\bm{L}_{Lz}^{\dagger}=0$, the above inequality implies that
$|r_{k}|$ $=\left|\frac{1}{m}\cdot m\bm{e}_{k}^{\top}\bm{r}\right|=\left|\frac{1}{m}\sum_{l=1}^{m}(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{r}\right|$
$=\left|\frac{1}{m}\sum_{l=1}^{m}(r_{k}-r_{l})\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
The proof is now completed.
Appendix D Proof of Theorem 2
We start with the proof of (5). By Lemma 4 we can express the MLE estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ as
$\widehat{\bm{\theta}}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$ (30)
for $\bm{B},\widehat{\bm{\epsilon}}$ defined in Section 4.3 and $\bm{r}\in\mathbb{R}^{m}$ is a residual term obeying $\|\bm{r}\|_{\infty}\leq C_{1}\kappa_{1}^{6}\log^{2}n/(np)$ for some constant $C_{1}$.
We first focus on the main term $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$. Expanding $\bm{B}$ and $\widehat{\bm{\epsilon}}$, we rewrite it as
$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\epsilon_{ij}^{(l)}\bm{L}_{Lz}^{\dagger}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
It is easy to see that conditional on $\mathcal{G}_{Y}$, $\{\bm{u}_{ij}^{(l)}\}_{i,j,l}$ is a set of independent zero-mean random variables. Thus we can expand $\mathbb{E}[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}]$ to be
$\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]$ $=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\bm{u}_{ij}^{(l)\top}\bm{u}_{ij}^{(l)}$
$=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathrm{Trace}\left(\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right)$
[…]
Here (i) follows from the equality
$z_{ij}=\mathrm{Var}(\epsilon_{ji}^{(l)})=\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}},$
and (ii) follows from the definition of $\bm{L}_{Lz}$. By Lemma 3,
$\frac{m}{2np}\leq(m-1)\lambda_{m-1}(\bm{L}_{Lz}^{\dagger})\leq\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})\leq m\|\bm{L}_{Lz}^{\dagger}\|\leq\frac{16\kappa_{1}\kappa_{2}m}{np}.$ (32)
Moreover, $\{\epsilon_{ij}^{(l)}\}_{i,j,l}$ is a set of sub-Gaussian random variable with variance proxy $1/z_{ij}$ (see, e.g., Section 2.5 in [Ver18]) and $1/z_{ij}\leq 4\kappa_{1}$ by Lemma 9. Applying Hanson-Wright inequality (see [RV13]), for any scalar $a>0$ we have
$\mathbb{P}\left[\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}-\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]\right|>a\right]$ (33)
$\quad\leq 2\exp\left[-C_{2}\left(\frac{a^{2}}{\left(4\kappa_{1}\right)^{4}\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|_{\mathrm{F}}^{2}}\wedge\frac{a}{\left(4\kappa_{1}\right)^{2}\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|}\right)\right]$
for some constant $C_{2}>0$. For $\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|_{\mathrm{F}}$ and $\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|$ we have
$\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|$ $\leq\|\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\|_{\mathrm{F}}$
$=\sqrt{\mathrm{Trace}\left(\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\right)}$
$=\sqrt{\mathrm{Trace}\left(\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\bm{B}^{\top}\right)}$
[…]
Here (i) follows from the fact that $\bm{B}\bm{B}^{\top}=\bm{L}_{Lz}$, (ii) follows from Lemma 3 and the fact that $\mathrm{Trace}(\bm{M})\leq m\|\bm{M}\|$ for any $m\times m$ matrix $\bm{M}$, and (iii) follows from Lemma 3. Now substitute $a=C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{m}\log(n)/(np)$ in (33) for some large enough constant $C_{3}$. We have that with probability at least $1-2n^{-10}$,
$\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}-\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]\right|\leq\frac{C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{m}\log(n)}{np}.$ (34)
Combining this with (31) and (32),
$\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $=\frac{\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}-\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})\right|}{\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|+\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}}$
$\leq\frac{C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{m}\log(n)/(np)}{\sqrt{m/(2np)}}$
$\leq C_{4}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}$
for some constant $C_{4}>0$.
Substituting (34) and (31) into (30), we have that for some constant $C_{1},C_{2}$,
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|+\|\bm{r}\|$
$\leq C_{4}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{1}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np}.$ (35)
The proof of (5) is now completed. For (6), the following lemma connects $\widehat{z}$ and $z$. The proof is deferred to the end of this section.
Lemma 7.
Instate the assumptions of Theorem 1, then with probability at least $1-C_{5}n^{-10}$ for some constant $C_{5}>0$,
$\|\bm{L}_{Lz}^{\dagger}-\bm{L}_{L\hat{z}}^{\dagger}\|\leq\frac{C_{6}\kappa_{1}^{7/2}\kappa_{2}^{2}}{(np)^{3/2}}$
for some large enough constant $C_{6}$.
Combining this lemma with Weryl’s inequality, we have that
$\left|\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})-\mathrm{Trace}(\bm{L}_{L\hat{z}}^{\dagger})\right|\leq\frac{C_{6}m\kappa_{1}^{7/2}\kappa_{2}^{2}}{(np)^{3/2}}.$
Then by (32),
$\left|\sqrt{\mathrm{Trace}(\bm{L}_{L\hat{z}}^{\dagger})}-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $=\frac{\left|\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})-\mathrm{Trace}(\bm{L}_{L\hat{z}}^{\dagger})\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}+\sqrt{\mathrm{Trace}(\bm{L}_{L\hat{z}}^{\dagger})}}$
$=\frac{C_{6}m\kappa_{1}^{7/2}\kappa_{2}^{2}/(np)^{3/2}}{\sqrt{m/(2np)}}\leq\frac{\sqrt{2}C_{6}\kappa_{1}^{7/2}\kappa_{2}^{2}\sqrt{m}}{np}.$
Using triangular inequality, we conclude that
$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\hat{z}}^{\dagger})}\right|\leq C_{4}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{(C_{1}\kappa_{1}^{6}+\sqrt{2}C_{6}\kappa_{1}^{7/2}\kappa_{2}^{2})\sqrt{m}\log^{2}(n)}{np}.$
Proof of Lemma 7.
Recall $\sigma$ is the sigmoid function and its derivative $\sigma^{\prime}$ is $1$-Lipschitz. By Theorem 1, for all $(i,j)$,
$\left|z_{ij}-\widehat{z}_{ij}\right|$ $=\left|\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})-\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})\right|$
$\leq\left|\left(\widehat{\theta}_{i}-\widehat{\theta}_{j}\right)-\left(\theta_{i}^{\star}-\theta_{j}^{\star}\right)\right|\leq C_{7}\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{\log(n)}{np}}$
for some constant $C_{7}>0$. Then
$\|\bm{L}_{Lz}-\bm{L}_{L\hat{z}}\|$ $=\max_{\bm{v}\in\mathbb{R}^{m}:\|\bm{v}\|=1}\left|\bm{v}^{\top}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\left(z_{ij}-\widehat{z}_{ij}\right)(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}\right|$
$\leq\max_{\bm{v}\in\mathbb{R}^{m}:\|\bm{v}\|=1}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\left|z_{ij}-\widehat{z}_{ij}\right|\bm{v}^{\top}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$\leq C_{7}\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{\log(n)}{np}}\cdot\|\bm{L}_{L}\|$
$\leq 3C_{7}\kappa_{1}\kappa_{2}^{1/2}\sqrt{np\log(n)},$
where the last line follows from Lemma 3. As $np\geq C_{8}\kappa_{1}^{4}\kappa_{2}^{3}\log^{2}(n)$ for some large enough constant $C_{8}$, $\|\bm{L}_{Lz}-\bm{L}_{L\hat{z}}\|\leq np/(32\kappa_{1}\kappa_{2})$. By Weryl’s inequality and Lemma 3,
$\lambda_{m-1}(\bm{L}_{L\hat{z}})\geq\lambda_{m-1}(\bm{L}_{Lz})-\|\bm{L}_{Lz}-\bm{L}_{L\hat{z}}\|\geq\frac{np}{16\kappa_{1}\kappa_{2}}-\frac{np}{32\kappa_{1}\kappa_{2}}=\frac{np}{32\kappa_{1}\kappa_{2}}.$
This implies that $\|\bm{L}_{Lz}^{\dagger}\|\leq 16\kappa_{1}\kappa_{2}/(np)$ and $\|\bm{L}_{L\hat{z}}^{\dagger}\|\leq 32\kappa_{1}\kappa_{2}/(np)$. Using the perturbation bound of pseudo-inverse (see Theorem 4.1 in [Wed73]), we have
$\|\bm{L}_{Lz}^{\dagger}-\bm{L}_{L\hat{z}}^{\dagger}\|$ $\leq 3\cdot\|\bm{L}_{Lz}^{\dagger}\|\cdot\|\bm{L}_{L\hat{z}}^{\dagger}\|\cdot\|\bm{L}_{Lz}-\bm{L}_{L\hat{z}}\|$
$\leq\frac{C_{9}\kappa_{1}^{3}\kappa_{2}^{5/2}}{(np)^{3/2}}$
for some constant $C_{9}>0$.
Appendix E Proofs for inference
In this section, we present the analysis for the inference results we presented in Section 3.3.
E.1 Proof of Theorem 3
We condition the whole analysis on the high probability event when Lemmas 4 and 3 holds. Let $\widehat{\bm{\delta}}\coloneqq\widehat{\bm{\theta}}-\bm{\theta}^{\star}$. Expanding (15), we have
$\widehat{\bm{\delta}}$ $=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
$=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})+\bm{r},$
where $\|\bm{r}\|_{\infty}\leq C_{1}\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C_{1}$.
We start with the case with $\widetilde{\bm{z}}=\bm{z}$. Consider $\bm{L}_{Lz}\widehat{\bm{\delta}}$. As $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$, $\lambda_{m-1}(\bm{L}_{Lz})>0$, and $\widehat{\bm{\delta}}^{\top}\bm{1}_{m}=0$,
$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}$ $=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}$
$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}+\bm{L}_{Lz}^{1/2}\bm{r}.$
Conditional on $(\mathcal{E}_{Y},\{L_{ij}\})$, $\bm{u}_{ij}^{(l)}$ are independent random variables. It is easy to see that $\bm{u}_{ij}^{(l)}$ is zero-mean and has covariance
$\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}.$
It is also bounded in third moment of its spectral norm by
$\mathbb{E}\left[\|\bm{u}_{ij}^{(l)}\|^{3}\right]\leq 2^{3/2}\|\bm{L}_{Lz}^{\dagger}\|^{3/2}\leq\frac{2^{15/2}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{3/2}},$
where the last inequality uses Lemma 3. Summing up across $i,j$ and $l$, we have
$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=(\bm{L}_{Lz}^{\dagger})^{1/2}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}$
$=(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{L}_{Lz}(\bm{L}_{Lz}^{\dagger})^{1/2}$
$=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
The last line holds since $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$ and $\lambda_{m-1}(\bm{L}_{Lz})>0$. Now using multivariate CLT (see, e.g., [Rai19]), we have that as $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$,
$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
As $m$ is fixed, Lemma 2 implies that $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$ is equivalent to $np\rightarrow\infty$. For the residual term, Lemma 4 and Lemma 3 imply that with probability at least $1-O(n^{-10})$,
$\|\bm{L}_{Lz}^{1/2}\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\|\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\sqrt{m}\|\bm{r}\|_{\infty}\leq\frac{C_{1}\sqrt{m}\kappa_{1}^{6}\log^{2}(n)}{\sqrt{np}}.$
Then as $np/\log^{4}(n)\rightarrow\infty$,
$\bm{L}_{Lz}^{1/2}\bm{r}\rightarrow\bm{0}$
in probability. Combining the limiting distribution of both two terms, we have
$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
In the case of $\widetilde{\bm{z}}=\widehat{\bm{z}}$, we introduce the following lemma. The proof is deferred to the end of this section.
Lemma 8.
Instate the assumptions of Theorem 1. Then with probability at least $1-C_{1}n^{-10}$ for some constant $C_{1}$, one has
$\|\bm{L}_{Lz}^{1/2}-\bm{L}_{L\hat{z}}^{1/2}\|\leq C_{2}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}$
for some constant $C_{2}$.
Now the above lemma and Theorem 1 imply
$\left\|\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right\|\leq\|\bm{L}_{L\widetilde{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|\cdot\|\widehat{\bm{\delta}}\|\leq C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{m\log^{2}(n)}{np}}$
for some constant $C_{3}>0$. Therefore, as $np/\log^{4}(n)\rightarrow\infty$,
$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\rightarrow 0$
in probability and hence
$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}=\left(\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right)+\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
Proof of Lemma 8.
Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then
$\|\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|=\|\bm{R}(\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|.$
We also observe that $(\bm{R}\bm{L}_{L\widetilde{z}}\bm{R}^{\top})^{1/2}=\bm{R}\bm{L}_{L\widetilde{z}}^{1/2}\bm{R}^{\top}$ for $\widetilde{z}\in\{z,\widehat{z}\}$. As shown in the proof of Lemma 7,
$\|\bm{L}_{L\hat{z}}-\bm{L}_{Lz}\|\leq C_{1}\kappa_{1}^{3/2}\sqrt{np\log(n)}$
for some constant $C_{1}>0$.
Lemma 3 implies that
$\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})=\left[\lambda_{m-1}(\bm{L}_{Lz})\right]{}^{1/2}\geq\frac{\sqrt{np}}{4\kappa_{1}^{1/2}\kappa_{2}^{1/2}}.$
Then by Lemma 2.1 in [Sch92],
$\|\bm{R}(\bm{L}_{L\widehat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|\leq\frac{\|\bm{R}(\bm{L}_{L\widehat{z}}-\bm{L}_{Lz})\bm{R}^{\top}\|}{\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})}\leq 4C_{1}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}.$
E.2 Proof of Corollary 1
Observe that
$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}=\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}.$
Then Theorem 3 implies
$\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}(\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top})(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{e}_{i}}{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}=1.$
The last inequality follows from the fact that $\bm{1}_{m}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}=0$. As a side note, we compare this with a similar result for the BTL model. Proposition 4.1 in [GSZ23] studies the uncertainty quantification of MLE in the BTL model with uniform observation. It says that
$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{\left[\bm{L}_{Lz}\right]_{ii}^{-1}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$
This is close to our result as $[\bm{L}_{Lz}]_{ii}^{-1}$ can be viewed as an approximation of $[\bm{L}_{Lz}^{\dagger}]_{ii}$ in the context of the BTL model. To see why, consider the simplified case where $\bm{\theta}^{\star}=\bm{0}_{n}$ and $L=1$. As $np\rightarrow\infty$, the Laplacian $\bm{L}_{Lz}$ concentrate to its expectation $\mathbb{E}[\bm{L}_{Lz}]=np\bm{I}_{n}-p\bm{1}_{n}\bm{1}_{n}^{\top}$. Then $\bm{L}^{\dagger}\approx(1/np)(\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{\top})$ and $[\bm{L}_{Lz}^{\dagger}]_{ii}\approx[\bm{L}_{Lz}]_{ii}^{-1}$
Appendix F Auxiliary lemmas
In this section, we gather some auxiliary results that are useful throughout this paper.
Lemma 9 (Range of $z_{ij}$).
Recall
$z_{ij}=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}=\frac{e^{\theta_{i}^{\star}-\theta_{j}^{\star}}}{(1+e^{\theta_{i}^{\star}-\theta_{j}^{\star}})^{2}}.$
For any $(i,j)$,
$\frac{1}{4\kappa_{1}}\leq z_{ij}\leq\frac{1}{4}.$
Proof.
Consider the function $f:[0,\infty)\rightarrow\mathbb{R}$ defined by $f(x)=x/(1+x)^{2}$. It has derivative $(1-x^{2})/(1+x)^{4}$, so it is increasing at $x\in[0,1)$ and decreasing at $x\in(1,\infty)$. By the definition of $\kappa_{1}$, $|\theta_{i}^{\star}-\theta_{j}^{\star}|\leq\log(\kappa_{1})$. Then
$\frac{1}{4\kappa_{1}}\leq f(e^{-\log(\kappa_{1})})\wedge f(e^{\log(\kappa_{1})})\leq z_{ij}\leq f(1)=\frac{1}{4}.$
Lemma 10 (Maximum eigenvalue of Laplacian).
Let $\bm{L}=\sum_{(i,j):i>j}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ be a weighted graph Laplacian. Then $\lambda_{1}(\bm{L})\leq 2\max_{i}\sum_{j}w_{ij}$.
Proof.
Let $\bm{v}\in\mathbb{R}^{m}$, then
$\bm{v}^{\top}\bm{L}\bm{v}$ $=\bm{v}^{\top}\sum_{(i,j):i>j}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{v}$
$=\sum_{(i,j):i>j}w_{ij}(v_{i}-v_{j})^{2}$
$\leq 2\sum_{(i,j):i>j}w_{ij}(v_{i}^{2}+v_{j}^{2})$
$\leq 2\sum_{i}\sum_{j\neq i}w_{ij}v_{j}^{2}$
$\leq 2\sum_{i}\max_{i}\sum_{j}w_{ij}\|\bm{v}\|^{2}$
So $\lambda_{1}(\bm{L})=\max_{\bm{v}\in\mathbb{R}^{m},\|\bm{v}\|=1}\bm{v}^{\top}\bm{L}\bm{v}\leq 2\max_{i}\sum_{j}w_{ij}$.∎
Lemma 11 (A quantitative version of Sylvester’s law of inertia, [Ost59]).
For any real symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$ and $\bm{S}\in\mathbb{R}^{n\times n}$ be a non-singular matrix. Then for any $i\in[n]$, $\lambda_{i}(\bm{S}\bm{A}\bm{S}^{\top})$ lies between $\lambda_{i}(\bm{A})\lambda_{1}(\bm{S}^{\top}\bm{S})$ and $\lambda_{i}(\bm{A})\lambda_{n}(\bm{S}^{\top}\bm{S})$.
Fact 2.
Let $\mathcal{G}$ be an arbitrary graph with $m$ vertices and let $\bm{L}_{w}$ be a weighted graph Laplacian defined by
$\bm{L}_{w}\coloneqq\sum_{i>j:(i,j)\in\mathcal{G}}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
If $w_{ij}>0$ for all $(i,j)\in\mathcal{G}$ and $\mathcal{G}$ is a connected graph, then $\bm{L}_{w}$ is rank $m-1$, $\bm{L}_{w}\bm{1}_{m}=\bm{0}_{m}$ and $\bm{L}_{w}^{\dagger}\bm{1}_{m}=0$. Moreover for any $i\in[n-1]$, $\lambda_{i}(\bm{L}_{w}^{\dagger})=\lambda_{n-i}(\bm{L}_{w})$.
Proof.
The fact that $\bm{L}_{w}$ is rank $m-1$ when $\mathcal{G}$ is connected is well-known. See e.g. [Spi07] for reference. Since $\bm{L}_{w}$ is a real symmetric matrix, it has an eigendecomposition $\bm{L}_{w}=\bm{U}\bm{\Sigma}\bm{U}^{\top}$ and then $\bm{L}_{w}^{\dagger}=\bm{U}\bm{\Sigma}^{\dagger}\bm{U}^{\top}$. The rest follows from this decomposition and the form of $\bm{L}_{w}$.∎

### v2

@@
 Abstract
-The Rasch model, a classical model in the item response theory, is widely used in psychometrics to model the relationship between individuals’ latent traits and their binary responses on assessments or questionnaires.
-In this paper, we introduce a new likelihood-based estimator—random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$) and its bootstrapped variant multiple random pairing MLE ($\mathsf{MRP\text{-}MLE}$) that faithfully estimate the item parameters in the Rasch model.
+The Rasch model, a classical model in the item response theory, is widely used in psychometrics to model the relationship between individuals’ latent traits and their binary responses to assessments or questionnaires.
+In this paper, we introduce a new likelihood-based estimator—random pairing maximum likelihood estimator ($\mathrm{RP\text{-}MLE}$) and its bootstrapped variant multiple random pairing MLE ($\mathrm{MRP\text{-}MLE}$) which faithfully estimate the item parameters in the Rasch model.
 The new estimators have several appealing features compared to existing ones.
@@
 Second, both estimators are provably minimax optimal in terms of finite sample $\ell_{\infty}$ estimation error.
-Lastly, $\mathsf{RP\text{-}MLE}$ admits precise distributional characterization that allows uncertainty quantification on the item parameters, e.g., construction of confidence intervals of the item parameters.
-The main idea underlying $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ is to randomly pair user-item responses to form item-item comparisons.
+Lastly, both admit precise distributional characterization that allows uncertainty quantification on the item parameters, e.g., construction of confidence intervals for the item parameters.
+The main idea underlying $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ is to randomly pair user-item responses to form item-item comparisons.
 This is carefully designed to reduce the problem size while retaining statistical independence.
@@
 1 Introduction
-The item response theory (IRT) [ER13] is a framework widely used in psychometrics to model the relationship between individuals’ latent traits (such as ability or personality) and their responses on assessments or questionnaires.
-It is particularly useful in the development, analysis, and scoring of tests and assessments; see a recent survey [CLLY21] for a statistical account of IRT.
-Among statistical models in IRT, the Rasch model [Ras60] is a simple but fundamental one for modeling binary responses.
-Specifically, for a user $t$ (e.g., test-taker) and an item $i$ (e.g., test problem), Rasch model assumes that the response user $t$ has to item $i$ is binary and obeys
-$\mathbb{P}\left[\text{user $t$ ``beats'' item $i$}\right]=\frac{e^{\zeta_{t}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}},$
+The item response theory (IRT) (ER, 13) is a framework widely used in psychometrics to model the relationship between individuals’ latent traits (such as ability or personality) and their responses to assessments or questionnaires.
+It is particularly useful in the development, analysis, and scoring of tests and assessments; see the recent survey in CLLY (25) for a statistical account of IRT.
+Among statistical models in IRT, the Rasch model (Ras, 60) is a simple but fundamental one for modeling binary responses.
+Specifically, for a user $t$ (e.g., test-taker) and an item $i$ (e.g., test problem), the Rasch model assumes that the response of user $t$ to item $i$ is binary and obeys
+$\mathbb{P}\left[\text{user $t$ ``loses to'' item $i$}\right]=\frac{e^{\theta_{i}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}},$
 where $\zeta_{t}^{\star},\theta_{i}^{\star}\in\mathbb{R}$ are latent traits of user $t$ and item $i$, respectively.
-The term “beats” here refers to positive responses, such as answering an exam question correctly, writing a favorable review of a product, etc.
+The term “loses to” here refers to negative responses, such as answering an exam question incorrectly, writing a negative review of a product, etc.
 We also call this response a comparison between user $t$ and item $i$.
-In this paper, we focus on estimating the item parameters $\bm{\theta}^{\star}$, which is one of the four main statistical tasks surrounding the Rasch model (or IRT more generally) listed in [CLLY21].
-Estimating the item parameters is also quite useful in practice.
+In this paper, we focus on estimating the item parameters $\bm{\theta}^{\star}$, which is one of the four main statistical tasks surrounding the Rasch model (or IRT more generally) listed in CLLY (25).
+Estimating the item parameters is practically important.
 For instance, in education testing, $\bm{\theta}^{\star}$ could reveal the difficulty of the exam questions, while in product reviews, $\bm{\theta}^{\star}$ could reveal the popularity of the products.
-Various methods have been proposed for estimating the item parameters $\bm{\theta}^{\star}$, including the joint maximum likelihood estimator (JMLE), the marginal maximum likelihood estimator, the conditional maximum likelihood estimator (CMLE), as well as the spectral estimator recently proposed in [NZ22, NZ23].
-We refer readers to a recent article [Rob21] for comparisons between different item parameter estimation methods.
+Various methods have been proposed for estimating the item parameters $\bm{\theta}^{\star}$, including the joint maximum likelihood estimator (JMLE), the marginal maximum likelihood estimator, the conditional maximum likelihood estimator (CMLE), and the spectral estimator recently proposed in NZ (22, 23).
+We refer readers to a recent article (Rob, 21) for comparisons between different item parameter estimation methods.
 However, three main gaps remain in tackling item estimation in the Rasch model:
   • Non-asymptotic guarantee.
-Apart from the recently proposed spectral estimator [NZ22, NZ23], most theoretical guarantees for the likelihood-based estimators are asymptotic.
+Apart from the recently proposed spectral estimator (NZ, 22, 23), most theoretical guarantees for the likelihood-based estimators are asymptotic.
 Since all the estimation procedures are necessarily applied with finite samples, the asymptotic guarantee alone fails to inform practitioners about the performance of different estimators when working with a limited number of samples.
   • Sparse observations.
-It is not uncommon to encounter situations where each user only responds to a handful of questions (or items).
-This brings the challenge of incomplete and sparse observations.
-Among the likelihood based estimators, CMLE is known to have numerical issues with incomplete observations [Mol95].
-While the spectral estimator [NZ22, NZ23] is capable of handling incomplete observations, it still requires the observations to be relatively dense.
+It is not uncommon to encounter situations where each user only responds to a handful of questions.
+This brings the challenge of incomplete or sparse observations.
+Many methods, such as CMLE, allow incomplete data (Mol, 95), but most of them lack theoretical support in the sparse regime.
+While the spectral estimator (NZ, 22, 23) is capable of handling incomplete observations, its theory still requires the observations to be relatively dense.
 We will elaborate on this point later.
@@
 However, existing results do not address this problem under sparse observations.
-An exception is the recent work [CLOX23], which is based on joint estimation and inference on the item parameters $\bm{\theta}^{\star}$ and the user parameters $\bm{\zeta}^{\star}$.
-Their sampling scheme is restrictive and requires a relatively dense sampling rate.
-In light of the above gaps, we raise the following natural question:
+An exception is the recent work by CLOX (23), which is based on joint estimation and inference on the item parameters $\bm{\theta}^{\star}$ and the user parameters $\bm{\zeta}^{\star}$.
+Their sampling scheme is more restrictive and requires a relatively dense sampling rate.
+In light of these gaps, we raise the following question:
 Can we develop an estimator for the item parameters $\bm{\theta}^{\star}$ that (1) enjoys optimal estimation guarantee in finite sample, and (2) is amenable to tight uncertainty quantification, when the observations are sparse?
 1.1 Main contributions
-The main contribution of our work is the proposal of a novel estimator named random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$ in short) that achieves the two desiderata listed above.
-In essence, $\mathsf{RP\text{-}MLE}$ compiles user-item comparisons to item-item comparisons by randomly pairing responses of the same user to different items.
+The main contribution of our work is the proposal of a novel estimator named random pairing maximum likelihood estimator ($\mathrm{RP\text{-}MLE}$ in short) that achieves the two desiderata listed above.
+In essence, $\mathrm{RP\text{-}MLE}$ compiles user-item comparisons to item-item comparisons by randomly pairing responses of the same user to different items.
 This pairing procedure is carefully designed to extract information of the item parameters while retaining statistical independence.
 After this compilation step, item parameters $\bm{\theta}^{\star}$ are estimated by the MLE $\widehat{\bm{\theta}}$ given the item-item comparisons.
-Even when the observations are extremely sparse, $\mathsf{RP\text{-}MLE}$ achieves the following:
-  • Regarding estimation, we show that both $\mathsf{RP\text{-}MLE}$ and its bootstrapped version enjoy optimal finite sample $\ell_{\infty}$ error guarantee.
-Compared to the conventional $\ell_{2}$ error guarantee, the $\ell_{\infty}$ guarantee, as an entrywise guarantee is more fine-grained.
-Consequently, we also show that $\mathsf{RP\text{-}MLE}$ can recover the top-$K$ items with minimal sample complexity.
+Even when the observations are extremely sparse, $\mathrm{RP\text{-}MLE}$ achieves the following:
+  • Regarding estimation, we show that both $\mathrm{RP\text{-}MLE}$ and its bootstrapped version enjoy optimal finite sample $\ell_{\infty}$ error guarantee.
+Compared to the conventional $\ell_{2}$ error guarantee, the $\ell_{\infty}$ guarantee, as an entrywise guarantee, is more fine-grained.
+Consequently, we also show that $\mathrm{RP\text{-}MLE}$ can recover the top-$K$ items with optimal sample complexity.
   • While the optimal $\ell_{\infty}$ error guarantee directly yields optimal $\ell_{2}$ guarantee, such guarantee is only correct in an order-wise sense.
-We provide a refined finite-sample $\ell_{2}$ error guarantee of $\mathsf{RP\text{-}MLE}$ that is precise even in the leading constant.
-  • Supplementing the estimation guarantee, we also build an inferential framework based on $\mathsf{RP\text{-}MLE}$ $\widehat{\bm{\theta}}$.
+We provide a refined finite-sample $\ell_{2}$ error guarantee of $\mathrm{RP\text{-}MLE}$ that is precise even in the leading constant.
+  • Supplementing the estimation guarantee, we also build an inferential framework based on $\mathrm{RP\text{-}MLE}$ $\widehat{\bm{\theta}}$.
 More specifically, we precisely characterize the asymptotic distribution of $\widehat{\bm{\theta}}$.
@@
 It often involves a probabilistic model that links categorical responses to latent traits of both users and items.
-Early endeavors include [Ras60] that introduces the Rasch model studied herein and [LNB68] that describes a more general framework using parametric models.
+In the early endeavors, Ras (60) introduced the Rasch model studied herein, and LNB (68) describes a more general framework using parametric models.
 Popular IRT models include the Rasch model, the two-parameter model (2PL), and the three-parameter logistic model (3PL).
-As response data widely appears in real life, IRT finds application in numerous fields including educational assessment [DC10], psychometrics [LNB68], political science [VHSA20], and medical assessment [FBC05].
-See also [CLLY21] for an overview of IRT from statisticians’ perspective.
+As response data widely appears in real life, IRT finds application in numerous fields including educational assessment (DC, 10), psychometrics (LNB, 68), political science (VHSA, 20), and medical assessment (FBC, 05).
+See CLLY (25) for an overview of IRT.
 Latent score estimation for Rasch model.
-An important statistical question in the Rasch model is to estimate the latent parameters of the items.
-As the Rasch model is an explicit probabilistic model, many methods are based on the principle of maximum likelihood estimation.
+An important statistical question in the Rasch model is to estimate the item parameters.
+As the Rasch model is an explicit probabilistic model, many estimation methods are based on the principle of maximizing likelihood.
 For instance, marginal MLE assumes a prior on the user parameters that is either given or optimized within a parametric distribution family.
@@
 both the item and user parameters.
-However, it is not consistent for estimating the item parameters when the number $m$ of items is fixed [Gho95].
-Interested readers may also consult [Lin99] for an overview of other classical estimators.
-Several methods are more relevant to our proposed estimator $\mathsf{RP\text{-}MLE}$ as they follow a similar philosophy to form item-item pairs from user-item responses.
-Pseudo MLE (PMLE) [Zwi95] maximizes the sum of the log-likelihood of all pairs of responses from the same users to different items.
-However, due to the dependency issue, no satisfying finite sample performance guarantee has been established.
+However, it is not consistent for estimating the item parameters when the number of items is fixed (Gho, 95).
+Interested readers may also consult Lin (99) for an overview of other classical estimators.
+Several methods are more relevant to our proposed estimator $\mathrm{RP\text{-}MLE}$ as they follow a similar philosophy to form item-item comparisons from user-item responses.
+Conditional MLE (CMLE) considers the tuple of all items that are related to a user instead of examining the induced item pairs.
+It maximizes the likelihood condition on the total number of positive response a user has.
+Theoretically, And (73) has shown that CMLE is asymptotically normal and consistent.
+However, no non-asymptotic rate in the setting of sparse observation has been established.
+Pseudo MLE (PMLE) (Zwi, 95) maximizes the sum of the log-likelihood of all pairs of responses from the same users to different items.
+However, due to the dependence, no satisfying finite sample performance guarantee has been established.
 Another related approach is the spectral method, in which a Markov chain on the space of items is formed and the item parameters are estimated via the stationary distribution of the Markov chain.
-The most recent works in this category are [NZ22, NZ23], which essentially use the same idea as pseudo MLE in forming item-item comparisons.
+The most recent works in this category are NZ (22, 23), which essentially use the same idea as pseudo MLE in forming item-item comparisons.
 The Bradley-Terry-Luce model with sparse comparisons.
-An informed reader may realize that the Rasch model resembles the Bradley-Terry-Luce (BTL) model [Luc59, BT52] in the ranking literature.
+An informed reader may realize that the Rasch model resembles the Bradley-Terry-Luce (BTL) model (Luc, 59; BT, 52) in the ranking literature.
 Indeed, one can view the Rasch model as a special case of the BTL model that distinguishes the two groups of users and items, and only allows inter-group comparisons.
-There has been a recent surge in interest in studying top-$K$ ranking in the BTL model [CS15, JKSO16, CFMW19, CGZ22, GSZ23] and its extensions [FHY22, FLWY22, FLWY23], especially under sparse observations of the pairwise responses.
-Most notably, under a uniform sampling scheme, [CFMW19] shows that (regularized) MLE and spectral methods are both optimal in top-$K$ ranking and [GSZ23] provides inference results for both methods.
-Going beyond uniform sampling, [Che23, LSR22] investigate the performance of MLE in the BTL model with a general comparison graph and later [YCOM24] improves the analysis to show the optimality of MLE for the BTL model in both uniform and semi-random sampling.
+There has been a recent surge in interest in studying top-$K$ ranking in the BTL model (SY, 99; YYX, 12; CS, 15; JKSO, 16; CFMW, 19; HYTC, 20; CGZ, 22; GSZ, 23; LFL, 23) and its extensions (HX, 23; HXC, 23; FHY24a, ; FHY24b, ; FLWY25b, ; FLWY25a, ), especially under sparse observations of the pairwise responses.
+Most notably, under a uniform sampling scheme, CFMW (19) shows that (regularized) MLE and spectral methods are both optimal in top-$K$ ranking and GSZ (23) provides inference results for both methods.
+Another line of research focuses on non-uniform sampling.
+HOX (14) and SBB+ (16) each studies the $\ell_{2}$ error of MLE with a general sampling graph.
+They obtain high probability upper bound and minimax lower bound that match under in some scenarios.
+In particular, SBB+ (16) extensively discusses the implication of their result in different graph topology.
+More recently, general sampling graph has also been studied in top-$K$ ranking.
+Several articles (HX, 23; Che, 23; LSR, 22) investigate the performance of MLE in the BTL model with a general comparison graph and later YCOM (24) improves the analysis to show the optimality of (weighted) MLE for the BTL model in both uniform and semi-random sampling.
 Notation.
-For a positive integer $n$, we denote $[n]=\{1,2,\ldots,n\}$.
-For any $a,b\in\mathbb{R}$, $a\wedge b$ means the minimum of $a,b$ and $a\vee b$ means the maximum of $a,b$.
-We use $\overset{\mathrm{d}}{\rightarrow}$ to denote convergence in distribution.
+For a positive integer $n$, we denote $n]=\{1,2,\ldots,n\}$.
+For any $\in\mathbb{R}, aba\wedge b$ means the minimum of $a,b$ and $a\vee b$ means the maximum of $a,b$.
+We use $\overset{\mathrm{a.s}}{\rightarrow}$, $\overset{\mathrm{p}}{\rightarrow}$, and $\overset{\mathrm{d}}{\rightarrow}$ to denote convergence almost surely, in probability, and in distribution respectively.
 For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$, we use $\lambda_{1}(\bm{A})\geq\lambda_{2}(\bm{A})\geq\cdots\geq\lambda_{n}(\bm{A})$ to denote its eigenvalues and $\bm{A}^{\dagger}$ to denote its Moore-Penrose pseudo-inverse.
@@
 Unless specified otherwise, $\log(\cdot)$ denotes the natural log.
-2 Problem setup and new estimator $\mathsf{RP\text{-}MLE}$
+2 Problem setup and new estimators
 In this section, we first introduce the formal setup of the item parameter estimation problem in the Rasch model.
-Then we present the newly proposed estimator $\mathsf{RP\text{-}MLE}$ along with the rationale behind its development.
+Then we present the news estimator $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ along with the rationale behind its development.
 2.1 Problem setup
@@
 Here $X_{ti}=1$ means user $t$ has negative response against item $i$ (e.g., unable to solve a problem). The goal is to estimate $\bm{\theta}^{\star}$, the item parameters.
-To model sparse observations/comparisons, we assume that $\mathbb{P}[(t,i)\text{ is compared}]=p$ independently for every $(t,i)\in[n]\times[m]$.
+To model sparse user-item responses, we assume that $\mathbb{P}[(t,i)\text{ is compared}]=p$ independently for every $(t,i)\in[n]\times[m]$.
 To put it in the language of graph theory, we denote the associated bipartite comparison graph to be $\mathcal{G}_{X}=(\mathcal{V}_{X},\mathcal{E}_{X})$, where $\mathcal{V}_{X}$ consists of $n$ users and $m$ items.
@@
 Our estimator and performance guarantee continue to work in this sampling scheme.
-Before moving on, we introduce the condition numbers that characterize the range of the latent traits.
-Let $\kappa_{1}$, $\kappa_{2}$, and $\kappa$ be defined by $\log\;(\kappa_{1})=\max_{ij}\{|\theta_{i}^{\star}-\theta_{j}^{\star}|\}$, $\log(\kappa_{2})=\max_{ti}\{|\zeta_{t}^{\star}-\theta_{i}^{\star}|\}$, and $\kappa=\max\{\kappa_{1},\kappa_{2}\}$, respectively.
+Before moving on, we define condition numbers to characterize the range of the latent traits.
+Let $\kappa_{1}$, $\kappa_{2}$, and $\kappa$ be defined by $\log\;(\kappa_{1})\coloneqq\max_{ij}\{|\theta_{i}^{\star}-\theta_{j}^{\star}|\}$, $\log(\kappa_{2})\coloneqq\max_{ti}\{|\zeta_{t}^{\star}-\theta_{i}^{\star}|\}$, and $\kappa\coloneqq\max\{\kappa_{1},\kappa_{2}\}$, respectively.
 2.2 Random pairing maximum likelihood estimator
-Algorithm 1 Random Pairing Maximum Likelihood Estimator ($\mathsf{RP\text{-}MLE}$)
+Algorithm 1 Random Pairing Maximum Likelihood Estimator ($\mathrm{RP\text{-}MLE}$)
   1. 1.
@@
     1. (a)
-      Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems.
+      Randomly split the $m_{t}$ problems taken by tester $t$ into $\lfloor m_{t}/2\rfloor$ pairs of problems.
     2. (b)
@@
   2. 2.
-    Let $\mathcal{E}_{Y}$ be a set of edges defined by $\mathcal{E}_{Y}\coloneqq\{(i,j):\sum_{t=1}L_{ij}^{t}\geq 1\}$ and let $\mathcal{G}_{Y}=([m],\mathcal{E}_{Y})$.
+    Let $\mathcal{E}_{Y}$ be a set of edges defined by $\mathcal{E}_{Y}\coloneqq\{(i,j):\sum_{t=1}^{n}L_{ij}^{t}\geq 1\}$ and let $\mathcal{G}_{Y}=([m],\mathcal{E}_{Y})$.
 For each $(i,j)\in\mathcal{E}_{Y}$, let $L_{ij}\coloneqq\sum_{t=1}^{n}L_{ij}^{t}$ and $Y_{ij}\coloneqq(1/L_{ij})\sum_{\{t:L_{ij}^{t}=1\}}Y_{ij}^{t}$.
   3. 3.
-    Compute MLE on $Y_{ij}$, i.e.,
-    $\widehat{\bm{\theta}}\coloneqq\arg\min_{\bm{\theta}:\bm{1}_{m}^{\top}\bm{\theta}=0}\mathcal{L}(\bm{\theta})$
-    where
+    Compute MLE on $Y_{ij}$, i.e., $\widehat{\bm{\theta}}\coloneqq\arg\min_{\bm{\theta}:\bm{1}_{m}^{\top}\bm{\theta}=0}\mathcal{L}(\bm{\theta})$, where
     $\mathcal{L}(\bm{\theta})=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\left(-Y_{ji}(\theta_{i}-\theta_{j})+\log(1+e^{\theta_{i}-\theta_{j}})\right).$ (2)
@@
     Return the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}$.
-In this section, we present our main method $\mathsf{RP\text{-}MLE}$; see Algorithm 1.
+In this section, we present our main method $\mathrm{RP\text{-}MLE}$; see Algorithm 1.
 The algorithm can be divided into two parts.
@@
 The second part—Steps 3 and 4—computes a standard MLE on the item-item comparisons.
-Some intuitions regarding the development of $\mathsf{RP\text{-}MLE}$ are in order.
+Some intuition regarding the development of $\mathrm{RP\text{-}MLE}$ is in order.
 Random pairing to construct item-item comparisons.
 The idea of pairing is that by matching the responses $X_{ti}$ with $X_{tj}$, we form a comparison between items $i$ and $j$ to directly extract information of item parameters $\theta_{i}^{\star}$ and $\theta_{j}^{\star}$.
-More specifically, the item-item comparisons $\bm{Y}$ follow the Bradley-Terry-Luce model [BT52, Luc59],i.e., $\mathbb{P}[Y_{ij}^{t}]=e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})$; see Section 4.1 for a formal argument.
+More specifically, the item-item comparisons $\bm{Y}$ follow the Bradley-Terry-Luce model (BT, 52; Luc, 59), i.e., $\mathbb{P}[Y_{ij}^{t}=1]=e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})$.
+For the claims made in this part, please see Section A.1 for a formal argument.
 By compiling user-item responses to item-item comparisons, we reduce the size of the data matrix from $n\times m$ to $m\times m$, and also the number of intrinsic parameters from $n+m$ to $m$, since the likelihood function (2) of $Y_{ij}^{t}$ is completely independent with the user parameter $\zeta_{t}^{\star}$.
 More importantly, the pairing is performed in a disjoint fashion.
-This ensures that all constructed item-item comparisons $Y_{ij}^{t}$ are independent with each other; see Section 4.1 for a formal statement.
-This is the key ingredient that enables us to improve over previous implementation of item-item comparisons, such as pseudo-likelihood [Cho82, Zwi95] and spectral methods [NZ22, NZ23].
+This ensures that all constructed item-item comparisons $Y_{ij}^{t}$ are independent with each other; see Section A.1 for a formal statement.
+This is the key ingredient that enables us to improve over previous implementation of item-item comparisons, such as pseudo-likelihood (Cho, 82; Zwi, 95) and spectral methods (NZ, 22, 23).
 A variant via bootstrapping.
@@
 Once $X_{ti}$ is paired with $X_{tj}$, we cannot pair $X_{ti}$ with another response $X_{tl}$.
-That being said, we will later show that the $\ell_{\infty}$ error of $\mathsf{RP\text{-}MLE}$ is still rate-optimal up to logarithmic factors.
+That being said, we will later show that the $\ell_{\infty}$ error of $\mathrm{RP\text{-}MLE}$ is still rate-optimal up to logarithmic factors.
 Hence the loss of information can at most incur a small constant factor in terms of estimation error.
-Nevertheless, we provide a remedy to this phenomenon in $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by running (in other words, bootstrapping) the $\mathsf{RP\text{-}MLE}$ multiple times with different random data splitting and averaging the resulting estimates.
-$\mathsf{MRP\text{-}MLE}$ trivially enjoys the same theoretical guarantee (cf.
-Theorem 1) while improving the estimation error in practice over $\mathsf{RP\text{-}MLE}$.
+Nevertheless, we provide a remedy to this phenomenon in $\mathrm{MRP\text{-}MLE}$ (Algorithm 2) by running (in other words, bootstrapping) the $\mathrm{RP\text{-}MLE}$ multiple times with different random data splitting and averaging the resulting estimates.
+$\mathrm{MRP\text{-}MLE}$ trivially enjoys the same non-asymptotic $\ell_{\infty}$ error rate (cf.
+Theorem 1) while improving the estimation error in practice over $\mathrm{RP\text{-}MLE}$.
 See Figure 4 in Section 4.1 for the empirical evidence.
-Algorithm 2 Multiple Random Pairing Maximum Likelihood Estimator ($\mathsf{MRP\text{-}MLE}$)
+Algorithm 2 Multiple Random Pairing Maximum Likelihood Estimator ($\mathrm{MRP\text{-}MLE}$)
   1. 1.
     Let $n_{\mathrm{split}}$ be the number of runs.
-For $i=1,\ldots,n_{\mathrm{split}}$, run $\mathsf{RP\text{-}MLE}$ (Algorithm 1), each time with an independent random splitting in Step 1.
+For $i=1,\ldots,n_{\mathrm{split}}$, run $\mathrm{RP\text{-}MLE}$ (Algorithm 1), each time with an independent random splitting in Step 1.
 Let the estimated latent scores be $\widehat{\bm{\theta}}^{(i)}$.
@@
     Estimate the latent score with
-    $\widehat{\bm{\theta}}=\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}^{(i)}.$
+    $\widehat{\bm{\theta}}_{\mathrm{MRP}}=\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}_{(i)}.$
   3. 3.
-    Estimate the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}$.
+    Estimate the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}_{\mathrm{MRP}}$.
 3 Main results
-In this section, we collect the main theoretical guarantees for $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$.
+In this section, we collect the main theoretical guarantees for $\mathrm{RP\text{-}MLE}$ and its variant $\mathrm{MRP\text{-}MLE}$.
 Section 3.1 focuses on the finite sample $\ell_{\infty}$ error bound.
-While one can translate $\ell_{\infty}$ error into an $\ell_{2}$ bound, in Section 3.2, we present a much sharper characterization of the $\ell_{2}$ error of $\mathsf{RP\text{-}MLE}$.
-Lastly in Section 3.3, we provide a distributional characterization of $\mathsf{RP\text{-}MLE}$.
-3.1 $\ell_{\infty}$ estimation error and top-$K$ recovery
+In Section 3.2, we present a non-asymptotic expansion that describes the distribution of $\mathrm{RP\text{-}MLE}$, and we apply this expansion to reach a Berry-Esseen type theorem and a much sharper characterization of the $\ell_{2}$ error of $\mathrm{RP\text{-}MLE}$.
+Lastly in Section 3.3, we prove the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ when $m$ and $p$ are fixed and draw a connection between $\mathrm{MRP\text{-}MLE}$ and (weighted) pseudo MLE.
+3.1 $\ell_{\infty}$ error bounds and top-$K$ recovery
 Without loss of generality, we assume that the scores of the items are ordered, i.e., $\theta_{1}^{\star}\geq\theta_{2}^{\star}\geq\cdots\geq\theta_{m}^{\star}$, and denote $\Delta_{K}\coloneqq\theta_{K}^{\star}-\theta_{K+1}^{\star}$.
 In words, $\Delta_{K}$ measures the difference between the difficulty levels of items $K$ and $K+1$.
-The following theorem provides $\ell_{\infty}$ error bounds and top-$K$ recovery guarantee for both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$.
-We defer its proof to Section 4.2.
+The following theorem provides $\ell_{\infty}$ error bounds and top-$K$ recovery guarantee for both $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$.
+We defer its proof to Section A.2.
 Theorem 1.
@@
 Suppose that there exists some constant $\alpha>0$ such that $m\leq n^{\alpha}$.
-Let $\widehat{\bm{\theta}}$ be the $\mathsf{RP\text{-}MLE}$ estimator.
+Let $\widehat{\bm{\theta}}$ be the $\mathrm{RP\text{-}MLE}$ estimator.
 With probability at least $1-O(n^{-10})$, $\widehat{\bm{\theta}}$ satisfies
@@
 Here $C_{2},C_{3}>0$ are some universal constants.
-All the claims continue to hold for $\mathsf{MRP\text{-}MLE}$ as long as there exists some constant $\beta>0$ such that $n_{\mathrm{split}}\leq n^{\beta}$.
+All the claims continue to hold for $\mathrm{MRP\text{-}MLE}$ as long as there exists some constant $\beta>0$ such that $n_{\mathrm{split}}\leq n^{\beta}$.
 Some remarks are in order.
 Finite sample minimax optimality.
-Intuitively, as we have $\Theta(mnp)$ sampled comparisons and $m$ parameters to estimate, the optimal $\ell_{\infty}$ error should be $O(\sqrt{m/(mnp)})=O(1/\sqrt{np})$.
-This is formalized in the following result from [NZ23].
-Proposition 1 (Minimax lower bound, Theorems 3.3 and 3.4 in [NZ23]).
-Assume that $np\geq C_{1}$ for some suffiently large constant $C_{1}>0$. For any $n$ and $m$, there exists a class of user and item parameters $\Theta$ such that
-$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{2}^{2}\geq\frac{C_{2}m}{np},\qquad\text{and}\qquad\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{\infty}^{2}\geq\frac{C_{2}}{np},$
+Based on results from the ranking literature, it is reasonable to guess that the optimal $\ell_{\infty}$ error is $O(\sqrt{m/(mnp)})=O(1/\sqrt{np})$.
+Here $m$ is the number of item parameters, and mnpmnp is the number of user-item comparisons.
+This guess is indeed correct, as formalized in the following result from NZ (23).
+Proposition 1 (Minimax lower bound, Theorems 3.3 and 3.4 in NZ (23)).
+Assume that $np\geq C_{1}$ for some sufficiently large constant $C_{1}>0$. For any $n$ and $m$, there exists a class of user and item parameters $\Theta$ such that
+$\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{2}^{2}\geq\frac{C_{2}m}{np},\qquad\text{and}\qquad\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{E}\left\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\right\|_{\infty}^{2}\geq\frac{C_{2}}{np},$
 where $C_{2}>0$ is some constant. Moreover if $np\leq C_{K}\log(m)/\Delta_{K}^{2}$ for some constant $C_{K}>0$, we have
-$\inf_{\hat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{P}\left[\widehat{\bm{\theta}}\text{ fails to identify all top-}K\text{ items}\right]\geq\frac{1}{2}.$
-Comparing our upper bounds with the lower bound in the proposition, we can see that both $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ are rate-optimal in $\ell_{\infty}$ estimation error and top-$K$ recovery sample complexity, up to logarithmic and $\kappa$ factors.
+$\inf_{\widehat{\bm{\theta}}}\sup_{(\bm{\zeta}^{\star},\bm{\theta}^{\star})\in\Theta}\mathbb{P}\left[\widehat{\bm{\theta}}\text{ fails to identify all top-}K\text{ items}\right]\geq\frac{1}{2}.$
+Comparing our upper bounds with the lower bound in the proposition, we can see that both $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ are rate-optimal in $\ell_{\infty}$ estimation error and top-$K$ recovery sample complexity, up to logarithmic and $\kappa$ factors.
 Sample size requirement.
@@
 In Theorem 1 we require an extra $\log^{2}(n)$ factor to suppress a quadratic error term that comes up in the analysis.
-This cubic log factor can possibly be loose, but it is a minor issue and we leave it as a problem for future research.
+This cubic log factor can possibly be loose, but it is a minor issue and we leave it to future research.
 Lastly, $m\leq n^{\alpha}$ and $n_{\mathrm{split}}\leq n^{\beta}$ are both minor as we only need these to allow union bounds over $m$ and $n_{\mathrm{split}}$.
-Comparison with [NZ23].
-The closest result to our paper in terms of $\ell_{\infty}$ guarantee for the Rasch model appears in the recent work [NZ23].
+Comparison with NZ (23).
+The closest result to our paper in terms of $\ell_{\infty}$ guarantee for the Rasch model appears in the recent work by NZ (23).
 Their spectral method uses a similar construction of the item-item comparisons but without disjoint pairing.
 To provide detailed comparisons, we restate their results below.
-Proposition 2 (Informal, Theorem 3.1 in [NZ23]).
+Proposition 2 (Informal, Theorem 3.1 in NZ (23)).
 Assume that $p\gtrsim\log(m)/\sqrt{n}$ and $mp\gtrsim\log(n)$, with probability at least $1-O(m^{-10}+n^{-10})$, spectral estimator $\widehat{\bm{\theta}}_{\mathrm{spectral}}$ satisfies
@@
 $mp\geq 2$ and $np\gtrsim\log^{3}(n)$.
-In fact, as we have argued earlier, it is nearly the sparsest possible.
+In fact, as we have argued earlier, it is nearly the sparsest possible regime for estimating item parameters.
 In addition, our methods enjoy a significantly better error rate dependency on $\kappa$.
-In Section 5.1, we provide empirical evidence for this improvement: when $\kappa$ is large, $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ outperform the spectral methods in [NZ23].
-3.2 Refined $\ell_{2}$ error characterization
+In Section 4.1, we provide empirical evidence for this improvement: when $\kappa$ is large, $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ outperform the spectral methods in NZ (23).
+Analysis via reduction to BTL model.
+As mentioned in Section 2.2, the random pairing in $\mathrm{RP\text{-}MLE}$ reduces the problem to the Bradley-Terry-Luce model with non-uniform sampling.
+This reduction results in an item-item comparison graph with nice spectral properties, and allows us to invoke the general theory of MLE in the BTL model established in the recent work by YCOM (24).
+See Section A.2 for the full analysis.
+3.2 Non-asymptotic expansion of $\mathrm{RP\text{-}MLE}$
+An important aspect of statistical estimators is the quantification of the variability.
+In this section we provide a non-asymptotic expansion, which precisely characterizes of the distribution of the $\mathrm{RP\text{-}MLE}$ estimator $\widehat{\bm{\theta}}$.
+We supplement this result with a Berry-Esseen theorem and an application to obtain a precise $\ell_{2}$ error characterization.
+We start with some necessary notation.
+Let $\sigma(x)=e^{x}/(1+e^{x})$ be the sigmoid function.
+Let $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$, $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$ and $\epsilon_{ij}^{t}\coloneqq Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$.
+Let $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$ be the total number of observed comparisons in $\mathcal{G}_{Y}$.
+We define $\bm{\bm{B}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ and $\widehat{\bm{\epsilon}}\in\mathbb{R}^{L_{\mathrm{total}}}$ via
+$\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\text{$\quad$(repeat $L_{ij}$ times for edge $(i,j)$)}$
+and
+$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{t}/\sqrt{z_{ij}},\cdots\right]_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}\in\mathbb{R}^{L_{\mathrm{total}}}.$
+Moreover, define a weighted graph Laplacian
+$\bm{L}_{L\tilde{z}}\coloneqq\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (3)
+for $\tilde{z}$ being $z$ or $\widehat{z}$, and let $\bm{L}_{L\tilde{z}}^{\dagger}$ be its pseudo-inverse.
+The following theorem characterizes the distribution of $\mathrm{RP\text{-}MLE}$ with a non-asymptotic expansion. The analysis is deferred to Section A.3 and the full proof is deferred to Section C.1.
+Theorem 2.
+Instate the assumptions of Theorem 1. With probability at least $1-O(n^{-10})$, the estimator $\widehat{\bm{\theta}}$ given by the Algorithm 1 can be written as
+$\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ $=-\left[\nabla^{2}\mathcal{L}(\bm{\theta}^{\star})\right]^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\star})+\bm{r}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r},$ (4)
+where $\bm{r}\in\mathbb{R}^{m}$ is a random vector obeying $\|\bm{r}\|_{\infty}\leq C\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C>0$.
+This theorem shows $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ can be well approximated by $\nabla^{2}\mathcal{L}(\bm{\theta}^{\star})]^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\star})$, a form that frequently appears in the analysis of maximum likelihood estimators.
+Moreover, it can be written as a linear transformation of the random vector $\widehat{\bm{\epsilon}}$, whose entries are independent outcomes of the item-item comparisons shifted and scaled to be mean-zero and variance-one.
+The term $\bm{L}_{Lz}^{\dagger}\bm{B}$ accounts for the geometry induced by the comparison graph $\mathcal{G}_{Y}$.
+As the residual term $\bm{r}$ has small magnitude, we may analyze the properties of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ by focusing on the leading term $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$.
+We compare our result with the inference result in CLOX (23).
+Theorem 8 therein studies the Rasch model and provides inferential results for a joint estimator of $\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}$.
+However, it requires a dense sampling scheme when $n\geq m$, with
+$p\gtrsim\sqrt{\frac{1}{m}}\vee\frac{\log^{2}(m)}{n}\vee\frac{n\log^{2}(n)}{m^{2}}.$
+It also requires that both $m$ and $n$ tend to infinity. These two assumptions are significantly more restrictive than ours.
+Normal approximation.
+The main term in (4) can be approximated with a normal random variable, allowing various applications such as hypothesis testing on $\bm{\theta}^{\star}$.
+Formally, we present the following Berry-Esseen type theorem.
+The proof is deferred to Section C.2.
+Proposition 3.
+Instate the assumptions of Theorem 1.
+Let $\bm{x}$ be a normal random variable in $\mathbb{R}^{m}$ with variance $\bm{L}_{Lz}^{\dagger}$.
+Let $\mathcal{C}_{m}$ be the set of all the measurable convex subset of $\{\bm{\theta}\in\mathbb{R}^{m}:\bm{\theta}^{\top}\bm{1}_{m}=0\}$.
+Then we have that
+$\sup_{A\in\mathcal{C}_{m}}\left|\mathbb{P}\left[\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\in A\mid\mathcal{G}_{Y}\right]-\mathbb{P}(\bm{x}\in A)\right|\leq C_{1}n^{-10}+C_{2}\frac{m^{5/4}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{1/2}}.$
+Refined $\ell_{2}$ error characterization.
+Another possible application of Theorem 2 is a refined characterization of the $\ell_{2}$ estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$.
 The $\ell_{\infty}$ error guarantee in Theorem 1 immediately implies an $\ell_{2}$ error bound
@@
 However, this guarantee is only correct in an order-wise sense.
-In this section, we present a refined characterization of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ that is precise in the leading constant.
-We start with necessary notation.
-Let $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$ and $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$.
-Define a matrix
-$\bm{L}_{L\tilde{z}}=\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (4)
-for $\tilde{z}\in\{z,\widehat{z}\}$ and let $\bm{L}_{L\tilde{z}}^{\dagger}$ be its pseudo-inverse.
-One can view $\bm{L}_{L\tilde{z}}$ as a weighted graph Laplacian.
-In the following theorem, we show that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrates tightly around $[\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})]^{1/2}$.
-We defer the analysis to Section 4.3 and the complete proof to Section D.
-Theorem 2.
+Here, we present a refined characterization of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ that is precise in the leading constant.
+In the following theorem, we show that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrates tightly around $\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})]^{1/2}$.
+We defer the complete proof to Section C.3.
+Proposition 4.
 Instate the assumptions of Theorem 1. Then for some constants $C_{1},C_{2}>0$, with probability at least $1-O(n^{-10})$, we have
-$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np},$ (5)
+$\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}\kappa_{1}^{6}\sqrt{m}\log^{2}(n)}{np};$ (6)
 $\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})}\right|$ $\leq C_{1}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{\log(n)}{np}}+\frac{C_{2}(\kappa_{1}^{6}+\kappa_{1}^{7/2}\kappa_{2}^{2})\sqrt{m}\log^{2}(n)}{np}.$ (7)
@@
 First, it provides both upper and lower bounds for $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$.
-Second, there is no hidden constant in front of the leading term $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$.
+Second, there is no hidden constant in front of the leading term $\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$.
 In addition, inspecting the proof of Theorem 4, we see that
 $\sqrt{\frac{m-1}{np}}\leq\sqrt{\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})}\leq 4\kappa_{1}^{1/2}\kappa_{2}^{1/2}\sqrt{\frac{m}{np}},$
-and the same holds for $[\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})]^{1/2}$.
-Consequently, the right hand sides of both (5) and (6) are lower order terms compared to $[\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$ when $n,m\rightarrow\infty$.
+and the same holds for $\mathrm{Trace}(\bm{L}_{L\widehat{z}}^{\dagger})]^{1/2}$.
+Consequently, the right hand sides of both (6) and (7) are lower order terms compared to $\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})]^{1/2}$ when $n,m\rightarrow\infty$.
 Indeed this recovers the naive $\ell_{2}$ bound (5) under appropriate sample size assumptions.
-3.3 Uncertainty quantification
-Now we move on to uncertainty quantification about the $\mathsf{RP\text{-}MLE}$ estimator $\widehat{\bm{\theta}}$ in order to perform, say, hypothesis testing on the item parameters $\bm{\theta}^{\star}$.
-Our starting point is the following characterization of the limiting distribution of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$.
-The analysis is deferred to Section 4.3 and the complete proof is deferred to Section E.
+Analysis via projected gradient descent trajectory.
+Inspired by Che (23), we analyze MLE via the projected gradient descent trajectory.
+This approach gives us an alternative to the leave-one-out type argument in CFMW (19); CGZ (22) and the leave-two-out argument in GSZ (23).
+In particular, unlike the leave-one-out and leave-two-out arguments, the projected gradient descent approach does not require independence in the sampling of the compared pairs.
+This is crucial to our analysis as the disjoint pairing in Step 1(a) of Algorithm 1 induces dependent item-item edges.
+See Section A.3 for the full analysis.
+3.3 Asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and pseudo MLE
+In this section, we consider the inference setting where $m$ and $p$ are fixed and $n$ tends to infinity.
+We establish the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and connect it to a weighted variant of pseudo MLE, which we call $\mathrm{WP\text{-}MLE}$.
+We show that these two estimators are asymptotically equal in distribution.
+We start by describing the setup of an infinite sequence of estimators $\{\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}\}$ and relevant notations.
+We will consider the user parameters to be fixed and study the asymptotic normality of $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ that accounts for the randomness in sampling, random pairing, and the comparison between paired outcomes.
+We use subscripts to denote the source of randomness in each expectation.
+For sampling, i.e., the generation of comparison graph $\mathcal{G}_{X}$, we use $\mathbb{E}_{\mathrm{s}}$; for random pairing, where we match item-item-user tuple, we use $\mathbb{E}_{\mathrm{r}}$; for forming an item-item comparison, i.e., determining whether the responses $X_{ti}$ and $X_{tj}$ are different for a given item-item-user tuple $(i,j,t)$, we use $\mathbb{E}_{\mathrm{d}}$; for comparison, i.e., given $X_{ti}\neq X_{tj}$, whether $X_{ti}<X_{tj}$, we use $\mathbb{E}_{\mathrm{c}}$.
+Multiple letters can be combined with + sign in the subscript.
+Let $\mathcal{L}_{k}(\bm{\theta})$ be the loss function for $\mathrm{RP\text{-}MLE}$ using $k$-th random splitting and $\mathcal{L}_{k}^{(t)}(\bm{\theta})$ be the sum of the terms corresponding to user $t$, i.e., $\mathcal{L}_{k}(\bm{\theta})\coloneqq\sum_{t=1}^{n}\mathcal{L}_{k}^{(t)}(\bm{\theta})$ and
+$\mathcal{L}_{k}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j:\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
+Here $\Omega_{k}$ denotes the set of paired item-item-user tuples for the $k$-th splitting.
+We assume there is an infinite sequence of users, which has an infinite sequence of user parameters $\{\zeta_{n}^{\star}\}_{n\in\mathbb{N}^{+}}$, random sampling $\{A_{it}:i\in[m]\}_{t=1}^{\infty}$, and responses $\{X_{it}:A_{it}=1\}_{t=1}^{\infty}$.
+Moreover we assume there are random splittings $\Omega_{k}$ for $k=1,\ldots,n_{\mathrm{split}}$.
+We label the estimators as $\widehat{\bm{\theta}}_{(k)}^{(n)}$ to denote the $\mathrm{MRP\text{-}MLE}$ with $k$-th splitting and the random sampling, responses, and random splittings associated with the first $n$ users.
+Similarly we use $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ to denote $\mathrm{WP\text{-}MLE}$ with first $n$ users.
+Moreover, we suppose that the following limits of average expectation and covariance matrices exist:
+$\bm{H}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star});$ (8a)
+$\bm{V}_{\mathrm{same}}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top};$ (8b)
+$\bm{V}_{\mathrm{diff}}^{\infty}$ $\coloneqq\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}$ (8c)
+Note that by the Cauchy-Schwarz inequality (see Section D.4 for a complete proof), we have
+$\bm{V}_{\mathrm{diff}}^{\infty}\preceq\bm{V}_{\mathrm{same}}^{\infty}.$ (9)
+The two sides are the same when $m_{t}=2$ for all $t$.
+With this setup of a infinite sequence of estimators, we may study the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ as $n$ tends to infinity, accounting for all randomness in $\mathbb{E}_{\mathrm{s+r+d+c}}$.
+We have the following result on $\mathrm{MRP\text{-}MLE}$ $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}=(1/n_{\mathrm{split}})\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}_{(i)}^{(n)}$.
+The analysis is deferred to Section A.4 and the full proof is deferred to Section D.1.
 Theorem 3.
-Instate the assumptions of Theorem 1 and fix $m$. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
-$\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
-Though simple, we record the following immediate consequence on the asymptotic normality of $\widehat{\theta}_{i}-\theta_{i}^{\star}$ for any single coordinate $i$.
-Corollary 1.
-Instate the assumptions of Theorem 1. As $np/\log^{4}(n)\rightarrow\infty$, for $\tilde{z}\in\{z,\widehat{z}\}$, we have
-$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$ (7)
-We compare our result with that in [CLOX23].
-Theorem 8 therein studies the Rasch model and provides an inference result for a joint estimator of $\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}$.
-However, it requires a dense sampling scheme with $p\gtrsim 1/\sqrt{\log(n)/n}$ and $p\gtrsim n\log^{2}(n)/m^{2}$.
-It also requires that both $m$ and $n$ tend to infinity.
-These two assumptions are more restrictive than ours.
-Theorem 3 is quite general as it allows various applications to inference.
-As an illustrative example, we demonstrate how to construct confidence intervals for a particular item parameter $\theta_{i}^{\star}$.
-Fix any $\alpha>0$, the level of significance. Define
-$\mathcal{C}_{i}^{+}(\alpha)$ $\coloneqq\widehat{\theta}_{i}+\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
-$\mathcal{C}_{i}^{-}(\alpha)$ $\coloneqq\widehat{\theta}_{i}-\Phi^{-1}(1-\alpha)\left[\bm{L}_{L\hat{z}}^{\dagger}\right]_{ii}^{1/2},$
-where $\Phi$ is the cdf for the standard normal distribution. Corollary 1 implies the following proposition.
+Instate the assumptions of Theorem 1.
+Consider $\mathrm{MRP\text{-}MLE}$ with $n_{\mathrm{split}}$ random splits with fixed $m$ and $p$.
+Suppose that the limits in (8) exist.
+Then as $n\rightarrow\infty$,
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right]\left(\bm{H}^{\infty}\right)^{\dagger}\right).$ (10)
+Theorem 3 provides an asymptotic result for inference of $\mathrm{MRP\text{-}MLE}$ when $m$ and $p$ are fixed and $n\rightarrow\infty$.
+In particular, it reveals the decrease in asymptotic covariance of $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ from $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ to $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ as $n_{\mathrm{split}}$ goes from $1$ to $\infty$.
+In what follows, we first connect this result with the asymptotic normality of $\mathrm{WP\text{-}MLE}$, which takes all possible item-item pairs with overlap instead of doing random pairing.
+Then we quantify the asymptotic variance of $\mathrm{MRP\text{-}MLE}$ in a special instance to see how much using multiple random splitting helps.
+Asymptotic normality of weighted pseudo MLE.
+Pseudo MLE (Zwi, 95) is a method for item parameter estimation similar to our approach $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$.
+Instead of random pairing, pseudo MLE forms all possible item-item pairs with overlaps.
+For the non-asymptotic analysis, the lack of independence between comparisons induced by the overlaps is clumsy.
+However, here we show that a variant of pseudo MLE is asymptotically normal and relates to our method $\mathrm{MRP\text{-}MLE}$.
+Now we formally introduce the weighted pseudo MLE, denoted as $\mathrm{WP\text{-}MLE}$.
+Let $m_{t}$ be the total number of responses of user $t$ and $\widetilde{m}_{t}$ be the largest even number smaller or equal to $m_{t}$.
+Consider the negative log-likelihood functions
+$\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j:\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right]$
+and
+$\mathcal{L}_{\mathrm{WP}}(\bm{\theta})\coloneqq\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
+Then $\mathrm{WP\text{-}MLE}$ is defined as
+$\widehat{\bm{\theta}}_{\mathrm{WP}}=\arg\min_{\bm{\theta}\in\mathbb{R}^{m},\bm{\theta}^{\top}\bm{1}_{m}=0}\mathcal{L}_{\mathrm{WP}}(\bm{\theta}).$ (11)
+Similar to $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$, we also use the notation $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ when appropriate.
+The intuition behind this reweighting is to account for the different numbers of responses for each user.
+It can be easily shown that $\mathbb{E}_{\mathrm{r}}\mathcal{L}_{k}(\bm{\theta})=\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$.
+Similar to the result for $\mathrm{MRP\text{-}MLE}$, we have the following asymptotic normality result for $\mathrm{WP\text{-}MLE}$.
+The proof is deferred to Section D.2.
+Theorem 4.
+Instate the assumptions of Theorem 1.
+Consider $\mathrm{WP\text{-}MLE}$ with fixed $m$ and $p$.
+Suppose that the limits in (8) exist.
+In addition, assume that for every fixed $\bm{\theta}$ obeying $\|\bm{\theta}-\bm{\theta}^{\star}\|_{\infty}\leq 10$, the following limit exists:
+$\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
+Then as $n\rightarrow\infty$,
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}\left(\bm{H}^{\infty}\right)^{\dagger}\right).$ (12)
+This theorem establishes the asymptotic normality of $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$.
+We observe that the asymptotic covariance of $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ is equal to the asymptotic covariance $\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}$ as $n_{\mathrm{split}}$ tends to infinity.
+This connects $\mathrm{WP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$, showing that they are asymptotically equivalent in distribution when $n\rightarrow\infty$ and $n_{\mathrm{split}}\rightarrow\infty$.
+Quantifying the shift of asymptotic covariance in $\mathrm{MRP\text{-}MLE}$.
+We have shown in (10) that the asymptotic covariance of $\mathrm{MRP\text{-}MLE}$ goes from $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ when $n_{\mathrm{split}}=1$ to $(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}$ when $n_{\mathrm{split}}\rightarrow\infty$.
+We have also established a qualified comparison in (9) that shows
+$(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}\preceq(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{same}}^{\infty}(\bm{H}^{\infty})^{\dagger}.$
+However, $\bm{V}_{\mathrm{same}}^{\infty}$ and $\bm{V}_{\mathrm{diff}}^{\infty}$ are not explicit.
+Here we make a quantified illustration in a special case to better understand this shift in asymptotic covariance.
+We make a few simplifications to make it straightforward.
+First, we suppose that the user parameters $\zeta_{t}^{\star}$ are independently drawn from a distribution $\pi$, and we denote the expectation with respect to the random user parameter with $\mathbb{E}_{\mathrm{u}}$.
+Second, we set the item parameter to be $\bm{\theta}^{\star}=\bm{0}_{m}$.
+Third, we assume the sampling model where each user response to $mp$ items uniformly at random, for some even integer $mp$.
+We also need a constant $\beta$, which is a scalar defined by
+$\beta\coloneqq\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{1}^{\star}}}{(e^{\zeta_{1}^{\star}}+1)^{2}}\right].$ (13)
+Note that $\zeta_{1}^{\star}$ in this definition can be replaced by $\zeta_{t}^{\star}$ for any $t$.
+In this special setting, we have the following result that quantifies the shift of asymptotic covariance for different $n_{\mathrm{split}}$.
+This proposition is special case for Theorem 3.
+The proof is deferred to Section D.3.
 Proposition 5.
-Suppose the statement in Theorem 3 holds. Then for any $(i,j)$, we have
-$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\geq\mathcal{C}_{i}^{+}(\alpha)\right]$ $=\alpha,$ (8a)
-$\lim_{np\rightarrow\infty}\mathbb{P}\left[\theta_{i}^{\star}\leq\mathcal{C}_{i}^{-}(\alpha)\right]$ $=\alpha,$ (8b)
-and
-$\lim_{np\rightarrow\infty}\mathbb{P}\left[\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\right]=\alpha.$ (9)
-This proposition can be extended to confidence intervals for any linear combination of $\bm{\theta}^{\star}$, such as $\theta_{i}^{\star}-\theta_{j}^{\star}$.
-These confidence intervals naturally enable hypothesis testings such as
-$\mathcal{H}_{0}:\theta_{j}^{\star}\geq\theta_{i}^{\star},\qquad\mathcal{H}_{1}:\theta_{j}^{\star}<\theta_{i}^{\star}$
-and
-$\mathcal{H}_{0}:\text{item $i$ is a top-$K$ item},\qquad\mathcal{H}_{1}:\text{item $i$ is not a top-$K$ item}.$
-Various other inference tasks have been considered in the ranking literature [LFL23][FLWY22, FLWY23].
-We expect our general distributional characterization in Theorem 3 to be readily applied for these tasks.
-4 Analysis
-In this section, we present the main steps in the analysis to obtain theoretical results in the previous section.
-Section 4.1 gives a complete argument on the reduction to the BTL model we alluded in Section 2.2, Section 4.2 provides the analysis of the $\ell_{\infty}$ error, and Section 4.3 provides the analysis of both the $\ell_{2}$ error and the asymptotic distribution as they share similar components in the analysis.
-4.1 Reduction to Bradley-Terry-Luce model
-A key idea in $\mathsf{RP\text{-}MLE}$ is the randomized pairing in Steps 1 and 2 of Algorithm 1.
+Instate the assumptions of Theorem 1. For fixed $m$ and $p$, as $n\rightarrow\infty$,
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\left(\frac{1}{n_{\mathrm{split}}}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\cdot\frac{mp}{2(mp-1)}\right)\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right);$ (14)
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\cdot\frac{mp}{2(mp-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right).$
+This proposition shows that in this special instance, the norm of the asymptotic covariance roughly scales as
+$\frac{1}{n_{\mathrm{split}}}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\cdot\frac{mp}{2(mp-1)}.$
+This equals $1$ when $n_{\mathrm{split}}=1$ and goes to mp/(2mp2)mp/(2mp-2) when $n_{\mathrm{split}}\rightarrow\infty$.
+The use of multiple random splitting in $\mathrm{MRP\text{-}MLE}$ or $\mathrm{WP\text{-}MLE}$ can shrink the asymptotic covariance of the estimators by a factor of mp/(2mp2)mp/(2mp-2).
+4 Experiment
+In this section, we demonstrate the empirical performance of $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ using both simulated and real data.
+4.1 Simulations
+We use simulated data to validate our theoretical results and compare our estimators with existing ones for the Rasch model.
+The data generating process follows the model specified in Section 2.1.
+Unless specified otherwise, in each trial, the user and item parameters are randomly drawn from
+$\widetilde{\bm{\zeta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{n}),\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{m}).$
+Afterwards, $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ is computed by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean.
+4.1.1 $\ell_{\infty}$ estimation error
+We investigate the $\ell_{\infty}$ estimation error with the following goals:
+  1. 1.
+    We validate the theoretical result in $\ell_{\infty}$ estimation error of $\mathrm{RP\text{-}MLE}$ in Theorem 1.
+  2. 2.
+    We show how much advantage the $\mathrm{MRP\text{-}MLE}$ brings through multiple runs of data splitting.
+  3. 3.
+    We compare our methods with existing comparison-based algorithms, including the case where $\kappa_{1},\kappa_{2}$ are large.
+Validating the theoretical result.
+Theorem 1 tells us that the $\ell_{\infty}$ error scales as $1/\sqrt{np}$.
+Figure 1 shows that $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ exhibits a near-linear relationship with respect to both $1/\sqrt{n}$ and $1/\sqrt{p}$, which is consistent with our theoretical predictions.
+(a) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{n}$. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $10000$ to $40000$.
+(b) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{p}$. The parameter is chosen to be $m=50,n=10000$ and $p$ varies from $1/9$ to $1$.
+Figure 1: Estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ of $\mathrm{RP\text{-}MLE}$ with varying $n$ and $p$. Each point represents the average of 1000 trials.
+Multiple runs in $\mathrm{MRP\text{-}MLE}$.
+As we have discussed after Theorem 1, the random data splitting could incur a small loss of information.
+We have introduced a remedy $\mathrm{MRP\text{-}MLE}$ (Algorithm 2) to address this by averaging over multiple runs with independent data splitting.
+Moreover, in Proposition 5 we have a quantitative characterization of the improvement in $\ell_{2}$ error achieved through multiple data splittings.
+Figure 2a shows that by averaging over more runs of data splittings, $\mathrm{MRP\text{-}MLE}$ achieves an improved $\ell_{\infty}$ estimation error that improves over PMLE and is close to $\mathrm{WP\text{-}MLE}$.
+In addition, we observe in Figure 2b that the improvement in squared $\ell_{2}$ error scales linearly with $1/n_{\mathrm{split}}$, consistent with the theoretical findings in Proposition 5.
+(a) The $\ell_{\infty}$ estimation error of $\mathrm{MRP\text{-}MLE}$ v.s.
+$n_{\mathrm{split}}$.
+The dash-dotted and dashed lines are the performance of PMLE and $\mathrm{WP\text{-}MLE}$, respectively.
+(b) The squared error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|^{2}$ of $\mathrm{MRP\text{-}MLE}$ v.s.
+$1/n_{\mathrm{split}}$ with varying $n_{\mathrm{split}}$.
+The dash-dotted and dashed lines are the performance of PMLE and $\mathrm{WP\text{-}MLE}$, respectively.
+Figure 2: Estimation error of $\mathrm{MRP\text{-}MLE}$ with varying number of data splittings.
+For each trial, we record $\|\frac{1}{k}\sum_{i=1}^{k}\widehat{\bm{\theta}}_{(i)}-\bm{\theta}^{\star}\|$ for $k=1,\ldots,100$.
+The parameters are chosen to be $m=50,p=0.2,n=10000$.
+The latent scores are all 0 and the each user is assigned with mpmp item uniformly-at-random.
+Each point is averaged over 1000 trials.
+Comparison with existing estimators.
+We compare our algorithms with two other comparison-based algorithms: the pseudo MLE (PMLE) and the spectral method from NZ (23).
+In Figure 4, We can see that the performance of $\mathrm{MRP\text{-}MLE}$ is comparable to PMLE and slightly outperforms the spectral method.
+Our proposed algorithm not only offers stronger theoretical guarantees but also demonstrates competitive practical performance.
+Performance with large $\kappa_{1},\kappa_{2}$.
+While we assume $\kappa=\max\{\kappa_{1},\kappa_{2}\}=O(1)$ is most of this article, scenarios with large $\kappa$ can be practically relevant.
+To evaluate the performance in such cases, we compare the $\ell_{\infty}$ error of different methods under different condition numbers.
+For a fixed $\kappa$, we draw the user and item parameters as
+$\widetilde{\bm{\zeta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))$
+and compute $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to have zero mean.
+Figure 4 illustrates the performance of different estimators as $\kappa$ varies.
+The MLE-based approaches including $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ achieve better $\ell_{\infty}$ error than the spectral method when $\kappa$ is large.
+Figure 3: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s.
+$n$ using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$, and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings.
+The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $5000$ to $20000$.
+The result is averaged over 1000 trials.
+Figure 4: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s.
+$\log(\kappa)$ using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$, and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings.
+The parameter is chosen to be $m=50,p=0.1,n=20000$ and $\kappa$ varies from $1$ to $e^{10}$.
+The result is averaged over 1000 trials.
+4.1.2 Top-$K$ recovery
+We investigate the performance of different algorithms in top-$K$ recovery.
+Set $\theta_{i}^{\star}=(1-K/m)\Delta_{K}$ for $i\leq K$ and $\theta_{i}^{\star}=(-K/m)\Delta_{K}$ otherwise.
+For any estimator $\widehat{\bm{\theta}}$, we define top-$K$ recovery rate to be
+$\frac{1}{K}\left|\left\{i\leq K:i\in A_{K}\right\}\right|,$
+where $A_{K}$ is an arbitrary $K$-element set such that $\widehat{\theta}_{i}\geq\widehat{\theta}_{j}$ for any $i\in A_{K},j\notin A_{K}$.
+We compare the top-$K$ recovery rate of PMLE and the spectral method in NZ (23) with $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ in Figure 6.
+The recovery rate of PMLE, spectral method and $\mathrm{MRP\text{-}MLE}$ is similar, indicating again that our algorithm performs well in practice.
+Figure 5: Top-$K$ recovery rate using spectral method, PMLE, $\mathrm{RP\text{-}MLE}$  and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings.
+The parameter is chosen to be $m=10000,m=50,p=0.1,K=5$ and $\Delta_{K}$ varies from 0.1 to 0.7.
+The result is averaged over 1000 trials.
+Figure 6: The relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ v.s.
+$n$ for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$.
+The parameter is chosen to be $p=0.1$, $n$ varies from $10000$ to $50000$ and $m=n/500$.
+The result is averaged over 1000 trials.
+4.1.3 Refined $\ell_{2}$ estimation error
+In Theorem 4 we have shown that the $\ell_{2}$ error concentrate around $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ for $\widetilde{z}\in\left\{\widehat{z},z\right\}$.
+In each trial we compute the following quantity
+$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}$
+for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$.
+This measures the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$.
+In Figure 6 we consider the regime where $p$ and $n/m$ is fixed.
+In this case, Theorem 4 implies that
+$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}\lesssim\frac{\sqrt{\frac{1}{np}}+\frac{\sqrt{m}}{np}}{\sqrt{\frac{m}{np}}}\lesssim\frac{1}{\sqrt{n}}.$
+[…]
+4.1.4 Confidence intervals for $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$
+Figure 7: Theoretical coverage rate $1-\alpha$ (dashed line) and empirical coverage rate $1-\widehat{\alpha}$ of the two-sided confidence intervals $\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$.
+The first 10 points on the left in each level are computed with $\mathrm{MRP\text{-}MLE}$ with $n_{\mathrm{split}}$ vaying from 10 to 100.
+The rightmost point on each level is computed with $\mathrm{WP\text{-}MLE}$.
+The parameters are set to be $n=10000,m=20,p=0.5$.
+The confidence interval is $1-\alpha=0.8,0.9,0.95,0.99$.
+Each point is averaged over $10000$ trials.
+The asymptotic normality results in Theorem 3 and 4 allow us to construct confidence intervals for $\theta_{i}^{\star}$ with $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$.
+For large enough $n_{\mathrm{split}}$, we can approximate the variance of both estimators with $\frac{1}{n}(\bm{H}^{\infty})^{\dagger}\bm{V}_{\mathrm{diff}}^{\infty}(\bm{H}^{\infty})^{\dagger}$, where $\bm{V}_{\mathrm{diff}}^{\infty}$ and $\bm{H}^{\infty}$ are estimated using the following plug-in estimates:
+[…]
+[…]
+For a given confidence level $1-\alpha$, the confidence interval $\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$ is
+$\mathcal{C}_{i}^{-}(\alpha/2)$ $\coloneqq\widehat{\theta}_{i}-z_{1-\alpha/2}\cdot\left[\frac{1}{n}(\widehat{\bm{H}}^{\infty})^{\dagger}\widehat{\bm{V}}_{\mathrm{diff}}^{\infty}(\widehat{\bm{H}}^{\infty})^{\dagger}\right]_{ii}^{1/2},$
+$\mathcal{C}_{i}^{+}(\alpha/2)$ $\coloneqq\widehat{\theta}_{i}+z_{1-\alpha/2}\cdot\left[\frac{1}{n}(\widehat{\bm{H}}^{\infty})^{\dagger}\widehat{\bm{V}}_{\mathrm{diff}}^{\infty}(\widehat{\bm{H}}^{\infty})^{\dagger}\right]_{ii}^{1/2}.$
+In Figure 7, we compare the empirical coverage rate
+$1-\widehat{\alpha}\coloneqq\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\{\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\}.$
+of these two-sided confidence intervals with the theoretical ones.
+We can see that when $n_{\mathrm{split}}$ is reasonably large, the empirical coverage rate matches well with the theoretical coverage rate for both $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$.
+4.2 LSAT dataset
+We study a real-world dataset (LSAT) on the Law School Admissions Test from DBL (70).
+LSAT has full observation of 1000 people answering 5 problems, with each person-item pair recording whether the answer was correct.
+The second row in Table 1 lists how many people answer each problem correctly.
+From the first look, Problem 3 appears to be the hardest question.
+We proceed to quantify the hardness of these problems under the Rasch model and infer how confident we are in claiming it is the hardest.
+Using $\mathrm{MRP\text{-}MLE}$  we compute a latent score estimate and construct two-sided confidence intervals at significance level $\alpha=0.01$ for each coordinate, following the methodology introduced in Section 4.1.4.
+The result is summarized in Table 1, where higher latent score correspond to greater difficulty.
+The estimated parameters align inversely with the total number of correct answers.
+Notably, the lower bound of the confidence interval for $\theta_{3}^{\star}$ is larger than the upper bounds of the confidence intervals of $\theta_{1}^{\star},\theta_{2}^{\star},\theta_{4}^{\star}$, and $\theta_{5}^{\star}$.
+With Bonferroni correction, we can conclude that with 95% confidence Problem 3 is the most difficult problem in this dataset.
+Problem 1 2 3 4 5
+Total correct 924 709 553 763 870
+$\theta$ estimate -1.2824 0.4511 1.2800 0.1926 -0.6413
+CI lower bound -1.5579 0.2696 1.0958 0.0017 -0.8711
+CI upper bound -1.0069 0.6327 1.4641 0.3834 -0.4116
+Table 1: Latent score estimate calculated using $\mathrm{MRP\text{-}MLE}$ with 20 data splittings and confidence interval calculated with the construction introduced in Section 4.1.4.
+Higher latent score here means higher difficulty.
+The significance level is chosen to be $\alpha=0.01$ for each coordinate.
+Now we assume Problem 3 is the top-$1$ item in latent score and investigate the top-$1$ recovery rate of different algorithms on LSAT under incomplete observation.
+In each trial we randomly select $\widetilde{n}$ people, and for each of them we randomly select their outcome on $\widetilde{m}$ problems.
+We then estimate $\bm{\theta}^{\star}$ using the subsampled data with different methods and compare the proportion of trials where the top-$1$ item is correctly identified.
+In Figure 8, we see results similar to the simulation.
+Our algorithm $\mathrm{MRP\text{-}MLE}$ has a similar recovery rate compared to PMLE and spectral method.
+Figure 8: Top-1 recovery rate using Spectral method, PMLE, $\mathrm{RP\text{-}MLE}$  and $\mathrm{MRP\text{-}MLE}$ using 20 data splittings.
+The parameters are chosen to be $\widetilde{m}=4$ and $\widetilde{n}$ varies from 100 to 200.
+The result is averaged over 10000 trials.
+5 Discussion
+In this paper, we propose two new likelihood-based estimators $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ for item parameter estimation in the Rasch model.
+Both enjoy optimal finite sample estimation guarantee and asymptotic normality that allows for tight uncertainty quantification.
+All this is achieved even when the user-item response data are extremely sparse (cf.
+NZ (23)).
+Below, we identify several questions that are interesting for further investigation:
+  • Does PMLE or CMLE achieve optimal theoretical guarantee?
+In our experiments, pseudo MLE has shown a similar performance to $\mathrm{MRP\text{-}MLE}$.
+This naturally leads to the question of whether PMLE can enjoy the same theoretical guarantee.
+This is relevant to our work because our methods can be viewed as a modification of pseudo MLE by incorporating random disjoint pairing to decouple statistical dependency among paired $Y_{ij}$’s.
+It remains unclear whether such dependency is a fundamental bottleneck.
+On the other hand, conditional MLE is another popular method used in practice.
+In Section A.1 of the supplement we will mention that CMLE can also be viewed as a reduction to a less studied item-only model.
+This reduction is more complicated for analysis as it constructs item tuples rather than item pairs.
+It would be interesting to know whether we can transfer the techniques we have used here to develop a non-asymptotic analysis for CMLE.
+  • Extending random pairing to other models in IRT.
+Some IRT models parameterize the latent score of users and items differently from the Rasch model.
+For instance, consider the two-parameter logistic model (2PL) with discrimination parameter on the users.
+It assumes that $X_{ti}$, the response of user $t$ to item $i$, follows the law
+    $\mathbb{P}[X_{ti}=1]=\frac{1}{1+\exp(a_{t}^{\star}(\zeta_{t}^{\star}-\theta_{i}^{\star}))},$
+    where $\bm{\theta}^{\star}$ is the latent scores of the items , $\bm{\zeta}^{\star}$ is the latent scores of the users, while $\bm{a}^{\star}$ is the discrimination parameters.
+Unlike the Rasch model, in the 2PL model,
+    $\mathbb{P}[X_{ti}>X_{tj}\mid X_{ti}\neq X_{tj}]=\frac{\exp(a_{t}^{\star}\theta_{i}^{\star})}{\exp(a_{t}^{\star}\theta_{i}^{\star})+\exp(a_{t}^{\star}\theta_{j}^{\star})}$
+    is not independent of the user discrimination parameter $\bm{a}^{\star}$.
+Therefore the reduction to the BTL model is no longer true in this case.
+However, one could employ a partially-Bayesian approach by putting a prior on $\bm{\alpha}^{\star}$ and maximize this marginally likelihood, which is a function of $\theta_{i}^{\star}-\theta_{j}^{\star}$ independent of $\bm{\zeta}$.
+It is interesting and non-trivial to extend the idea of random pairing to the 2PL model.
+  • Extension to joint estimation of user and item parameters.
+It is sometimes of interest to estimate both the user and the item parameters.
+We expect our method $\mathrm{MRP\text{-}MLE}$ continues to work with slight modifications.
+In a high level, the idea is to estimate the mean-shifted parameters $\bm{\theta}^{\star}-(1/m)\bm{1}_{m}\bm{1}_{m}^{\top}\bm{\theta}^{\star}$ and $\bm{\zeta}^{\star}-(1/m)\bm{1}_{m}\bm{1}_{m}^{\top}\bm{\zeta}^{\star}$ using $\mathrm{MRP\text{-}MLE}$ twice.
+In the end, one estimates the difference in the means using MLE over the comparison outcomes.
+We leave the detailed investigation to future work.
+Appendix A Analysis
+In this section, we present the main steps to obtain theoretical results in the previous section.
+Section A.1 provides a complete argument on the reduction to the BTL model we mentioned in Section 2.2.
+Section A.2 provides the analysis of the $\ell_{\infty}$ error, Section A.3 provides the analysis of the non-asymptotic expansion, and Section A.4 sketches the proof of the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$.
+A.1 Reduction to Bradley-Terry-Luce model
+A key component in $\mathrm{RP\text{-}MLE}$ is the random pairing in Steps 1 and 2 of Algorithm 1.
 It compiles the user-item responses $\bm{X}$ to item-item comparisons $\bm{Y}$.
@@
 The following fact provides the distribution of $Y_{ij}^{t}$ conditional on $L_{ij}^{t}=1$.
-We defer its proof to Section A.
+We defer its proof to Section E.1.
 Fact 1.
@@
 $\mathbb{P}[X_{ti}<X_{tj}\mid L_{ij}^{t}=1]=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}},$
-and
 $\mathbb{P}[L_{ij}^{t}=1]\geq\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}.$ (15)
@@
   2. 2.
-    Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems. (Step 1(a) of Algorithm 1)
+    Randomly split the $m_{t}$ problems taken by user $t$ into $\lfloor m_{t}/2\rfloor$ pairs of problems. (Step 1(a) of Algorithm 1)
   3. 3.
-    For all $(t,i,j)$, items $i$ and $j$ are compared by user $t$ if $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}=1$.
+    For all $(i,j,t)$, items $i$ and $j$ are compared by user $t$ if $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}=1$.
   4. 4.
@@
 This justifies that (2) is truly the likelihood function of the BTL model conditional on the comparison graph $\mathcal{E}_{Y}$.
-4.2 Analysis of $\ell_{\infty}$ error
-We have seen that analyzing $\mathsf{RP\text{-}MLE}$ under the Rasch model can be reduced to analyzing the MLE under the BTL model.
-This reduction allows us to invoke the general theory of MLE in the BTL model established in the recent work [YCOM24].
+In addition, we would like to comment on another popular method conditional MLE, which can also be viewed as a reduction to a item-only model.
+The CMLE maximizes the likelihood conditioned on total number of positive responses.
+It can be computed that
+$\mathbb{P}\left[X_{ti_{1}},X_{ti_{2}},\cdots,X_{ti_{m_{t}}}\mid\sum_{l=1}^{m_{t}}X_{ti_{l}}=k\right]=\frac{\prod_{l=1}^{m_{t}}e^{\theta_{i_{l}}^{\star}}\mathds{1}\{X_{ti_{l}}=1\}}{\sum_{\bm{\alpha}\in\{0,1\}^{m_{t}}}\prod_{l=1}^{m_{t}}e^{\theta_{i_{l}}^{\star}}\mathds{1}\{\alpha_{l}=1\}}.$
+It is easy to see that the conditional probability is not dependent on the user parameters $\bm{\zeta}^{\star}$.
+However, the model CMLE reduces to is less studied than the BTL model, especially in the setting of sparse observations.
+A.2 Analysis for entrywise error bound
+We have seen that analyzing $\mathrm{RP\text{-}MLE}$ under the Rasch model can be reduced to analyzing the MLE under the BTL model.
+This reduction allows us to invoke the result in the recent work [47] established for MLE in the BTL model with a general comparison graph.
 To facilitate the presentation, we introduce the necessary notation.
@@
 $\bm{L}_{L}\coloneqq\sum_{i,j:i>j}L_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
-The following lemma adapts Theorem 3 of the recent work [YCOM24] to our setting.
-Lemma 1 (Theorem 3 in [YCOM24]).
+The following lemma adapts Theorem 3 of the recent work [47] to our setting.
+Lemma 1 (Theorem 3 in [47]).
 Assume that $\mathcal{G}_{Y}$ is connected, and that
-$[\lambda_{m-1}(\bm{L}_{L})]^{5}\geq C_{1}\kappa_{1}^{4}(d_{\max})^{4}\log^{2}(n)$ (11)
+$\lambda_{m-1}(\bm{L}_{L})]^{5}\geq C_{1}\kappa_{1}^{4}(d_{\max})^{4}\log^{2}(n)$ (16)
 for some large enough constant $C_{1}>0$. Then with probability at least $1-n^{-10}$, we have
@@
 Suppose $mp\geq 2$, $np\geq C\kappa_{2}^{2}\log(n)$ for some large enough constant $C$, and $m\leq n^{\alpha}$ for some constant $\alpha>0$. With probability at least $1-10n^{-10}$, we have
-$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L})\leq\lambda_{1}(\bm{L}_{L})\leq 3np$ (13)
-and
+$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L})\leq\lambda_{1}(\bm{L}_{L})\leq 3np,$ (18)
 $\frac{np}{16\kappa_{1}\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{Lz})\leq\lambda_{1}(\bm{L}_{Lz})\leq np.$ (19)
-4.2.1 Proof of Theorem 1
+A.2.1 Proof of Theorem 1
 Now we are ready to prove Theorem 1.
-We focus on analyzing $\mathsf{RP\text{-}MLE}$, as the analysis of $\mathsf{MRP\text{-}MLE}$ follows immediately from the union bound of the different data splitting and the triangular inequality:
+We focus on analyzing $\mathrm{RP\text{-}MLE}$, as the analysis of $\mathrm{MRP\text{-}MLE}$ follows immediately from the union bound of the different data splitting and the triangular inequality:
 $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}\leq\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\|\widehat{\bm{\theta}}^{(i)}-\bm{\theta}^{\star}\|_{\infty}.$
@@
 $np\geq\frac{16C_{2}^{2}\kappa_{1}^{2}\kappa_{2}\log(n)}{\Delta_{K}^{2}}.$
-4.3 Analysis for refined $\ell_{2}$ error and uncertainty quantification
-To make the main text concise, we sketch the key ideas that lead to both Theorems 2 and 3.
-We start with some necessary notation.
-For convenience, for any $(i,j)$, renumber the set of comparisons $\{Y_{ji}^{t}\}_{t:L_{ij}^{t}=1}$ as $\{Y_{ji}^{(l)}\}_{l=1,\ldots,L_{ij}^{t}}$ in an arbitrary order.
-Let $\epsilon_{ij}^{(l)}\coloneqq Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$ where $\sigma(x)=e^{x}/(1+e^{x})$ is the sigmoid function.
-Let $\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ (repeat $L_{ij}$ times for edge $(i,j)$), where $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$ is the total number of observed effective comparisons in $\mathcal{G}_{Y}$.
-Define
-$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{(l)}/\sqrt{z_{ij}},\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y},l=1,\ldots,L_{ij}^{t}}\in\mathbb{R}^{L_{\mathrm{total}}}.$
-The key step in establishing Theorems 2 and 3 is the following refined characterization of the estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$.
-Lemma 4.
-Instate the assumptions of Theorem 1. With probability at least $1-O(n^{-10})$, the estimator $\widehat{\bm{\theta}}$ given by the Algorithm (1) can be written as
-$\widehat{\bm{\theta}}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r},$ (15)
-where $\bm{r}\in\mathbb{R}^{m}$ is a random vector obeying $\|\bm{r}\|_{\infty}\leq C\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C>0$.
-As the residual term $\bm{r}$ has small magnitude, we may analyze the properties of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ by focusing on the leading term $-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$.
-Proof outline of Theorem 2.
-Using the approximation $\widehat{\bm{\theta}}-\bm{\theta}^{\star}\approx-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$, we can turn our attention to the concentration of
-$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|^{2}\approx\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}=\widehat{\bm{\epsilon}}^{\top}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}.$
-This is a quadratic form of sub-Gaussian random variables.
-The Hanson-Wright inequality tells us that $\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}$ concentrates around its expectation—$\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})$.
-The complete proof is deferred to Section D for Theorem 2.
-Proof outline of Theorem 3.
-With Lemma 4 in hand, we can expand $\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ to reach
-$\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ $\approx-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$
-$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
-The main component of the last expression is a sum of bounded random variables $\bm{u}_{ij}^{(l)}$ that are independent conditional on $\mathcal{G}_{Y}$. This sum has zero mean and variance
-$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
-Therefore applying multivariate CLT, we obtain
-$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
-See the complete proof in Section E.
-4.4 Proof outline of Lemma 4
-Due to the importance and usefulness of Lemma 4, we provide a sketch of its proof and leave the full one to Section C.
-The proof is inspired by the proof of Theorem 1 in [Che23], which analyzes MLE via the trajectory of the preconditioned gradient descent (PGD) dynamic starting from ground truth.
+A.3 Analysis for non-asymptotic expansion
+To make the main text concise, we provide a sketch of the proof of Theorem 2 and leave the full one to Section C.1.
+The proof is inspired by the proof of Theorem 1 in [5], which analyzes MLE via the trajectory of the preconditioned gradient descent (PGD) dynamic starting from ground truth.
 More precisely, letting $\bm{\theta}^{0}=\bm{\theta}^{\star}$, we consider the PGD iterates defined by
@@
 where $\eta>0$ is the step size of PGD.
-[Che23] shows that this dynamic converges to $\widehat{\bm{\theta}}$.
-We proceed one step further by establishing precise distributional characterization of $\hat{\bm{\theta}}$ via analyzing PGD.
+[5] shows that this dynamic converges to $\widehat{\bm{\theta}}$.
+We proceed one step further by establishing precise distributional characterization of $\widehat{\bm{\theta}}$ via analyzing PGD.
 With Taylor expansion, the gradient can be decomposed into
-$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
+$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}(\bm{\theta}^{t}-\bm{\theta}^{\star})-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
 where $\bm{r}^{t}$ is a residual vector with small magnitude. Then the PGD update becomes
 $\bm{\theta}^{t+1}-\bm{\theta}^{\star}=\left(1-\eta\right)(\bm{\theta}^{t}-\bm{\theta}^{\star})-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right).$
-We establish Lemma 4 by solving this recursive relation.
+We establish Theorem 2 by solving this recursive relation.
 More specifically, as $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ does not depend on $t$ and $\|\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\|_{\infty}$ can be controlled for each step $t$, taking $t\rightarrow\infty$, we see that
@@
 for some residual term $\bm{r}$ that is well controlled in $\ell_{\infty}$ norm.
-5 Experiment
-In this section, we demonstrate the empirical performance of $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$ using both simulated and real data.
-5.1 Simulations
-We begin by using simulated data to validate our theoretical results and compare our estimators with existing ones for the Rasch model.
-The data generating process follows the model specified in Section 2.1. Unless specified otherwise, in each trial, the user and item parameters are randomly drawn from
-$\widetilde{\bm{\zeta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{n}),\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathcal{N}(0,\bm{I}_{m}).$
-Afterwards, $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ is computed by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean.
-5.1.1 $\ell_{\infty}$ estimation error
-Here, we investigate the $\ell_{\infty}$ estimation error.
-First, we validate the theoretical result in $\ell_{\infty}$ estimation error of $\mathsf{RP\text{-}MLE}$.
-Second, we show how much advantage the $\mathsf{MRP\text{-}MLE}$ brings through multiple runs of data splitting.
-Lastly, we compare our methods with existing comparison-based algorithms, including the case where $\kappa_{1},\kappa_{2}$ are large.
-Validating the theoretical result.
-Theorem 1 tells us that the $\ell_{\infty}$ error scales as $1/\sqrt{np}$.
-As shown in Figure 1, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ is very close to being linear w.r.t.
-both $1/\sqrt{n}$ and $1/\sqrt{p}$.
-This is consistent with our theoretical predictions.
-(a) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ with v.s.
-$1/\sqrt{n}$.
-The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $10000$ to $40000$.
-Each point represents the average of 1000 trials.
-(b) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s.
-$1/\sqrt{p}$.
-The parameter is chosen to be $m=50,n=10000$ and $p$ varies from $1/9$ to $1$.
-Each point represents the average of 1000 trials.
-Figure 1: Estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ of $\mathsf{RP\text{-}MLE}$ with varying $n$ and $p$.
-Multiple runs in $\mathsf{MRP\text{-}MLE}$.
-As we have discussed after Theorem 1, the random data splitting could incur a small loss of information.
-We have provided a remedy $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by averaging over multiple runs with independent data splitting.
-In Figure 2, we show the empirical evidence of its effectiveness.
-For each trial, we run 50 independent data splitting and obtain the 50 estimates $\widehat{\bm{\theta}}^{(i)}$.
-We then report the $\ell_{\infty}$ error of the running average, i.e., $\|\frac{1}{k}\sum_{i=1}^{k}\widehat{\bm{\theta}}^{(i)}-\bm{\theta}^{\star}\|$ for $k=1,\ldots,50$.
-As a baseline for comparison, we also run pseudo MLE on the same datasets, which uses all possible pairs of comparison.
-We can see that by averaging over more runs of data splittings, $\mathsf{MRP\text{-}MLE}$ gives a better estimation.
-In particular, one can see that with sufficient number of independent data splittings, the average estimation error matches and even beats pseudo MLE.
-Figure 2: The $\ell_{\infty}$ estimation error of $\mathsf{MRP\text{-}MLE}$ with different numbers of data splitting.
-The dashed line is the performance of Pseudo MLE.
-The parameters are chosen to be $m=50,p=0.1,n=10000$.
-Each point is averaged over 1000 trials.
-Comparison with existing estimator.
-We then compare our algorithms with two other comparison-based algorithms: pseudo MLE (PMLE) and the spectral method in [NZ23].
-We plot the $\ell_{\infty}$ error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ against $n$ in Figure 4.
-We can see that all four algorithms follow a similar trend.
-The performance of $\mathsf{MRP\text{-}MLE}$ is on par with PMLE and better than the spectral method.
-Our proposed algorithm not only enjoys an improved theoretical guarantee but is also competitive in practice.
-Performance with large $\kappa_{1},\kappa_{2}$.
-While we assume $\kappa=\max\{\kappa_{1},\kappa_{2}\}=O(1)$ is most of this article, the scenario with large $\kappa$ can be practically relevant.
-Here we compare the $\ell_{\infty}$ error of different methods when the condition number is large.
-Fixing $\kappa$, we draw the user and item parameters with
-$\widetilde{\bm{\zeta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))\qquad\text{and}\qquad\widetilde{\bm{\theta}}^{\star}\sim\mathrm{Unif}(0,\log(\kappa))$
-and compute $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean.
-Figure 4 shows the performance of different estimators as $\kappa$ varies.
-The MLE-based approaches including $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ achieve better $\ell_{\infty}$ error than the spectral method when $\kappa$ is large.
-Figure 3: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s.
-$n$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings.
-The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $5000$ to $20000$.
-The result is averaged over 1000 trials.
-Figure 4: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s.
-$\log(\kappa)$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings.
-The parameter is chosen to be $m=50,p=0.1,n=20000$ and $\kappa$ varies from $1$ to $e^{10}$.
-The result is averaged over 1000 trials.
-5.1.2 Top-$K$ recovery
-In this part, we investigate the performance of different algorithms in top-$K$ recovery.
-We set $\theta_{i}^{\star}=(1-K/m)\Delta_{K}$ for $i\leq K$ and $\theta_{i}^{\star}=(-K/m)\Delta_{K}$ otherwise.
-For any estimator $\widehat{\bm{\theta}}$, we define top-$K$ recovery rate to be
-$\frac{1}{K}\left|\left\{i\leq K:i\in A_{K}\right\}\right|,$
-where $A_{K}$ is an arbitrary $K$-element set such that $\widehat{\theta}_{i}\geq\widehat{\theta}_{j}$ for any $i\in A_{K},j\notin A_{K}$.
-We compare the top-$K$ recovery rate of PMLE and the spectral method in [NZ23] with $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ in Figure 6.
-The recovery rate of PMLE, spectral method and $\mathsf{MRP\text{-}MLE}$ is similar, indicating again that our algorithm performs well in practice.
-Figure 5: Top-$K$ recovery rate using spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings.
-The parameter is chosen to be $m=10000,m=50,p=0.1,K=5$ and $\Delta_{K}$ varies from 0.1 to 0.7.
-The result is averaged over 1000 trials.
-Figure 6: The relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ v.s.
-$n$ for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$.
-The parameter is chosen to be $p=0.1$, $n$ varies from $10000$ to $50000$ and $m=n/500$.
-The result is averaged over 1000 trials.
-5.1.3 Refined $\ell_{2}$ estimation error
-In this part of the section, we show the empirical evidence of Theorem 2.
-In Theorem 2 we have shown that the $\ell_{2}$ error concentrate around $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ for $\widetilde{z}\in\left\{\widehat{z},z\right\}$.
-In each trial we compute the following quantity
-$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}$
-for both $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$.
-This measures the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$.
-In Figure 6 we consider the regime where $p$ and $n/m$ is fixed.
-In this case, Theorem 2 implies that
-$\frac{\left|\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|-\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}\right|}{\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}}\lesssim\frac{\sqrt{\frac{1}{np}}+\frac{\sqrt{m}}{np}}{\sqrt{\frac{m}{np}}}\lesssim\frac{1}{\sqrt{n}}.$
-In other words, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrate tightly around $\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})$.
-We can see that the deviation is very small between $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$.
-In both cases, the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ decreases as $n$ and $m$ increase as expected.
-5.1.4 Inference
-Lastly, we validate our inferential results in Section 3.3.
-Asymptotic normality.
-We first check the asymptotic normality presented in Theorem 3 and Corollary 1. For simplicity, we focus on the first coordinate. Theorem 3 and Corollary 1 have claimed that
-$\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1-\frac{1}{m}\right)$
-and
-$\frac{\widehat{\theta}_{1}-\theta_{1}^{\star}}{\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$
-For simplicity we set both $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ to be all zero vectors.
-In Figure 7, we present the QQ-plot comparing the generated instances and the theoretical distributions.
-In both cases, the spread of the term of interest matches the normality claimed in Theorem 3 and Corollary 1.
-Figure 7: Quantile-quantile plots of $\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}$ and $(\widehat{\theta}_{1}-\theta_{1}^{\star})/\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}$ v.s.
-zero-mean normal distributions.
-The estimate $\widehat{\bm{\theta}}$ is computed with $\mathsf{RP\text{-}MLE}$.
-The parameters are set to be $n=10000,m=10,p=0.5$.
-Data is collected over 10000 trials.
-Coverage of the confidence intervals.
-We compare the empirical coverage rate of the two-sided confidence intervals in Proposition 3 with the theoretical ones.
-Without loss of generality we consider the confidence interval for the first coordinate.
-In each trial, we compute the empirical coverage rate
-$1-\widehat{\alpha}=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\{\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\}$
-and report the average $1-\widehat{\alpha}$ over the trials.
-We do this for multiple $\alpha$’s and compare $1-\alpha$ with $1-\widehat{\alpha}$ in Table 1.
-The empirical coverage rate $1-\widehat{\alpha}$ is very close to the theoretical coverage rate $1-\alpha$.
-$1-\alpha$ 0.8 0.9 0.95 0.98 0.99 0.995 0.999
-$1-\widehat{\alpha}$ 0.801145 0.900635 0.95108 0.980395 0.99025 0.99523 0.99904
-Table 1: Theoretical coverage rate $1-\alpha$ and empirical coverage rate $1-\widehat{\alpha}$ of the two-sided confidence intervals $[\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$.
-The parameters are set to be $n=10000,m=20,p=0.1$.
-The result is averaged over $10000$ trials.
-5.2 LSAT real data
-We study a real dataset (LSAT) on the Law School Admissions Test from [DBL70].
-LSAT has full observation of 1000 people answering 5 problems.
-Foro each person-item pair it records whether the answer is correct or not.
-The second row in Table 2 lists how many people answer each problem correctly.
-From the first look, Problem 3 seems to be the hardest question.
-We proceed to quantify the hardness of these problems under the Rasch model and infer how confident we are in claiming it is the hardest.
-We use $\mathsf{RP\text{-}MLE}$ to get an estimate of the latent scores and calculate two-sided confidence intervals at level $\alpha=0.01$ for each coordinate with the framework in Section 3.3.
-See Table 2 for the result.
-Higher latent score here means higher difficulty.
-The estimated parameters agree with the total number of correct answers in reverse order.
-We can also see that the confidence interval lower bound of $\theta_{3}^{\star}$ is larger than the largest confidence interval upper bound of $\theta_{1}^{\star},\theta_{2}^{\star},\theta_{4}^{\star},\theta_{5}^{\star}$.
-With Bonferroni correction, we have a 0.95 confidence level to say Problem 3 is the most difficult problem.
-Problem 1 2 3 4 5
-Total correct 924 709 553 763 870
-$\theta$ estimate -1.2139 0.4093 1.1123 0.2977 -0.6053
-CI lower bound -1.5957 0.1449 0.8321 0.0234 -0.9050
-CI upper bound -0.8322 0.6736 1.3926 0.5720 -0.3057
-Table 2: Latent score estimate calculated using $\mathsf{RP\text{-}MLE}$ and confidence interval calculated with the framework in Section 3.3.
-The significance level is chosen to be $\alpha=0.01$ for each coordinate.
-Now we assume Problem 3 is the top-$1$ item and investigate the top-$1$ recovery rate of different algorithms on LSAT under incomplete observation.
-In each trial we randomly select $\widetilde{n}$ people, and for each of them we randomly select their outcome on $\widetilde{m}$ problems.
-We then estimate $\bm{\theta}^{\star}$ using the subsampled data with different methods and compare their top-1 recovery rate, i.e., the proportion of trials where the top-$1$ item is correctly identified.
-In Figure 8 we see results similar to the simulation.
-Our algorithm $\mathsf{MRP\text{-}MLE}$ has a similar recovery rate compared to PMLE and spectral method.
-Figure 8: Top-1 recovery rate using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings.
-The parameters are chosen to be $\widetilde{m}=4$ and $\widetilde{n}$ varies from 100 to 200.
-The result is averaged over 10000 trials.
-6 Discussion
-In this paper, we propose two new likelihood-based estimators $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ for item parameter estimation in the Rasch model.
-Both enjoy optimal finite sample estimation guarantee and $\mathsf{RP\text{-}MLE}$ is further amenable to tight uncertainty quantification.
-All this is achieved even when the user-item response data are extremely sparse (cf.
-[NZ23]).
-Below, we identify several questions that are interesting for further investigation:
-  • Does PMLE achieve optimal theoretical guarantee?
-In our experiments, pseudo MLE has shown a similar performance to $\mathsf{MRP\text{-}MLE}$.
-This naturally leads to the question of whether PMLE can enjoy the same theoretical guarantee.
-This is relevant to our work because our methods can be viewed as a modification of pseudo MLE by incorporating random disjoint pairing to decouple statistical dependency among paired $Y_{ij}$’s.
-It remains unclear whether such dependency is a fundamental bottleneck.
-  • Inference for multiple data splitting.
-In $\mathsf{MRP\text{-}MLE}$ (Algorithm 2), we average the estimates over multiple runs of data splitting to improve the empirical result over $\mathsf{RP\text{-}MLE}$.
-While this is effective for estimation, we do not have an immediate analog for inference as the estimations $\{\widehat{\bm{\theta}}^{(i)}\}_{i\in[n_{\mathrm{split}}]}$ are dependent on each other.
-It is an interesting question whether we can characterize the distribution of $\widehat{\bm{\theta}}$ in $\mathsf{MRP\text{-}MLE}$ and whether it leads to better inference, e.g., tighter confidence intervals.
-  • Extending random pairing to other models in IRT.
-Some IRT models parameterize the latent score of users and items differently from the Rasch model.
-For instance, the two-parameter logistic model (2PL) assumes that $X_{ti}$, the response of user $t$ to item $i$, follows the law
-    $\mathbb{P}[X_{ti}=1]=\frac{\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})}{1+\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})},$
-    where $\bm{\theta}^{\star}$ is the latent score of the items while $\bm{a}^{\star}$ and $\bm{b}^{\star}$ are two parameter vectors for the users. Unlike the Rasch model, in the 2PL model,
-    $\mathbb{P}[X_{ti}>X_{tj}\mid X_{ti}\neq X_{tj}]=\frac{\exp(b_{t}^{\star}\theta_{i}^{\star})}{\exp(b_{t}^{\star}\theta_{i}^{\star})+\exp(b_{t}^{\star}\theta_{j}^{\star})}$
-    is not independent of the user parameter $\bm{b}^{\star}$.
-Therefore reduction to the BTL model is no longer true in this case.
-It is interesting and non-trivial to extend the idea of random pairing to the 2PL model.
-Acknowledgement
-YY and CM were partially supported by the National Science Foundation via grant DMS-2311127.
-Appendix A Proof of Fact 1
-Expanding the probability of the random events, we have that
-$\mathbb{P}[X_{ti}<X_{tj}\mid X_{ti}\neq X_{tj}]$ $=\frac{\mathbb{P}[X_{ti}=0,X_{tj}=1]}{\mathbb{P}[X_{ti}=0,X_{tj}=1\text{\text{ or }}X_{ti}=1,X_{tj}=0]}$
-$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\Bigg{/}\left(\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\right)$
-$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{e^{\zeta_{t}^{\star}}(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}$
-$=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}.$
-Now consider $\mathbb{P}[X_{ti}\neq X_{tj}]$, we have
-$\mathbb{P}[X_{ti}\neq X_{tj}]$ $=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}$
-$=\frac{e^{\theta_{j}^{\star}-\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}}}{(1+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}})(1+e^{\theta_{j}^{\star}-\zeta_{t}^{\star}})}.$ (16)
-Let $f:[1/\kappa_{2},\kappa_{2}]^{2}\rightarrow\mathbb{R}$ defined by
-$f(a,b)\coloneqq\frac{a+b}{(1+a)(1+b)}.$
-Its partial derivatives are
-$\frac{\partial}{\partial a}f(a,b)=\frac{b^{2}-1}{(1+a)^{2}(1+b)^{2}}\quad\text{and}\quad\frac{\partial}{\partial b}f(a,b)=\frac{a^{2}-1}{(1+a)^{2}(1+b)^{2}}.$
-It is now easy to see that the minimum or maximum of $f$ can only happen if $(a,b)=(1,1)$ or $(a,b)\in\{1/\kappa_{2},\kappa_{2}\}^{2}$.
-After comparing the value of $f$ at these points, we conclude that $f$ achieves minimum at
-$f(1/\kappa_{2},1/\kappa_{2})=f(\kappa_{2},\kappa_{2})=\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}.$
-By the definition of $\kappa_{2}$, $|\theta_{l}^{\star}-\zeta_{t}^{\star}|\leq\log(\kappa_{2})$ for any $l\in[m]$. Then (16) fits the definition of $f$ and the proof is completed.
+A.4 Analysis for asymptotic normality
+The analysis of the asymptotic normality when $m,p$ are fixed is standard for maximum likelihood estimators.
+Here we illustrate the idea on $\mathrm{RP\text{-}MLE}$ for one random splitting.
+By mean value theorem
+$\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\widehat{\bm{\theta}}_{k}^{(n)})+\left[\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\right]\mathrm{d}\tau(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})$
+$=\left[\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}).$
+Here the second row comes from the optimality condition of $\widehat{\bm{\theta}}_{k}^{(n)}$.
+Now under some regularity conditions, we can use the consistency of $\widehat{\bm{\theta}}_{k}^{(n)}$ to show
+$\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\approx\bm{H}^{\infty}.$ (20)
+Then $\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}\approx(\bm{H}^{\infty})^{\dagger}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$.
+Note that $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ is zero-mean and independent between different user $t$.
+This independence also holds for $\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ and $\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$.
+Then we can invoke central limit theorem to reach the desired result.
 Appendix B Degree and spectral properties of the comparison graphs
@@
 B.1 Degree range of $\mathcal{G}_{X}$
-Recall that $n_{t}$ is the number of neighbors of user $t$ in $\mathcal{G}_{X}$.
-Furthermore, we denote $m_{i}$ as the number of users that is compared with problem $i$ and at least another item, i.e.,
-$m_{i}\coloneqq\left|\{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\}\right|.$ (17)
-The following lemma controls the size of $n_{t}$ and $m_{i}$.
+Recall that $m_{t}$ is the number of neighbors of user $t$ in $\mathcal{G}_{X}$.
+Furthermore, we denote $n_{i}$ as the number of users that is compared with problem $i$ and at least another item, i.e.,
+$n_{i}\coloneqq\left|\{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2\}\right|.$ (21)
+The following lemma controls the size of $m_{t}$ and $n_{i}$.
 Lemma 4 (Degree bounds in $\mathcal{G}_{X}$).
 Suppose that $np\geq C\log(n)$ for some large enough constant $C>0$ and that $m\leq n^{\alpha}$ for some constant $\alpha>0$. Then with probability at least $1-2n^{-10}$, for all $i\in[m]$, we have
-$\frac{1}{4}np\leq m_{i}\leq\frac{3}{2}np.$ (18)
+$\frac{1}{4}np\leq n_{i}\leq\frac{3}{2}np.$ (22)
 Moreover, with probability at least $1-n^{-10}$, for all $t\in[n]$, we have
-$n_{t}\leq\left(\frac{3}{2}mp\right)\vee 165\log(n).$ (19)
+$m_{t}\leq\left(\frac{3}{2}mp\right)\vee 165\log(n).$ (23)
 Proof.
@@
 Fix any $t,i$. One has
-$\mathbb{P}\left[t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\right]$ $=\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]-\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\text{ and }n_{t}=1\right]$
+$\mathbb{P}\left[t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2\right]$ $=\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]-\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\text{ and }m_{t}=1\right]$
 $=p-p(1-p)^{m-1}$
@@
 $\geq\frac{1}{2}p,$ (24)
-as long as $mp\geq 2$. Let $\mu_{i}\coloneqq\mathbb{E}[m_{i}]$. By the linearity of expectation, we have
+as long as $mp\geq 2$. Let $\mu_{i}\coloneqq\mathbb{E}[n_{i}]$. By the linearity of expectation, we have
 $np/2\leq\mu_{i}\leq\sum_{t}\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]=np.$ (25)
 Fix $i\in[m]$. Since the sampling is independent with different $t$, by the Chernoff bound,
-$\mathbb{P}[|m_{i}-\mu_{i}|\leq(1/2)\mu_{i}]$ $\leq 2e^{-\frac{1}{12}\mu_{i}}\leq 2e^{-\frac{1}{24}np}\leq m^{-1}n^{-10}$
+$\mathbb{P}[|n_{i}-\mu_{i}|\leq(1/2)\mu_{i}]$ $\leq 2e^{-\frac{1}{12}\mu_{i}}\leq 2e^{-\frac{1}{24}np}\leq m^{-1}n^{-10}$
 as long as $np\geq C\log(n)$ for large enough constant $C$. Applying (25) and union bound on $i\in[m]$ yields (22).
 Moving on to (23), we first consider the case where $mp\geq 110\log(n)$. By Chernoff bound,
-$\mathbb{P}[n_{t}\geq(3/2)mp]\leq e^{-\frac{1}{10}mp}\leq n^{-11}.$
-In the case of $mp<110\log(n)$, the quantity $\mathbb{P}\left[n_{t}\geq 165\log(n)\right]$ clear decreases as $p$ decreases.
+$\mathbb{P}[m_{t}\geq(3/2)mp]\leq e^{-\frac{1}{10}mp}\leq n^{-11}.$
+In the case of $mp<110\log(n)$, the quantity $\mathbb{P}\left[m_{t}\geq 165\log(n)\right]$ clear decreases as $p$ decreases.
 So we may use the case $mp=110\log(n)$ to bound this quantity and conclude that
-$\mathbb{P}\left[n_{t}\geq 165\log(n)\right]\leq n^{-11}.$
+$\mathbb{P}\left[m_{t}\geq 165\log(n)\right]\leq n^{-11}.$
 Finally we apply union bound on $t\in[n]$ to reach (23).
@@
 $=\sum_{t:(t,i)\in\mathcal{E}_{X}}\sum_{j:j\neq i}L_{ij}^{t}$
-=$\overset{(\text{i})}{=}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}L_{ij}^{t}$
-$\overset{(\text{ii})}{\leq}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}R_{ij}^{t}\leq m_{i}\leq\frac{3}{2}np.$
-Here (i) holds since $L_{ij}^{t}$ can only be 0 when $n_{t}\leq 1$, and (ii) holds since $L_{ij}^{t}\leq R_{ij}^{t}$ by definition.
-For the lower bound of $d_{i}$, notice that for any $(t,i)$, $\sum_{j}L_{ij}^{t}$ is either 00 or 1.
+$\overset{(\text{i})}{=}\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\sum_{j:j\neq i}L_{ij}^{t}$
+$\overset{(\text{ii})}{\leq}\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\sum_{j:j\neq i}R_{ij}^{t}\leq n_{i}\leq\frac{3}{2}np.$
+Here (i) holds since $L_{ij}^{t}$ can only be 0 when $m_{t}\leq 1$, and (ii) holds since $L_{ij}^{t}\leq R_{ij}^{t}$ by definition.
+For the lower bound of $d_{i}$, notice that for any $(t,i)$, $\sum_{j}L_{ij}^{t}$ is either 0 or 1.
 Fix $\mathcal{E}_{X}$ and only consider randomness on $L_{ij}^{t}$.
 By Hoeffding’s inequality,
-$\mathbb{P}\left\{d_{i}-\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j:j\neq i}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\leq-\frac{1}{12\kappa_{2}}np\right\}$ $\leq\exp\left(-\frac{(1/72)\kappa_{2}^{-2}n^{2}p^{2}}{m_{i}}\right)$
+$\mathbb{P}\left\{d_{i}-\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\mathbb{E}\left[\sum_{j:j\neq i}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\leq-\frac{1}{12\kappa_{2}}np\right\}$ $\leq\exp\left(-\frac{(1/72)\kappa_{2}^{-2}n^{2}p^{2}}{n_{i}}\right)$
 $\leq\exp\left(-\frac{\kappa_{2}^{-2}np}{108}\right)$
@@
 $\geq\sum_{j}\mathbb{P}\left[R_{ij}^{t}=1\mid(t,i)\in\mathcal{E}_{X}\right]\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}$
-$\geq\frac{2\lfloor n_{t}/2\rfloor}{n_{t}}\cdot\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq\frac{1}{3\kappa_{2}}$
-as long as $n_{t}\geq 2$. The first inequality here uses Fact 1. Then by definition of $m_{i}$ in (17),
-$\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\geq\frac{1}{3\kappa_{2}}m_{i}.$ (23)
+$\geq\frac{2\lfloor m_{t}/2\rfloor}{m_{t}}\cdot\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq\frac{1}{3\kappa_{2}}$
+as long as $m_{t}\geq 2$. The first inequality here uses Fact 1. Then by definition of $n_{i}$ in (21),
+$\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\geq\frac{1}{3\kappa_{2}}n_{i}.$ (27)
 Combining (26), (27) and (22),
-$d_{i}\geq\frac{1}{3\kappa_{2}}m_{i}-\frac{1}{12\kappa_{2}}np\geq\frac{1}{24\kappa_{2}}np.$
+$d_{i}\geq\frac{1}{3\kappa_{2}}n_{i}-\frac{1}{12\kappa_{2}}np\geq\frac{1}{24\kappa_{2}}np.$
 Applying union bound over $i\in[m]$ yields the desired result.
@@
 For the upper bound, it is clear from Lemmas 12 and 2 that $\lambda_{1}(\bm{L}_{L})\leq 2\max_{i}d_{i}\leq 3np$.
-For the lower bound, we use the matrix Chernoff inequality (see Section 5 of [Tro15]).
+For the lower bound, we use the matrix Chernoff inequality (see Section 5 of [43]).
 Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$.
@@
 For any $t\in[n]$, by (23),
-$0\leq\lambda_{m-1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})\leq\lambda_{1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})=\lambda_{1}(\bm{L}_{L}^{t})\leq 2$
+$0\leq\lambda_{m-1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})\leq\lambda_{1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})=\lambda_{1}(\bm{L}_{L}^{t})\leq 2.$
 The last inequality follows from Lemma 12 since $\sum_{j}\bm{L}_{ij}^{t}\leq 1$ for any $t\in[n]$ and $i\in[m]$. By Fact 1,
@@
 $\geq\frac{1}{2\kappa_{2}}\lambda_{m-1}\left(\bm{R}\sum_{t=1}^{n}\sum_{i>j}\mathbb{E}R_{ij}^{t}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{R}^{\top}\right)$
-Moreover $\sum_{i>j}\mathbb{E}R_{ij}^{t}\geq\frac{1}{2}(\mathbb{E}[n_{t}]-1)=(mp-1)/2$, where the $-1$ accounts for possible unpaired $X_{ti}$.
+Moreover $\sum_{i>j}\mathbb{E}R_{ij}^{t}\geq\frac{1}{2}(\mathbb{E}[m_{t}]-1)=(mp-1)/2$, where the 1-1 accounts for possible unpaired $X_{ti}$.
 By symmetry $\mathbb{E}R_{ij}^{t}$ is the same for any $(i,j)$.
@@
 $=\frac{1}{4\kappa_{1}}\lambda_{m-1}(\bm{L}_{L})\geq\frac{np}{16\kappa_{1}\kappa_{2}}.$
-Appendix C Proof of Lemma 4
+Appendix C Proofs for Section 3.2
+In this section, we provide the proofs for the results related to the non-asymptotic expansion in Section 3.2.
+C.1 Proof of Theorem 2
 We study the MLE $\widehat{\bm{\theta}}$ by analyzing the iterates of preconditioned gradient descent starting from the ground truth.
 Let $\bm{\theta}_{0}=\bm{\theta}^{\star}$ be the starting point and $\eta>0$ be the stepsize that is small enough.
-The preconditioned gradient descent iterates are given by
-$\bm{\theta}^{t+1}=\bm{\theta}^{t}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{t}).$ (24)
-Consider the Taylor expansion of $\nabla\mathcal{L}(\bm{\theta}^{t})$, we have
-$\nabla\mathcal{L}(\bm{\theta}^{t})$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-Y_{ji}^{(l)}+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
-$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
-$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}+\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})(\delta_{i}^{t}-\delta_{j}^{t})+\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}\right)(\bm{e}_{i}-\bm{e}_{j})\right).$
-Here $\bm{\delta}^{t}\coloneqq\bm{\theta}^{t}-\bm{\theta}^{\star}$ and for all $(i,j)$, $\xi_{ij}^{t}\in\mathbb{R}$ is some number that lies between $\theta_{i}^{\star}-\theta_{j}^{\star}$ and $\theta_{i}^{t}-\theta_{j}^{t}$.
-As $\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=z_{ij}$ and $\delta_{i}^{t}-\delta_{j}^{t}=(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{\delta}^{t}$, we have
-$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}\bm{\delta}^{t}-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
-where $\bm{r}^{t}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}[\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}(\bm{e}_{i}-\bm{e}_{j})]$.
+At iteration $\tau$, the preconditioned gradient descent update is given by
+$\bm{\theta}^{\tau+1}=\bm{\theta}^{\tau}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\tau}).$ (28)
+Recall the definitions: $\sigma(x)=e^{x}/(1+e^{x})$ is the sigmoid function.
+We also have $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$, $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$ and $\epsilon_{ij}^{t}\coloneqq Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$.
+The total number of observed effective comparisons in $\mathcal{G}_{Y}$ is $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$.
+We have defined $\bm{\bm{B}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ and $\widehat{\bm{\epsilon}}\in\mathbb{R}^{L_{\mathrm{total}}}$ as
+$\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\text{$\quad$(repeat $L_{ij}$ times for edge $(i,j)$)}$
+and
+$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{t}/\sqrt{z_{ij}},\cdots\right]_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}\in\mathbb{R}^{L_{\mathrm{total}}}.$
+Moreover, define a weighted graph Laplacian
+$\bm{L}_{L\tilde{z}}\coloneqq\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+for $\tilde{z}$ being $z$ or $\widehat{z}$, and $\bm{L}_{L\tilde{z}}^{\dagger}$ is its pseudo-inverse.
+Consider the Taylor expansion of $\nabla\mathcal{L}(\bm{\theta}^{\tau})$, we have
+$\nabla\mathcal{L}(\bm{\theta}^{\tau})$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-Y_{ji}^{t}+\sigma(\theta_{i}^{\tau}-\theta_{j}^{\tau})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
+$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-\epsilon_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})+\sigma(\theta_{i}^{\tau}-\theta_{j}^{\tau})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
+$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-\epsilon_{ji}^{t}+\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})(\delta_{i}^{\tau}-\delta_{j}^{\tau})+\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{\tau})(\delta_{i}^{\tau}-\delta_{j}^{\tau})^{2}\right)(\bm{e}_{i}-\bm{e}_{j})\right).$
+Here $\bm{\delta}^{\tau}\coloneqq\bm{\theta}^{\tau}-\bm{\theta}^{\star}$ and for all $(i,j)$, $\xi_{ij}^{\tau}\in\mathbb{R}$ is some number that lies between $\theta_{i}^{\star}-\theta_{j}^{\star}$ and $\theta_{i}^{\tau}-\theta_{j}^{\tau}$.
+As $\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=z_{ij}$ and $\delta_{i}^{\tau}-\delta_{j}^{\tau}=(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{\delta}^{\tau}$, we can rewrite $\nabla\mathcal{L}(\bm{\theta}^{\tau})$ as
+$\nabla\mathcal{L}(\bm{\theta}^{\tau})=\bm{L}_{Lz}\bm{\delta}^{\tau}-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{\tau},$
+where $\bm{r}^{\tau}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\cdot[\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{\tau})(\delta_{i}^{\tau}-\delta_{j}^{\tau})^{2}(\bm{e}_{i}-\bm{e}_{j})]$.
 Feeding this into (28), we have
-$\bm{\delta}^{t+1}=\left(1-\eta\right)\bm{\delta}^{t}-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right)$ (25)
-By definition $\bm{\delta}^{0}=\bm{0}$. Applying this recursive relation $t-1$ times we obtain
-$\bm{\delta}^{t}$ $=-\eta\sum_{i=0}^{t-1}(1-\eta)^{i}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}$
-$=-\left[1-(1-\eta)^{t}\right]\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}.$
-At this point, we invoke an existing result on these terms that have been studied in [YCOM24] for a more general setting.
-The proof of Lemma 2 combined with Lemmas 3 and 4 in [YCOM24] reveal the following properties of (24).
+$\bm{\delta}^{\tau+1}=\left(1-\eta\right)\bm{\delta}^{\tau}-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right)$ (29)
+By definition $\bm{\delta}^{0}=\bm{0}$. Applying this recursive relation $\tau-1$ times we obtain
+$\bm{\delta}^{\tau}$ $=-\eta\sum_{i=0}^{\tau-1}(1-\eta)^{i}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}$
+$=-\left[1-(1-\eta)^{\tau}\right]\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}.$
+At this point, we invoke an existing result on these terms that have been studied in [47] for a more general setting.
+The proof of Lemma 6 combined with Lemmas 7 and 8 in [47] reveal the following properties of (28).
+The proof is deferred to Section C.1.1.
 Lemma 5.
@@
   2. 2.
-    There exist some $\alpha_{1},\alpha_{2}$ obeying $0<\alpha_{1}\leq\alpha_{2}$ such that any $t\in\mathbb{N}$,
-    $\|\bm{\theta}^{t}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}}\leq(1-\eta\alpha_{1})^{t}\|\bm{\theta}^{0}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}},$
+    There exist some $\alpha_{1},\alpha_{2}$ obeying $0<\alpha_{1}\leq\alpha_{2}$ such that any $\tau\in\mathbb{N}$,
+    $\|\bm{\theta}^{\tau}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}}\leq(1-\eta\alpha_{1})^{\tau}\|\bm{\theta}^{0}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}},$
     provided that $0<\eta\leq 1/\alpha_{2}$.
   3. 3.
-    For any $k,l$ and iteration $t\geq 0$,
-    $\left|\left(\theta_{k}^{t}-\theta_{l}^{t}\right)-\left(\theta_{k}^{\star}-\theta_{l}^{\star}\right)\right|\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$ (27)
+    For any $k,l$ and iteration $\tau\geq 0$,
+    $\left|\left(\theta_{k}^{\tau}-\theta_{l}^{\tau}\right)-\left(\theta_{k}^{\star}-\theta_{l}^{\star}\right)\right|\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$ (31)
     for some constant $C_{2}>0$.
   4. 4.
-    For any $k,l$ and iteration $t\geq 0$,
-    $\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{3}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{L}))^{3}}$
+    For any $k,l$ and iteration $\tau\geq 0$,
+    $\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right|\leq C_{3}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{L}))^{3}}$
     for some constant $C_{3}>0$.
@@
 $\frac{np}{4\kappa_{1}\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{Lz})\qquad\text{and}\qquad d_{\max}\leq\frac{3}{2}np.$
-Then the condition (26) holds as long as $np\geq C_{4}\kappa_{1}^{2}\kappa_{2}\log^{3}(n)$ for some large enough constant $C_{4}$. Invoke Lemma 6 to see that for any $(k,l,t)$,
-$\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np},$ (28)
+Then the condition (30) holds as long as $np\geq C_{4}\kappa_{1}^{2}\kappa_{2}\log^{3}(n)$ for some large enough constant $C_{4}$. Invoke Lemma 5 to see that for any $(k,l,\tau)$,
+$\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np},$ (32)
 where $C_{5}>0$ is some constant. Furthermore, the convergence given by Lemma 5 implies that
-$\widehat{\bm{\delta}}$ $=\lim_{t\rightarrow\infty}\bm{\delta}^{t}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\underbrace{\eta\lim_{t\rightarrow\infty}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}}_{\eqqcolon\bm{r}}.$ (29)
-[…]
-$|r_{k}-r_{j}|\leq\eta\lim_{t\rightarrow 0}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}=C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
+$\widehat{\bm{\delta}}$ $=\lim_{\tau\rightarrow\infty}\bm{\delta}^{\tau}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\underbrace{\eta\lim_{\tau\rightarrow\infty}\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}}_{\eqqcolon\bm{r}}.$ (33)
+It remains to control the $\ell_{\infty}$ norm of $\bm{r}$. For any $(k,l$), (32) shows that
+$|r_{k}-r_{j}|\leq\eta\lim_{\tau\rightarrow 0}\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}=C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
 As $\bm{1}^{\top}\bm{L}_{Lz}^{\dagger}=0$, the above inequality implies that
@@
 The proof is now completed.
-Appendix D Proof of Theorem 2
-We start with the proof of (5). By Lemma 4 we can express the MLE estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ as
+C.1.1 Proof of Lemma 5
+Lemma 5 is a direct combination of Lemmas 7, Lemma 8, and the proof of Lemma 6 in [47].
+The results therein describe controls the error for the same dynamic with a more general setting.
+Thus it is directly applicable to our case.
+For clarity, in this section we explain the connection and relate our notations with the ones used in [47].
+For any $(k,l)\in[m]^{2}$, let
+$Q_{kl}\coloneqq C_{1}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{Lz}))^{3}}\qquad\text{and}\qquad B_{kl}\coloneqq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{Lz})}},$
+where $C_{1},C_{2}$ are some constants.
+Note that our setting in this paper is unweighted and has different number of observation for each edge in $\mathcal{G}_{Y}$.
+Moreover $\bm{L}_{wz}$ in [47] correspond to $\bm{L}_{Lz}$ in this paper.
+Then Lemma 3 and Lemma 4 in [47] imply that $Q_{kl}$ and $B_{kl}$ satisfies equation (9) therein.
+Furthermore, as we have assumed in(30), as long as $C_{2}$ is large enough, $Q_{kl}\leq 4B_{kl}$ for any $(k,l)$.
+Thus we can invoke Lemma 2 as well as all its proof.
+Lemma 20 in [47] implies the first two items in Lemma 5 herein.
+Lemma 19 in [47] implies the item 3.
+Finally, item 4 appears in the second-to-last equation block in the proof of Lemma 19 in [47].
+C.2 Proof of Proposition 3
+We condition the whole analysis on the high probability event when Lemmas 2 and 3 hold, which happens with probability at least $1-O(n^{-10})$.
+We can expand $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ as
+$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left[Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})+\bm{r}$
+where $\|\bm{r}\|_{\infty}\leq C_{1}\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C_{1}$.
+Consider $\bm{L}_{Lz}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$. As $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$, $\lambda_{m-1}(\bm{L}_{Lz})>0$, and $\widehat{\bm{\delta}}^{\top}\bm{1}_{m}=0$,
+$(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ $=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
+$=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\underbrace{\left[Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{t}}+\bm{r}.$
+Conditional on $(\mathcal{E}_{Y},\{t:L_{ij}^{t}=1\})$, $\bm{u}_{ij}^{t}$ are independent random variables. It is also easy to see that $\bm{u}_{ij}^{t}$ is zero-mean and has covariance
+$\mathbb{E}\left[\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right]$ $=z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
+When rescaled by $(\bm{L}_{Lz}^{\dagger})^{1/2}$, it is also bounded in the third moment of spectral norm by
+$\mathbb{E}\left[\|(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{u}_{ij}^{t}\|^{3}\right]\leq 2^{3/2}\|\bm{L}_{Lz}^{\dagger}\|^{3/2}\leq\frac{2^{15/2}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{3/2}},$ (34)
+where the last inequality uses Lemma 3. Summing up across $i,j$ and $l$, we have
+$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathbb{E}\left[\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right]$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+$=\bm{L}_{Lz}$
+The last line holds since $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$ and $\lambda_{m-1}(\bm{L}_{Lz})>0$.
+Now using multivariate Berry–Esseen theorem (see, e.g., [36]), let $\widetilde{\bm{x}}$ be a random variable such that
+$\bm{x}\sim\mathcal{N}\left(\bm{0}_{m},\bm{L}_{Lz}^{\dagger}\right),$
+then
+$\sup_{A\in\mathcal{C}_{m}}\left|\mathbb{P}\left[-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\in A\mid\mathcal{G}_{Y}\right]-\mathbb{P}(\bm{x}\in A)\right|$ $\leq C_{2}m^{1/4}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathbb{E}\left[\|(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{u}_{ij}^{t}\|^{3}\right]$
+$\overset{\text{(i)}}{\leq}C_{3}\frac{m^{5/4}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{1/2}}.$
+Here (i) uses 2 and (34), and $C_{2},C_{3}$ are some absolute constants. We then reach the desired conclusion by adding the probability upper bound that Lemmas 2 and 3 fail.
+C.3 Proof of Proposition 4
+We start with the proof of (6). By Theorem 2 we can express the MLE estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ as
 $\widehat{\bm{\theta}}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$ (35)
@@
 We first focus on the main term $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$. Expanding $\bm{B}$ and $\widehat{\bm{\epsilon}}$, we rewrite it as
-$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\epsilon_{ij}^{(l)}\bm{L}_{Lz}^{\dagger}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
-It is easy to see that conditional on $\mathcal{G}_{Y}$, $\{\bm{u}_{ij}^{(l)}\}_{i,j,l}$ is a set of independent zero-mean random variables.
+$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\underbrace{\epsilon_{ij}^{t}\bm{L}_{Lz}^{\dagger}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{t}}.$
+It is easy to see that conditional on $(\mathcal{E}_{Y},\{t:L_{ij}^{t}=1\})$, $\{\bm{u}_{ij}^{t}\}_{i,j,t}$ is a set of independent zero-mean random variables.
 Thus we can expand $\mathbb{E}[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}]$ to be
-$\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]$ $=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\bm{u}_{ij}^{(l)\top}\bm{u}_{ij}^{(l)}$
-$=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathrm{Trace}\left(\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right)$
-[…]
-[…]
+$\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]$ $=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\bm{u}_{ij}^{t\top}\bm{u}_{ij}^{t}$
+$=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathrm{Trace}\left(\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right)$
+$\overset{\text{(i)}}{=}\mathrm{Trace}\left(\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\bm{L}_{Lz}^{\dagger}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{L}_{Lz}^{\dagger}\right)$
+$\overset{\text{(ii)}}{=}\mathrm{Trace}(\bm{L}_{Lz}^{\dagger}).$ (36)
 Here (i) follows from the equality
-$z_{ij}=\mathrm{Var}(\epsilon_{ji}^{(l)})=\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}},$
+$z_{ij}=\mathrm{Var}(\epsilon_{ji}^{t})=\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}},$
 and (ii) follows from the definition of $\bm{L}_{Lz}$. By Lemma 3,
 $\frac{m}{2np}\leq(m-1)\lambda_{m-1}(\bm{L}_{Lz}^{\dagger})\leq\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})\leq m\|\bm{L}_{Lz}^{\dagger}\|\leq\frac{16\kappa_{1}\kappa_{2}m}{np}.$ (37)
-Moreover, $\{\epsilon_{ij}^{(l)}\}_{i,j,l}$ is a set of sub-Gaussian random variable with variance proxy $1/z_{ij}$ (see, e.g., Section 2.5 in [Ver18]) and $1/z_{ij}\leq 4\kappa_{1}$ by Lemma 9.
-Applying Hanson-Wright inequality (see [RV13]), for any scalar $a>0$ we have
+Moreover, $\{\epsilon_{ij}^{t}\}_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}$ is a set of sub-Gaussian random variable with variance proxy $1/z_{ij}$ (see, e.g., Section 2.5 in [44] for the definition of sub-Gaussian random variable) and $1/z_{ij}\leq 4\kappa_{1}$ by Lemma 11.
+Applying Hanson-Wright inequality (see [39]), for any scalar $a>0$ we have
 $\mathbb{P}\left[\left|\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}-\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]\right|>a\right]$ (38)
@@
 $=\sqrt{\mathrm{Trace}\left(\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\bm{B}^{\top}\right)}$
-[…]
-[…]
+$\overset{\text{(i)}}{=}\sqrt{\mathrm{Trace}\left(\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\right)}$
+$\overset{\text{(ii)}}{\leq}\frac{\sqrt{m}}{\lambda_{m-1}(\bm{L}_{Lz})}\overset{\text{(iii)}}{\leq}\frac{16\kappa_{1}\kappa_{2}\sqrt{m}}{np}.$
 Here (i) follows from the fact that $\bm{B}\bm{B}^{\top}=\bm{L}_{Lz}$, (ii) follows from Lemma 3 and the fact that $\mathrm{Trace}(\bm{M})\leq m\|\bm{M}\|$ for any $m\times m$ matrix $\bm{M}$, and (iii) follows from Lemma 3.
@@
 This implies that $\|\bm{L}_{Lz}^{\dagger}\|\leq 16\kappa_{1}\kappa_{2}/(np)$ and $\|\bm{L}_{L\hat{z}}^{\dagger}\|\leq 32\kappa_{1}\kappa_{2}/(np)$.
-Using the perturbation bound of pseudo-inverse (see Theorem 4.1 in [Wed73]), we have
+Using the perturbation bound of pseudo-inverse (see Theorem 4.1 in [46]), we have
 $\|\bm{L}_{Lz}^{\dagger}-\bm{L}_{L\hat{z}}^{\dagger}\|$ $\leq 3\cdot\|\bm{L}_{Lz}^{\dagger}\|\cdot\|\bm{L}_{L\hat{z}}^{\dagger}\|\cdot\|\bm{L}_{Lz}-\bm{L}_{L\hat{z}}\|$
@@
 for some constant $C_{9}>0$.
-Appendix E Proofs for inference
-In this section, we present the analysis for the inference results we presented in Section 3.3.
-E.1 Proof of Theorem 3
-We condition the whole analysis on the high probability event when Lemmas 4 and 3 holds. Let $\widehat{\bm{\delta}}\coloneqq\widehat{\bm{\theta}}-\bm{\theta}^{\star}$. Expanding (15), we have
-$\widehat{\bm{\delta}}$ $=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
-$=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})+\bm{r},$
-where $\|\bm{r}\|_{\infty}\leq C_{1}\kappa_{1}^{6}\log^{2}(n)/(np)$ for some constant $C_{1}$.
-We start with the case with $\widetilde{\bm{z}}=\bm{z}$.
-Consider $\bm{L}_{Lz}\widehat{\bm{\delta}}$.
-As $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$, $\lambda_{m-1}(\bm{L}_{Lz})>0$, and $\widehat{\bm{\delta}}^{\top}\bm{1}_{m}=0$,
-$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}$ $=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}$
-$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}+\bm{L}_{Lz}^{1/2}\bm{r}.$
-Conditional on $(\mathcal{E}_{Y},\{L_{ij}\})$, $\bm{u}_{ij}^{(l)}$ are independent random variables. It is easy to see that $\bm{u}_{ij}^{(l)}$ is zero-mean and has covariance
-$\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}.$
-It is also bounded in third moment of its spectral norm by
-$\mathbb{E}\left[\|\bm{u}_{ij}^{(l)}\|^{3}\right]\leq 2^{3/2}\|\bm{L}_{Lz}^{\dagger}\|^{3/2}\leq\frac{2^{15/2}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{3/2}},$
-where the last inequality uses Lemma 3. Summing up across $i,j$ and $l$, we have
-$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=(\bm{L}_{Lz}^{\dagger})^{1/2}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}$
-$=(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{L}_{Lz}(\bm{L}_{Lz}^{\dagger})^{1/2}$
-$=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
-The last line holds since $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$ and $\lambda_{m-1}(\bm{L}_{Lz})>0$.
-Now using multivariate CLT (see, e.g., [Rai19]), we have that as $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$,
-$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
-As $m$ is fixed, Lemma 2 implies that $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$ is equivalent to $np\rightarrow\infty$.
-For the residual term, Lemma 4 and Lemma 3 imply that with probability at least $1-O(n^{-10})$,
-$\|\bm{L}_{Lz}^{1/2}\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\|\bm{r}\|\leq\|\bm{L}_{Lz}^{1/2}\|\cdot\sqrt{m}\|\bm{r}\|_{\infty}\leq\frac{C_{1}\sqrt{m}\kappa_{1}^{6}\log^{2}(n)}{\sqrt{np}}.$
-Then as $np/\log^{4}(n)\rightarrow\infty$,
-$\bm{L}_{Lz}^{1/2}\bm{r}\rightarrow\bm{0}$
-in probability. Combining the limiting distribution of both two terms, we have
-$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
-In the case of $\widetilde{\bm{z}}=\widehat{\bm{z}}$, we introduce the following lemma. The proof is deferred to the end of this section.
+Appendix D Proofs for Section 3.3
+In this section we prove the results for Section 3.3, including Theorem 3, Theorem 4, Proposition 5, and (9).
+To facilitate our analysis, we decompose the loss function by each user and random split.
+For $\mathrm{WP\text{-}MLE}$, recall that we let the loss function associated with user $t$ to be
+$\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}\right.\\ \left.\hskip 80.00012pt{}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
+We have also defined the loss function associated with random split $k$ and user $t$ for $\mathrm{MRP\text{-}MLE}$ as
+$\mathcal{L}_{k}^{(t)}(\bm{\theta})=-\sum_{\begin{subarray}{c}(i,j):i>j,\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right],$
+where $\Omega_{k}$ denotes the set of paired item-item-user tuples for the $k$-th split.
+D.1 Proof of Theorem 3
+We first fix a random split $k$. By mean value theorem
+$\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\widehat{\bm{\theta}}_{k}^{(n)})+\left[\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})$
+$=\left[\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}).$ (41)
+Here the second line comes from the optimality condition for $\widehat{\bm{\theta}}_{k}^{(n)}$.
+We first consider the Hessian part.
+We start by showing the almost surely convergence of $\widehat{\bm{\theta}}_{k}^{(n)}$ and controlling the $\ell_{\infty}$ norm of $\widehat{\bm{\theta}}_{k}^{(n)}$.
+By Theorem 1, we know that with probability at least $1-O(n^{-10})$, $\|\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}\|\leq C\kappa_{1}\kappa_{2}^{1/2}\sqrt{\frac{m\log(n)}{np}}$ for some constant $C$.
+This and the Borel-Cantelli lemma implies that $\widehat{\bm{\theta}}_{k}^{(n)}\overset{\mathrm{a.s.}}{\rightarrow}\bm{\theta}^{\star}$ and $\|\widehat{\bm{\theta}}_{k}^{(n)}\|_{\infty}\leq 2\kappa$ for large enough $n$.
+Furthermore, observe that the Hessian
+$\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta})=\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\frac{e^{\theta_{i}}e^{\theta_{j}}}{(e^{\theta_{i}}+e^{\theta_{j}})^{2}}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+is Lipschitz-continuous for $\{\bm{\theta}:\|\bm{\theta}\|_{\infty}\leq 2\kappa\}$ with a uniform Lipschitz constant for all $k,t$. Then we conclude that
+$\int_{\tau=0}^{1}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\overset{\mathrm{a.s.}}{\rightarrow}\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}).$ (42)
+Recall that the sampling, random pairing, and comparisons are all independent between different users.
+Moreover, the Hessian and its second moment are both bounded.
+Then we can invoke the strong law of large number to say that
+$\frac{1}{n}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\overset{\mathrm{a.s.}}{\rightarrow}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})=\bm{H}^{\infty},$ (43)
+where $\bm{H}^{\infty}$ is defined in (8a).
+Note that this equality holds trivially for $k=1$ and then by symmetry holds all $k$.
+We also make the following claim.
+The proof is deferred to the end of this section.
 Lemma 7.
-Instate the assumptions of Theorem 1. Then with probability at least $1-C_{1}n^{-10}$ for some constant $C_{1}$, one has
-$\|\bm{L}_{Lz}^{1/2}-\bm{L}_{L\hat{z}}^{1/2}\|\leq C_{2}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}$
-for some constant $C_{2}$.
-Now the above lemma and Theorem 1 imply
-$\left\|\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right\|\leq\|\bm{L}_{L\widetilde{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|\cdot\|\widehat{\bm{\delta}}\|\leq C_{3}\kappa_{1}^{3}\kappa_{2}\sqrt{\frac{m\log^{2}(n)}{np}}$
-for some constant $C_{3}>0$. Therefore, as $np/\log^{4}(n)\rightarrow\infty$,
-$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\rightarrow 0$
-in probability and hence
-$\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}=\left(\bm{L}_{L\widehat{z}}^{1/2}\widehat{\bm{\delta}}-\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\right)+\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
+Instate the assumptions of Theorem 1, we have that $\lambda_{m-1}\left(\bm{H}_{k}^{\infty}\right)>0$ and $\bm{H}_{k}^{\infty\top}\bm{1}_{m}=\bm{0}_{m}$.
+Now we move to the gradient part. Recall that the gradient
+$\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta})$
+$=\mathbb{E}_{\mathrm{s+r+d+c}}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)(\bm{e}_{i}-\bm{e}_{j})\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
+$=\bm{0}_{m}$
+is a zero-mean random variable and independent across different $t$. Recall that for all $k_{1}\in[n_{\mathrm{split}}]$,
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{same}}^{\infty}.$
+Note that this equality holds trivially for $k=1$ and then by symmetry holds all $k$. Similarly for all $k_{1}\neq k_{2}$,
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{2}}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{diff}}^{\infty}.$
+Note that $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ is bounded for all $k,t$.
+We can then invoke the central limit theorem for triangular arrays (see Theorems 3.4.10 and 3.10.6 in [12]) to reach that
+$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}\left[\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right]\overset{\mathrm{d}}{\rightarrow}$
+$\hskip 100.00015pt\mathcal{N}\left(\bm{0},\frac{1}{n_{\mathrm{split}}^{2}}\sum_{k_{1}=1}^{n_{\mathrm{split}}}\sum_{k_{2}=1}^{n_{\mathrm{split}}}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{k_{1}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{k_{2}}^{(t)}(\bm{\theta}^{\star})^{\top}\right).$
+We can then rewrite the above formula as
+$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}\left[\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right]\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right).$ (44)
+We now combine the gradient and the Hessian parts together. By (41), (42) and (43), we know that
+$\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})-\bm{H}^{\infty}(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})\overset{\mathrm{a.s.}}{\rightarrow}\bm{0}.$
+Recall that $\bm{\theta}^{\star\top}\bm{1}_{m}=\widehat{\bm{\theta}}_{k}^{(n)\top}\bm{1}_{m}=0$ by design and $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})^{\top}\bm{1}_{m}=0$ by definition.
+Combining this with Lemma 7, we can take the pseudo-inverse of $\bm{H}^{\infty}$ to reach
+$\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}\overset{\mathrm{a.s.}}{\rightarrow}\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right].$
+Now we average across all random splittings to conclude that
+$\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\overset{\mathrm{a.s.}}{\rightarrow}-\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n}\sum_{t=1}^{n}\frac{1}{n_{\mathrm{split}}}\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})\right].$
+Combining this with (44), we have that
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}^{\infty}\right)^{\dagger}\left[\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same}}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff}}^{\infty}\right]\left(\bm{H}^{\infty}\right)^{\dagger}\right).$
 Proof of Lemma 7.
-Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then
-$\|\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2}\|=\|\bm{R}(\bm{L}_{L\hat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|.$
-We also observe that $(\bm{R}\bm{L}_{L\widetilde{z}}\bm{R}^{\top})^{1/2}=\bm{R}\bm{L}_{L\widetilde{z}}^{1/2}\bm{R}^{\top}$ for $\widetilde{z}\in\{z,\widehat{z}\}$. As shown in the proof of Lemma 7,
-$\|\bm{L}_{L\hat{z}}-\bm{L}_{Lz}\|\leq C_{1}\kappa_{1}^{3/2}\sqrt{np\log(n)}$
-for some constant $C_{1}>0$.
-Lemma 3 implies that
-$\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})=\left[\lambda_{m-1}(\bm{L}_{Lz})\right]{}^{1/2}\geq\frac{\sqrt{np}}{4\kappa_{1}^{1/2}\kappa_{2}^{1/2}}.$
-Then by Lemma 2.1 in [Sch92],
-$\|\bm{R}(\bm{L}_{L\widehat{z}}^{1/2}-\bm{L}_{Lz}^{1/2})\bm{R}^{\top}\|\leq\frac{\|\bm{R}(\bm{L}_{L\widehat{z}}-\bm{L}_{Lz})\bm{R}^{\top}\|}{\lambda_{m-1}(\bm{R}\bm{L}_{Lz}^{1/2}\bm{R}^{\top})}\leq 4C_{1}\kappa_{1}^{2}\kappa_{2}^{1/2}\sqrt{\log(n)}.$
-E.2 Proof of Corollary 1
-Observe that
-$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}=\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{L}_{L\tilde{z}}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})}{\sqrt{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}}.$
-Then Theorem 3 implies
-$\frac{\bm{e}_{i}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}(\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top})(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}\bm{e}_{i}}{[\bm{L}_{L\tilde{z}}^{\dagger}]_{ii}}=1.$
-The last inequality follows from the fact that $\bm{1}_{m}^{\top}(\bm{L}_{L\tilde{z}}^{\dagger})^{1/2}=0$.
-As a side note, we compare this with a similar result for the BTL model.
-Proposition 4.1 in [GSZ23] studies the uncertainty quantification of MLE in the BTL model with uniform observation.
-It says that
-$\frac{\widehat{\theta}_{i}-\theta_{i}^{\star}}{\sqrt{\left[\bm{L}_{Lz}\right]_{ii}^{-1}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$
-This is close to our result as $[\bm{L}_{Lz}]_{ii}^{-1}$ can be viewed as an approximation of $[\bm{L}_{Lz}^{\dagger}]_{ii}$ in the context of the BTL model.
-To see why, consider the simplified case where $\bm{\theta}^{\star}=\bm{0}_{n}$ and $L=1$.
-As $np\rightarrow\infty$, the Laplacian $\bm{L}_{Lz}$ concentrate to its expectation $\mathbb{E}[\bm{L}_{Lz}]=np\bm{I}_{n}-p\bm{1}_{n}\bm{1}_{n}^{\top}$.
-Then $\bm{L}^{\dagger}\approx(1/np)(\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{\top})$ and $[\bm{L}_{Lz}^{\dagger}]_{ii}\approx[\bm{L}_{Lz}]_{ii}^{-1}$
-Appendix F Auxiliary lemmas
+It suffices to show that $\lambda_{m-1}(\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}))\geq\beta n$ for some $\beta>0$ that does not depend on $n$. Recall that
+$\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\left(-\frac{e^{\theta_{i}^{\star}+\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}\right)(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
+$=\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[-z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right].$
+Invoking Lemma 3, we have that with probability at least $1-O(n^{-10})$, $\lambda_{m-1}(\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}))\geq np/(16\kappa_{1}\kappa_{2})$, so we are done.
+D.2 Proof for Theorem 4
+The proof for the asymptotic normality of $\mathrm{WP\text{-}MLE}$ is very similar to the proof for $\mathrm{MRP\text{-}MLE}$.
+We start with a lemma on the asymptotic consistency of $\mathrm{WP\text{-}MLE}$.
+Lemma 8.
+Instate the assumptions of Theorem 1. Then as $n\rightarrow\infty$,
+$\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}\overset{\mathrm{p}}{\rightarrow}\bm{\theta}^{\star}.$
+Recall that in the assumption of Theorem 3, we do not specified the limit of Hessian and gradient of the loss function of the $\mathrm{WP\text{-}MLE}$.
+In the following lemma, we show that they are implied by the limit of Hessian and gradient of the $\mathrm{MRP\text{-}MLE}$ loss function.
+Lemma 9.
+Let $\bm{H}^{\infty}$ and $\bm{V}_{\mathrm{diff}}^{\infty}$ be defined as in Theorem 3. Then we have that
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})=\bm{H}^{\infty}$ (45)
+and
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})^{\top}=\bm{V}_{\mathrm{diff}}^{\infty}.$ (46)
+The proof of these two lemmas are deferred to Section D.2.1 and Section D.2.2.
+The rest of the proof for the asymptotic normality of $\mathrm{WP\text{-}MLE}$ is identical to the proof of $\mathrm{MRP\text{-}MLE}$. For conciseness we omit it here.
+D.2.1 Proof of Lemma 8
+Recall that
+$\frac{1}{n}\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$
+$=-\frac{1}{n}\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}\right$.
+$\hskip 150.00023pt{}\left.+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
+Let $\bm{\Theta}\coloneqq\{\bm{\theta}\in\mathbb{R}^{m}:\bm{1}_{m}^{\top}\bm{\theta}=0,\|\bm{\theta}-\bm{\theta}^{\star}\|_{\infty}\leq 10\}$.
+Let $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}\coloneqq\arg\min_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$.
+We make the following claims.
+  1. 1.
+    As $n\rightarrow\infty$, the convergence of
+    $\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\rightarrow\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})$
+    is uniform on $\bm{\Theta}$.
+  2. 2.
+    Property 1 implies that $\frac{1}{n}\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$ converges uniformly in probability to $\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})$ (see [32] for the definition of this type of convergence).
+  3. 3.
+    $\bm{\theta}^{\star}$ is the unique minimizer of $\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})$.
+It is obvious that $\bm{\theta}^{\star}\in\bm{\Theta}$.
+Then with these properties, we can invoke Theorem 2.1 in [32] to conclude that $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ is consistent, i.e., $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}\overset{\mathrm{p}}{\rightarrow}\bm{\theta}^{\star}$ as $n\rightarrow\infty$.
+This further shows that $\|\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\|\leq 5$ with probability at least $1-n^{-10}$ for large enough $n$, and therefore it is a global minimizer in $\bm{\Theta}$ and a local minimizer in $\{\bm{\theta}\in\mathbb{R}^{m}:\bm{1}_{m}^{\top}\bm{\theta}=0\}$.
+Similar to the case of $\mathrm{MRP\text{-}MLE}$ (see Lemma 7, we omit the full proof here for conciseness), $\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$ is strictly convex (modulo $\bm{1}_{m}^{\top}\bm{\theta}=0$) for large enough $n$.
+Then $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ being a local minimum means that it is also the global minimum, i.e., and $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}=\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}$ a.s.
+with probability at least $1-n^{-10}$ for large enough $n$.
+Then, we have that $\widetilde{\bm{\theta}}_{\mathrm{WP}}^{(n)}\overset{\mathrm{p}}{\rightarrow}\bm{\theta}^{\star}$ as $n\rightarrow\infty$ implies $\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}\overset{\mathrm{p}}{\rightarrow}\bm{\theta}^{\star}$ as $n\rightarrow\infty$.
+We now prove these properties in order.
+Proof of Claim 1.
+Observe that for any $\theta\in\bm{\Theta}$,
+$\left|\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right]\right|$
+$\hskip 200.0003pt\leq-\log\left(\frac{1}{e^{\kappa_{1}+20}+1}\right)\leq\kappa_{1}+21.$
+Then
+$\left|\frac{1}{n}\mathcal{L}_{\mathrm{WP}}(\bm{\theta})\right|$ $\leq\frac{1}{n}\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\cdot$
+$\qquad\qquad\left|\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right|$
+$\leq\frac{1}{n}\sum_{t=1}^{n}\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}(\kappa_{1}+21)}{m_{t}(m_{t}-1)}$
+$\leq\frac{1}{n}\sum_{t=1}^{n}\binom{m_{t}}{2}\frac{\kappa_{1}+21}{m_{t}-1}\leq\frac{m_{t}(\kappa_{1}+21)}{2}\leq\frac{m(\kappa_{1}+21)}{2}.$ (47)
+The last line holds since there can at most be $m_{t}$ pairs of responses for each user $t$.
+Therefore $\frac{1}{n}\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$ is uniformly bounded.
+Similarly, we can show that it is also uniformly Lipschitz.
+Then by Arzelà–Ascoli theorem, the pointwise convergence implies that the convergence is uniform on the compact set $\bm{\Theta}$.
+Proof of Claim 2.
+By Hoeffding’s inequality and (47), for any $\bm{\theta}$,
+$\mathbb{P}\left[\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\right|>\epsilon\right]\leq 2\exp\left(-\frac{\epsilon^{2}}{nm(\kappa_{1}+1)}\right).$
+Let $\bm{\Theta}_{\delta}\subset\bm{\Theta}$ such that $\bm{\Theta}_{\delta}$ is a minimal $\delta$-covering of $\bm{\Theta}$ with respect to $\|\cdot\|_{\infty}$.
+It is well known that $|\bm{\Theta}_{\delta}|\leq\mathcal{N}(\bm{\Theta},\|\cdot\|_{\infty},\delta/2)$ where $\mathcal{N}(\bm{\Theta},\|\cdot\|_{\infty},\delta/2)$ is the packing number.
+By volume argument,
+$|\bm{\Theta}_{\delta}|\leq\mathcal{N}(\bm{\Theta},\|\cdot\|_{\infty},\delta/2)\leq\left(\frac{10}{\delta/2}\right)^{m-1}.$
+We claim that
+$\sup_{\bm{\theta}\in\bm{\Theta}}\inf_{\bm{\theta}^{\prime}\in\bm{\Theta}_{\delta}}\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\prime})\right|\leq\frac{m(\kappa_{1}+20)}{2}\cdot\delta.$ (48)
+The proof is deferred to the end of this section. Take $\delta=2\epsilon/(m(\kappa_{1}+20))$. We have that
+$\sup_{\bm{\theta}\in\bm{\Theta}}\inf_{\bm{\theta}^{\prime}\in\bm{\Theta}_{\delta}}\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\prime})\right|\leq\frac{m(\kappa_{1}+20)}{2}\cdot\delta=\epsilon.$ (49)
+Now for any $\epsilon,$ by union bound
+$\mathbb{P}\left[\sup_{\bm{\theta}\in\bm{\Theta}_{\delta}}\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\right|>\epsilon\right]\leq 2\left(\frac{10}{\delta/2}\right)^{m-1}\exp\left(-\frac{2\epsilon^{2}}{nm^{2}(\kappa_{1}+21)}\right).$ (50)
+Moreover, by assumption,
+$\sup_{\bm{\theta}\in\bm{\Theta}_{\delta}}\left|\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})\right|\leq\epsilon.$ (51)
+for large enough $n$. Combining (49), (50), (51), we may conclude that
+$\mathbb{P}\left[\sup_{\bm{\theta}\in\bm{\Theta}}\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})\right|>3\epsilon\right]\leq 2\left(\frac{10}{\delta/2}\right)^{m-1}\exp\left(-\frac{2\epsilon^{2}}{nm^{2}(\kappa_{1}+21)}\right)$
+for large enough $n$, which implies that $\frac{1}{n}\mathcal{L}_{\mathrm{WP}}(\bm{\theta})$ converges uniformly in probability to $\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})$ on $\bm{\Theta}$.
+Proof of Claim 3.
+We first show that $\bm{\theta}^{\star}$ is the minimizer for $\mathbb{E}_{\mathrm{c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$.
+Compute $\mathbb{E}_{\mathrm{c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$, we have that
+$\mathbb{E}_{\mathrm{c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})=-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\frac{e^{\theta_{i}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right].$
+Observe that $\log(x)p+\log(1-x)(1-p)$ as a function of $x$ is minimized at $x=p$, so $\bm{\theta}^{\star}$ is the minimizer for all terms inside $\cdot$] and $\mathbb{E}_{\mathrm{c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$ itself.
+As $\bm{\theta}^{\star}$ is the minimizer for $\mathbb{E}_{\mathrm{c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$, $\bm{\theta}^{\star}$ is also a minimizer of $\mathbb{E}_{\mathrm{s+r+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$ and
+$\overline{\mathcal{L}}_{\mathrm{WP}}(\bm{\theta})=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
+The uniqueness comes from the strict convexity of $\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})$.
+This is similar to the case in $\mathrm{MRP\text{-}MLE}$ (see Lemma 7, we omit the full proof here for conciseness).
+Proof of (48).
+Since $\bm{\Theta}_{\delta}$ is a $\delta$-covering, for any $\bm{\theta}\in\bm{\Theta}$, there exists $\bm{\theta}^{\prime}\in\bm{\Theta}_{\delta}$ such that $\|\bm{\theta}-\bm{\theta}^{\prime}\|_{\infty}\leq\delta$.
+Observe that $x\mapsto 1/(1+e^{-x})$ as a function is $(\kappa_{1}+20)$-Lipschitz if $|x|\leq\kappa_{1}+20$.
+Then similar to (47)
+$\left|\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})-\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\prime})\right|\leq\frac{m(\kappa_{1}+20)}{2}\cdot\delta.$
+D.2.2 Proof of Lemma 9
+Recall that the loss function for $\mathrm{WP\text{-}MLE}$ associated with user $t$ is
+$\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})=-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}\right.\\ \left.\hskip 80.00012pt{}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right],$
+and the loss function for $\mathrm{MRP\text{-}MLE}$ associated with user $t$ is
+$\mathcal{L}_{k}^{(t)}(\bm{\theta})=-\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
+As the random splitting only affect $\Omega_{k}$, we can see that
+$\mathbb{E}_{\mathrm{r}}\mathcal{L}_{k}^{(t)}(\bm{\theta})$
+$=-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\mathbb{E}_{\mathrm{r}}\mathds{1}_{\{(i,j,t)\in\Omega_{k}\}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right]$
+$=-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right]$
+$=\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
+Here
+$\mathbb{E}_{\mathrm{r}}\mathds{1}_{\{(i,j,t)\in\Omega_{k}\}}=\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}$
+comes from the fact that we select $\widetilde{m}_{t}/2$ pairs out of $m_{t}(m_{t}-1)/2$ total pairs and each pair is equally likely to be selected due to symmetry.
+Similarly, we can deduct the same thing for the gradient and the Hessian, i.e., for any $k\in[n_{\mathrm{split}}]$,
+$\mathbb{E}_{\mathrm{r}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta})$ $=\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta});$ (52a)
+$\mathbb{E}_{\mathrm{r}}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta})$ $=\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$ (52b)
+For the gradient, note that each random splitting is independent, so we have that
+$\mathbb{E}_{\mathrm{r}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}$ $=\left[\mathbb{E}_{\mathrm{r}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\right]\left[\mathbb{E}_{\mathrm{r}}\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})\right]^{\top}$
+$=\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})^{\top}$
+Then (52a) implies that
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})^{\top}=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{\mathrm{1}}^{(t)}(\bm{\theta})\nabla\mathcal{L}_{\mathrm{2}}^{(t)}(\bm{\theta})^{\top}=\bm{V}_{\mathrm{diff}}^{\infty}.$
+For the Hessian, (52b) implies that
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})=\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})=\bm{H}^{\infty}.$
+D.3 Proof of Proposition 5
+Consider the expectation of the Hessian and the covariance, we claim that the gradient and Hessian of this special case has expectation
+$\bm{H}_{\pi}^{\infty}\coloneqq\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla^{2}\mathcal{L}_{1}^{(1)}(\bm{\theta}^{\star})=\frac{\beta mp}{4(m-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right];$
+$\bm{V}_{\mathrm{same},\pi}^{\infty}\coloneqq\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(1)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(1)}(\bm{\theta}^{\star})^{\top}=\frac{\beta mp}{2(m-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right];$
+$\bm{V}_{\mathrm{diff},\pi}^{\infty}\coloneqq\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(1)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(1)}(\bm{\theta}^{\star})^{\top}=\frac{\beta m^{2}p^{2}}{4(m-1)(mp-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right].$
+The rest of the proof of this corollary consists of two parts.
+In the first part, we verify that with probability 1, the conditions in Theorem 3 holds.
+In the second part, we compute the asymptotic covariance explicitly.
+We then invoke Theorem 3 to conclude that with probability 1,
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{MRP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}_{\pi}^{\infty}\right)^{\dagger}\left[\frac{1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{same},\pi}^{\infty}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\bm{V}_{\mathrm{diff},\pi}^{\infty}\right]\left(\bm{H}_{\pi}^{\infty}\right)^{\dagger}\right)$
+$=\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\left(\frac{1}{n_{\mathrm{split}}}+\frac{n_{\mathrm{split}}-1}{n_{\mathrm{split}}}\cdot\frac{mp}{2(mp-1)}\right)\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right)$
+and
+$\sqrt{n}\left(\widehat{\bm{\theta}}_{\mathrm{WP}}^{(n)}-\bm{\theta}^{\star}\right)$ $\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(\bm{0},\left(\bm{H}_{\pi}^{\infty}\right)^{\dagger}\bm{V}_{\mathrm{diff},\pi}^{\infty}\left(\bm{H}_{\pi}^{\infty}\right)^{\dagger}\right)$
+$=\mathcal{N}\left(\bm{0},\frac{8(m-1)}{\beta mp}\left(\frac{mp}{2(mp-1)}\right)\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right]\right)$
+D.3.1 Verifying the condition of Theorem 3 and 4
+Consider the terms in (8a)
+$\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})$
+as a function of the random variable $\delta_{t}^{\star}$.
+It has mean $\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})$.
+Since it is bounded, by strong law of large number, as $n\rightarrow\infty$
+$\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\overset{\mathrm{a.s.}}{\rightarrow}\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})=\bm{H}_{\pi}^{\infty}.$
+Similarly we have that
+$\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}\overset{\mathrm{a.s.}}{\rightarrow}\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{same},\pi}^{\infty}$
+and
+$\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}\overset{\mathrm{a.s.}}{\rightarrow}\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}=\bm{V}_{\mathrm{diff},\pi}^{\infty}.$
+For $\mathrm{WP\text{-}MLE}$, we further verify that for any $\bm{\theta}$, invoking the strong law of large number, as $n\rightarrow\infty$,
+$\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+d+c}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\overset{\mathrm{a.s.}}{\rightarrow}\mathbb{E}_{\mathrm{s+d+c+u}}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}).$
+D.3.2 Computing the Hessian
+In this section we compute $\bm{H}_{\pi}^{\infty}$. Recall that
+$\bm{H}_{\pi}^{\infty}=\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla^{2}\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})=\mathbb{E}_{\mathrm{s+d+c+u}}\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$
+and
+$\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{\begin{subarray}{c}(i,j):i>j,\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\left(\frac{e^{\theta_{i}^{\star}+\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}\right)(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
+$=\sum_{\begin{subarray}{c}(i,j):i>j\end{subarray}}\frac{\mathds{1}\{t,i),(t,j)\in\mathcal{G}_{X}\}}{mp-1}\left[\frac{1}{4}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
+Taking expectation, we have
+$\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{\begin{subarray}{c}(i,j):i>j\end{subarray}}\frac{\mathbb{E}_{\mathrm{s}}\mathds{1}\{(t,i),(t,j)\in\mathcal{G}_{X}\}}{4(mp-1)}\cdot$
+$\qquad\mathbb{E}_{\mathrm{d+c+u}}\left[\mathds{1}\{X_{ti}\neq X_{tj}\}\mid(t,i),(t,j)\in\mathcal{G}_{X}\right](\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
+For sampling we have
+$\mathbb{E}_{\mathrm{s}}\mathds{1}\{(t,i),(t,j)\in\mathcal{G}_{X}\}=\frac{\binom{mp}{2}}{\binom{m}{2}}=\frac{mp(mp-1)}{m(m-1)}.$
+For comparison we have
+$\mathbb{E}_{\mathrm{d+c}}\left[\mathds{1}\{X_{ti}\neq X_{tj}\}\mid(t,i),(t,j)\in\mathcal{G}_{X}\right]$
+$\quad=\mathbb{P}_{\mathrm{d+c}}\left[X_{ti}>X_{tj}\mid(t,i),(t,j)\in\mathcal{G}_{X}\right]+\mathbb{P}_{\mathrm{d+c}}\left[X_{ti}<X_{tj}\mid(t,i),(t,j)\in\mathcal{G}_{X}\right]$
+$\quad=\frac{e^{\theta_{i}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}})}\frac{e^{\zeta_{t}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}})}+\frac{e^{\zeta_{t}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}})}\frac{e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}})}$
+$\quad=\frac{e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+1)^{2}},$
+where the last line uses the assumption $\bm{\theta}^{\star}=\bm{0}_{m}$. Combining these with the definition of $\beta$ in (13), we have
+$\mathbb{E}_{\mathrm{s+c+u}}\nabla^{2}\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{\begin{subarray}{c}(i,j):i>j\end{subarray}}\frac{p}{4(m-1)}\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+1)^{2}}\right](\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+$=\sum_{\begin{subarray}{c}(i,j):i>j\end{subarray}}\frac{\beta p}{4(m-1)}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+$=\frac{\beta mp}{4(m-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right].$
+D.3.3 Intra-split covariance
+In this section we compute $\bm{V}_{\mathrm{same},\pi}^{\infty}$. Recall that
+$\bm{V}_{\mathrm{same},\pi}^{\infty}=\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}$
+We expand the term $\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})$ with
+$\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})=\sum_{\begin{subarray}{c}(i,j):i>j\\ (i,j,t)\in\Omega_{k}\end{subarray}}\underbrace{\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)(\bm{e}_{i}-\bm{e}_{j})\mathds{1}\{X_{ti}\neq X_{tj}\}\right]}_{\eqqcolon\bm{u}_{ij}^{(t)}}.$
+We then have that
+$\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}$ $=\mathbb{E}_{\mathrm{s+r+d+c+u}}\sum_{(i,j):i>j}\mathds{1}\{(i,j,1)\in\Omega_{1}\}\bm{u}_{ij}^{(t)}\bm{u}_{ij}^{(t)\top}$
+$=\sum_{(i,j):i>j}a_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$ (53)
+where
+$a_{ij}$ $=\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j,t)\in\Omega_{1}\}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}\mathds{1}\{X_{ti}\neq X_{1j}\}\right]$
+$=\mathbb{P}_{\mathrm{s+r}}\left[(i,j,t)\in\Omega_{1}\right]\mathbb{E}_{\mathrm{d+c+u}}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}\mathds{1}\{X_{ti}\neq X_{1j}\}\mid(i,j,1)\in\Omega_{1}\right].$
+The last equality holds since comparison and user parameter draw are independent since sampling and random pairing. Observe the fact that
+$\sum_{(i,j):i>j}\mathds{1}\{(i,j,1)\in\Omega_{1}\}=\text{number of pairs in a splitting}=m_{t}/2=mp/2.$
+By symmetry,
+$\mathbb{P}_{\mathrm{s+r}}\{(i,j,t)\in\Omega_{1}\}$ $=\binom{m}{2}^{-1}\mathbb{E}_{\mathrm{s+r}}\sum_{(i,j):i>j}\mathds{1}\{(i,j,1)\in\Omega_{1}\}$
+$=\frac{p}{m-1}.$
+Condition on the event $(i,j,t)\in\Omega_{1}$ (which we omit in the formulas below for formatting),
+$\mathbb{E}_{\mathrm{d+c+u}}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}\mathds{1}\{X_{ti}\neq X_{tj}\}\right]$
+$\quad=\mathbb{E}_{\mathrm{u}}\left[\mathbb{P}(X_{ti}<X_{tj})\left(-1+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}+\mathbb{P}(X_{ti}>X_{tj})\left(\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}\right]$
+$\quad=\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{1}^{\star}}}{e^{\zeta_{1}^{\star}}+e^{\theta_{i}^{\star}}}\frac{e^{\theta_{j}^{\star}}}{e^{\zeta_{1}^{\star}}+e^{\theta_{j}^{\star}}}\frac{e^{2\theta_{i}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}+\frac{e^{\theta_{i}^{\star}}}{e^{\zeta_{1}^{\star}}+e^{\theta_{i}^{\star}}}\frac{e^{\zeta_{1}^{\star}}}{e^{\zeta_{1}^{\star}}+e^{\theta_{j}^{\star}}}\frac{e^{2\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}}\right]$
+$\quad=\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{1}^{\star}+\theta_{i}^{\star}+\theta_{j}^{\star}}}{(e^{\zeta_{1}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{1}^{\star}}+e^{\theta_{j}^{\star}})(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}\right].$
+Substitute in the assumption $\bm{\theta}^{\star}=\bm{0}_{m}$, we have
+$\mathbb{E}_{\mathrm{d+c+u}}\left[\left(-\mathds{1}\{X_{ti}<X_{tj}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right)^{2}\mathds{1}\{X_{ti}\neq X_{tj}\}\mid(i,j,1)\in\Omega_{1}\right]$
+$\quad=\mathbb{E}_{\mathrm{u}}\left[\frac{e^{\zeta_{1}^{\star}}}{(e^{\zeta_{1}^{\star}}+1)(e^{\zeta_{1}^{\star}}+1)(1+1)}\right]=\frac{\beta}{2}.$ (54)
+Then
+$a_{ij}=\frac{\beta p}{2(m-1)}.$
+In conclusion
+$\bm{V}_{\mathrm{same},\pi}^{\infty}$ $=\sum_{(i,j):i>j}\frac{\beta p}{2(m-1)}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+$=\frac{\beta mp}{2(m-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right].$
+D.3.4 Inter-split covariance
+In this section we compute $\bm{V}_{\mathrm{diff},\pi}^{\infty}$. Recall
+$\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{(i,j):i>j}\mathds{1}\{(i,j,t)\in\Omega_{1}\}\bm{u}_{ij};$
+$\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{(i,j):i>j}\mathds{1}\{(i,j,t)\in\Omega_{2}\}\bm{u}_{ij}.$
+Then
+$\bm{V}_{\mathrm{diff},\pi}^{\infty}$ $=\mathbb{E}_{\mathrm{s+r+d+c+u}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}$
+$=\sum_{(i_{1},j_{1}):i_{1}>j_{1}}\sum_{(i_{2},j_{2}):i_{2}>j_{2}}\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i_{1},j_{1},t)\in\Omega_{1}\}\mathds{1}\{(i_{2},j_{2},t)\in\Omega_{2}\}\bm{u}_{i_{1}j_{1}}\bm{u}_{i_{2}j_{2}}^{\top}.$ (55)
+Since $\bm{u}_{ij}$ is zero-mean for all $(i,j)$, it suffices to only consider the terms where $i_{1},i_{2},j_{1},j_{2}$ has some overlap.
+We first consider the terms where $(i_{1},j_{1})=(i_{2},j_{2})=(i,j)$.
+$\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j,t)\in\Omega_{1}\}\mathds{1}\{(i,j,t)\in\Omega_{2}\}\bm{u}_{ij}\bm{u}_{ij}^{\top}$
+$\quad=\mathbb{P}_{\mathrm{s+r}}\left[(i,j,t)\in\Omega_{1}\cap\Omega_{2}\right]\mathbb{E}_{\mathrm{d+c+u}}\left[\bm{u}_{ij}\bm{u}_{ij}^{\top}\mid\mathds{1}\{(i,j,t)\in\Omega_{1}\cap\Omega_{2}\}\right].$ (56)
+Similar to (53) and (54), we have $\mathbb{E}_{\mathrm{d+c+u}}\left[\bm{u}_{ij}\bm{u}_{ij}^{\top}\mid\mathds{1}\{(i,j,t)\in\Omega_{1}\cap\Omega_{2}\}\right]=0.5\beta(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$.
+On the other hand,
+$\mathbb{P}_{\mathrm{s+r}}\left[(i,j,t)\in\Omega_{1}\cap\Omega_{2}\right]$ $=\mathbb{P}_{\mathrm{s}}\left[A_{ti}=A_{tj}=1\right]\mathbb{P}_{\mathrm{r}}\left[(i,j,t)\in\Omega_{1}\cap\Omega_{2}\mid A_{ti}=A_{tj}=1\right]$
+$=\frac{\binom{mp}{2}}{\binom{m}{2}}\left(\frac{1}{mp-1}\right)^{2}$
+$=\frac{p}{(m-1)(mp-1)}.$
+Then
+$\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j,t)\in\Omega_{1}\}\mathds{1}\{(i,j,t)\in\Omega_{2}\}\bm{u}_{ij}\bm{u}_{ij}^{\top}=\frac{\beta p}{2(m-1)(mp-1)}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$ (57)
+Now we consider the terms where only two of the four indices overlaps.
+Without loss of generality, we take $i_{1}=i_{2}=i$ and remove the $i_{1}>j_{1},i_{2}>j_{2}$ restriction.
+In other words, we consider
+$\sum_{\begin{subarray}{c}(i,j_{1},j_{2}):\\ i\neq j_{1},i\neq j_{2},j_{1}\neq j_{2}\end{subarray}}\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j_{1},t)\in\Omega_{1}\}\mathds{1}\{(i,j_{2},t)\in\Omega_{2}\}\bm{u}_{ij_{1}}\bm{u}_{ij_{2}}^{\top}$
+$\quad=\sum_{\begin{subarray}{c}(i,j_{1},j_{2}):\\ i\neq j_{1},i\neq j_{2},j_{1}\neq j_{2}\end{subarray}}\mathbb{P}_{\mathrm{s+r}}\left[(i,j_{1},t)\in\Omega_{1},(i,j_{2},t)\in\Omega_{2}\right]\mathbb{E}_{\mathrm{u+c}}\left[\bm{u}_{ij_{1}}\bm{u}_{ij_{2}}^{\top}\mid(i,j_{1},t)\in\Omega_{1},(i,j_{2},t)\in\Omega_{2}\right].$
+We first deal with $\mathbb{P}_{\mathrm{s+r}}\{(i,j_{1},t)\in\Omega_{1},(i,j_{2},t)\in\Omega_{2}\}$,
+$\mathbb{P}_{\mathrm{s+r}}\{(i,j_{1},t)\in\Omega_{1},(i,j_{2},t)\in\Omega_{2}\}$ $=\mathbb{P}_{\mathrm{s}}\left[A_{ti}=A_{tj_{1}}=A_{tj_{2}}=1\right]$
+$\quad\cdot\mathbb{P}_{\mathrm{r}}\left[(i,j_{1},t)\in\Omega_{1},(i,j_{2},t)\in\Omega_{2}\mid A_{ti}=A_{ti}=A_{ti}=1\right]$
+$=\frac{\binom{mp}{3}}{\binom{m}{3}}\left(\frac{1}{mp-1}\right)^{2}$
+$=\frac{p(mp-2)}{(m-1)(m-2)(mp-1)}.$
+Now we consider the $\bm{u}_{ij_{1}}\bm{u}_{ij_{2}}^{\top}$ part. Given $(i,j_{1},t)\in\Omega_{1}$ and $(i,j_{2},t)\in\Omega_{2}$ (for notation simplicity we omit this from now on), we compute
+$\mathbb{E}_{\mathrm{d+c}}\left[\delta_{ij_{1}}\delta_{ij_{2}}\mathds{1}\{X_{ti}\neq X_{tj_{1}},X_{ti}\neq X_{1j_{2}}\}(\bm{e}_{i}-\bm{e}_{j_{1}})(\bm{e}_{i}-\bm{e}_{j_{2}})^{\top}\right].$ (58)
+where
+$\delta_{ij}\coloneqq\left(-\mathds{1}\{X_{1i}<X_{1j}\}+\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}\right).$
+For the scope of this proof we let $E_{1}$, $E_{2}$ be two events, where $E_{1}\coloneqq\{X_{ti}=1,X_{tj_{1}}=0,X_{1j_{2}}=0\}$ and $E_{2}\coloneqq\{X_{ti}=0,X_{tj_{1}}=1,X_{tj_{2}}=1\}$.
+Then we can express (58) as
+$\mathbb{E}_{\mathrm{d+c}}\left[\delta_{ij_{1}}\delta_{ij_{2}}\mathds{1}\{X_{ti}\neq X_{tj_{1}},X_{ti}\neq X_{1j_{2}}\}\right]$
+$\quad=\mathbb{P}(E_{1})\frac{e^{\theta_{j_{1}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}}}\frac{e^{\theta_{j_{2}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}}}+\mathbb{P}(E_{2})\left[-1+\frac{e^{\theta_{j_{1}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}}}\right]\left[-1+\frac{e^{\theta_{j_{2}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}}}\right]$
+$\quad=\frac{e^{\theta_{i}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\zeta_{t}^{\star}}}{e^{\theta_{j_{1}}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\zeta_{t}^{\star}}}{e^{\theta_{j_{2}}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\theta_{j_{1}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}}}\frac{e^{\theta_{j_{2}}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}}}$
+$\quad\quad+\frac{e^{\zeta_{t}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\theta_{j_{1}}^{\star}}}{e^{\theta_{j_{1}}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\theta_{j_{2}}^{\star}}}{e^{\theta_{j_{2}}^{\star}}+e^{\zeta_{t}^{\star}}}\frac{e^{\theta_{i}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}}}\frac{e^{\theta_{i}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}}}$
+$\quad=\frac{e^{\zeta_{t}^{\star}+\theta_{i}^{\star}+\theta_{j_{1}}^{\star}+\theta_{j_{2}}^{\star}}}{(e^{\theta_{j_{1}}^{\star}}+e^{\zeta_{t}^{\star}})(e^{\theta_{j_{2}}^{\star}}+e^{\zeta_{t}^{\star}})(e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}})(e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}})}.$
+Furthermore we taken expectation $\mathbb{E}_{\mathrm{u}}$ to reach
+$\mathbb{E}_{\mathrm{d+c+u}}\left[\delta_{ij_{1}}\delta_{ij_{2}}\mathds{1}\{X_{1i}\neq X_{1j_{1}},X_{1i}\neq X_{1j_{2}}\}\right]=\beta\cdot\frac{e^{\theta_{i}^{\star}+\theta_{j_{1}}^{\star}+\theta_{j_{2}}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j_{1}}^{\star}})(e^{\theta_{i}^{\star}}+e^{\theta_{j_{2}}^{\star}})}=\frac{\beta}{4}.$
+Then
+$\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j_{1},t)\in\Omega_{1}\}\mathds{1}\{(i,j_{2},t)\in\Omega_{2}\}\bm{u}_{ij_{1}}\bm{u}_{ij_{2}}^{\top}$
+$\quad=\frac{\beta p(mp-2)}{4(m-1)(m-2)(mp-1)}(\bm{e}_{i}-\bm{e}_{j_{1}})(\bm{e}_{i}-\bm{e}_{j_{2}})^{\top}.$ (59)
+We can now compute $\bm{V}_{\mathrm{diff},\pi}^{\infty}$ with (55), (57) and (59),
+$\bm{V}_{\mathrm{diff},\pi}^{\infty}$ $=\sum_{(i_{1},j_{1}):i_{1}>j_{1}}\sum_{(i_{2},j_{2}):i_{2}>j_{2}}\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i_{1},j_{1},t)\in\Omega_{1}\}\mathds{1}\{(i_{2},j_{2},t)\in\Omega_{2}\}\bm{u}_{i_{1}j_{1}}\bm{u}_{i_{2}j_{2}}^{\top}.$
+$=\sum_{(i_{1},j_{1}):i_{1}>j_{1}}\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j,t)\in\Omega_{1}\}\mathds{1}\{(i,j,t)\in\Omega_{2}\}\bm{u}_{ij}\bm{u}_{ij}^{\top}$
+$\quad+\sum_{(i,j_{1},j_{2}):i\neq j_{1},i\neq j_{2},j_{1}\neq j_{2}}\mathbb{E}_{\mathrm{s+r+d+c+u}}\mathds{1}\{(i,j_{1},t)\in\Omega_{1}\}\mathds{1}\{(i,j_{2},t)\in\Omega_{2}\}\bm{u}_{ij_{1}}\bm{u}_{ij_{2}}^{\top}$
+$=\sum_{(i_{1},j_{1}):i_{1}>j_{1}}\frac{\beta p}{2(m-1)(mp-1)}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
+$\quad+\sum_{(i,j_{1},j_{2}):i\neq j_{1},i\neq j_{2},j_{1}\neq j_{2}}\frac{\beta p(mp-2)}{4(m-1)(m-2)(mp-1)}(\bm{e}_{i}-\bm{e}_{j_{1}})(\bm{e}_{i}-\bm{e}_{j_{2}})^{\top}.$
+Simplifying this expression gives us
+$\bm{V}_{\mathrm{diff},\pi}^{\infty}$ $=\frac{\beta mp}{2(m-1)(mp-1)}\bm{I}_{m}-\frac{\beta p}{2(m-1)(mp-1)}\bm{1}_{m}\bm{1}_{m}^{\top}$
+$\quad+\frac{\beta mp(mp-2)}{4(m-1)(mp-1)}\bm{I}_{m}-\frac{\beta p(mp-2)}{4(m-1)(mp-1)}\bm{1}_{m}\bm{1}_{m}^{\top}$
+$=\frac{\beta m^{2}p^{2}}{4(m-1)(mp-1)}\left[\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right].$
+D.4 Proof of (9)
+We first introduce a more general result in the following lemma. The proof is deferred to the end of this section.
+Lemma 10.
+Let $\bm{A}\in\mathbb{R}^{n\times d}$, $\bm{B}\in\mathbb{R}^{n\times d}$ be real-value random matrices.
+Suppose that $\mathbb{E}\bm{B}\bm{B}^{\top}=\mathbb{E}\bm{A}\bm{A}^{\top}$ and $\mathbb{E}\bm{A}\bm{B}^{\top}$ is symmetric.
+Then
+$\mathbb{E}\bm{A}\bm{B}^{\top}\preceq\mathbb{E}\bm{A}\bm{A}^{\top}.$
+Now let $\bm{A}$ be $\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})$ and $\bm{B}$ be $\frac{1}{n}\sum_{t=1}^{n}\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})$.
+By symmetry of $\bm{A}$ and $\bm{B}$, $\mathbb{E}\bm{B}\bm{B}^{\top}=\mathbb{E}\bm{A}\bm{A}^{\top}$ and $\mathbb{E}\bm{A}\bm{B}^{\top}=\mathbb{E}\bm{B}\bm{A}^{\top}$ so $\mathbb{E}\bm{A}\bm{B}^{\top}$ is symmetric.
+Observe that $\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})=\bm{0}$ and fact that $\nabla\mathcal{L}_{1}^{(t_{1})}(\bm{\theta}^{\star}),\nabla\mathcal{L}_{1}^{(t_{2})}(\bm{\theta}^{\star})$ are independent for any $t_{1}\neq t_{2}$.
+Then
+$\mathbb{E}\bm{A}\bm{A}^{\top}$ $=\frac{1}{n^{2}}\sum_{t_{1}=1}^{n}\sum_{t_{2}=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t_{1})}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t_{2})}(\bm{\theta}^{\star})^{\top}$
+$=\frac{1}{n^{2}}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}.$
+Similarly,
+$\mathbb{E}\bm{A}\bm{B}^{\top}=\frac{1}{n^{2}}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}$
+and
+$\mathbb{E}\bm{B}\bm{B}^{\top}=\frac{1}{n^{2}}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}.$
+Invoking Lemma 10, we have that
+$\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{2}^{(t)}(\bm{\theta}^{\star})^{\top}\preceq\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\mathrm{s+r+d+c}}\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})\nabla\mathcal{L}_{1}^{(t)}(\bm{\theta}^{\star})^{\top}.$
+As this holds for every $n$,
+$\bm{V}_{\mathrm{diff}}^{\infty}\preceq\bm{V}_{\mathrm{same}}^{\infty}.$
+The proof is now completed.
+Proof of Lemma 10
+Let $\bm{v}\in\mathbb{R}^{d}$ be an arbitrary vector. It suffices to show
+$\bm{v}^{\top}\left(\mathbb{E}\bm{A}\bm{B}^{\top}\right)\bm{v}\leq\bm{v}^{\top}\left(\mathbb{E}\bm{A}\bm{A}^{\top}\right)\bm{v}$
+By Cauchy-Schwarz inequality and linearity of expectations,
+$\bm{v}^{\top}\left(\mathbb{E}\bm{A}\bm{B}^{\top}\right)\bm{v}$ $=\mathbb{E}\bm{v}^{\top}\bm{A}\bm{B}^{\top}\bm{v}$
+$\leq\sqrt{\mathbb{E}\bm{v}^{\top}\bm{A}\bm{A}^{\top}\bm{v}}\cdot\sqrt{\mathbb{E}\bm{v}^{\top}\bm{B}\bm{B}^{\top}\bm{v}}$
+$=\mathbb{E}\bm{v}^{\top}\bm{A}\bm{A}^{\top}\bm{v}=\bm{v}^{\top}\left(\mathbb{E}\bm{A}\bm{A}^{\top}\right)\bm{v}.$
+Appendix E Auxiliary lemmas
 In this section, we gather some auxiliary results that are useful throughout this paper.
@@
 $\leq 2\sum_{i}\sum_{j\neq i}w_{ij}v_{j}^{2}$
-$\leq 2\sum_{i}\max_{i}\sum_{j}w_{ij}\|\bm{v}\|^{2}$
+$\leq 2\sum_{i}\max_{i}\sum_{j}w_{ij}\|\bm{v}\|^{2}.$
 So $\lambda_{1}(\bm{L})=\max_{\bm{v}\in\mathbb{R}^{m},\|\bm{v}\|=1}\bm{v}^{\top}\bm{L}\bm{v}\leq 2\max_{i}\sum_{j}w_{ij}$.∎
-Lemma 11 (A quantitative version of Sylvester’s law of inertia, [Ost59]).
+Lemma 13 (A quantitative version of Sylvester’s law of inertia, [35]).
 For any real symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$ and $\bm{S}\in\mathbb{R}^{n\times n}$ be a non-singular matrix.
@@
 $\bm{L}_{w}\coloneqq\sum_{i>j:(i,j)\in\mathcal{G}}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
-If $w_{ij}>0$ for all $(i,j)\in\mathcal{G}$ and $\mathcal{G}$ is a connected graph, then $\bm{L}_{w}$ is rank $m-1$, $\bm{L}_{w}\bm{1}_{m}=\bm{0}_{m}$ and $\bm{L}_{w}^{\dagger}\bm{1}_{m}=0$.
+If $w_{ij}>0$ for all $(i,j)\in\mathcal{G}$ and $\mathcal{G}$ is a connected graph, then $\bm{L}_{w}$ is rank m1m-1, $\bm{L}_{w}\bm{1}_{m}=\bm{0}_{m}$ and $\bm{L}_{w}^{\dagger}\bm{1}_{m}=0$.
 Moreover for any $i\in[n-1]$, $\lambda_{i}(\bm{L}_{w}^{\dagger})=\lambda_{n-i}(\bm{L}_{w})$.
 Proof.
-The fact that $\bm{L}_{w}$ is rank $m-1$ when $\mathcal{G}$ is connected is well-known.
+The fact that $\bm{L}_{w}$ is rank m1m-1 when $\mathcal{G}$ is connected is well-known.
 See e.g.
-[Spi07] for reference.
+[41] for reference.
 Since $\bm{L}_{w}$ is a real symmetric matrix, it has an eigendecomposition $\bm{L}_{w}=\bm{U}\bm{\Sigma}\bm{U}^{\top}$ and then $\bm{L}_{w}^{\dagger}=\bm{U}\bm{\Sigma}^{\dagger}\bm{U}^{\top}$.
 The rest follows from this decomposition and the form of $\bm{L}_{w}$.∎
+E.1 Proof of Fact 1
+Expanding the probability of the random events, we have
+[…]
+$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\cdot$
+$\qquad\left(\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\right)^{-1}$
+$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{e^{\zeta_{t}^{\star}}(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}.$
+Now consider $\mathbb{P}[X_{ti}\neq X_{tj}]$, we have
+$\mathbb{P}[X_{ti}\neq X_{tj}]$ $=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}$
+$=\frac{e^{\theta_{j}^{\star}-\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}}}{(1+e^{\theta_{i}^{\star}-\zeta_{t}^{\star}})(1+e^{\theta_{j}^{\star}-\zeta_{t}^{\star}})}.$ (60)
+Let $f:[1/\kappa_{2},\kappa_{2}]^{2}\rightarrow\mathbb{R}$ defined by
+$f(a,b)\coloneqq\frac{a+b}{(1+a)(1+b)}.$
+Its partial derivatives are
+$\frac{\partial}{\partial a}f(a,b)=\frac{b^{2}-1}{(1+a)^{2}(1+b)^{2}}\quad\text{and}\quad\frac{\partial}{\partial b}f(a,b)=\frac{a^{2}-1}{(1+a)^{2}(1+b)^{2}}.$
+It is now easy to see that the minimum or maximum of $f$ can only happen if $(a,b)=(1,1)$ or $(a,b)\in\{1/\kappa_{2},\kappa_{2}\}^{2}$.
+After comparing the value of $f$ at these points, we conclude that $f$ achieves minimum at
+$f(1/\kappa_{2},1/\kappa_{2})=f(\kappa_{2},\kappa_{2})=\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}.$
+By the definition of $\kappa_{2}$, $|\theta_{l}^{\star}-\zeta_{t}^{\star}|\leq\log(\kappa_{2})$ for any $l\in[m]$. Then (60) fits the definition of $f$ and the proof is completed.